### v1

# RadarLLM: Adapting Pretrained Large Language Models for Marine Radar Target Detection with Preference-aware Loss

Abstract
Recent advances in pre-trained large language models (LLMs) have demonstrated their capacities to capture universal knowledge, making them promising general-purpose optimization solvers for wireless signal processing. Motivated by these findings, we take the first step towards fine-tuning pre-trained LLMs for the effective analysis of radar signal features in marine target detection tasks. Nevertheless, directly fine-tuning pre-trained LLMs on marine target detection tasks tends to suffer from pronounced overfitting, particularly in challenging low signal-to-clutter ratio (SCR) scenarios. This overfitting primarily stems from the model’s tendency to memorize spurious or noisy feature patterns rather than learning discriminative structures that generalize well to unseen data. To address this challenge, we introduce RadarLLM, a novel fine-tuning framework that utilizes an effective preference-aware loss. Unlike conventional training strategies that uniformly optimize all feature tokens, this loss function selectively optimizes different feature patches based on their online evaluated learning values, thus guiding the model to focus on the most generalizable patterns during optimization. We theoretically demonstrate the effectiveness of the evaluated learning values by transforming the problem as selecting useful feature tokens. Extensive experiments on real-world marine radar datasets show that 1) the proposed loss function is much better than the original one, with particularly significant gains in challenging low SCR scenarios and 2) RadarLLM consistently outperforms state-of-the-art baselines across diverse detection scenarios, with particularly notable gains under limited training data conditions.
Index Terms:
Marine small target detection, radar signal processing, pre-trained large language models (LLMs), preference-aware loss.
I Introduction
With the growth of global economic integration, maritime transportation has become crucial for international trade and security. Real-time and precise detection of maritime targets is essential for intelligent monitoring systems and various applications such as search and rescue, environmental monitoring, and national security. Radar technology, known for its ability to operate under diverse weather conditions and around the clock, plays a key role in maritime target detection. However, radar systems are challenged by sea clutter, which significantly obscures small targets and complicates detection. The complex nature of sea clutter makes it difficult to distinguish weak targets from sea clutter, posing a persistent challenge in radar target detection. Detecting small maritime targets, including small boats, frogmen, and fairway buoys, remains an active research area, as these targets have small radar cross sections and are difficult to identify in cluttered environments.
The early methods for detecting marine targets are typically developed based on the fundamental concept of constant false alarm rate (CFAR), with specific implementations such as CA-CFAR and GO-CFAR. These methods operate by estimating clutter statistics from reference cells to dynamically adjust detection thresholds. Although computationally efficient, these methods exhibit significant limitations in complex marine environments. Empirical studies reveal that their actual false alarm rates (FAR) often deviate drastically from preset parameters, sometimes exceeding theoretical values by orders of magnitude. This inadequacy stems from their reliance on simplistic statistical assumptions, which fail to hold in heterogeneous sea clutter scenarios characterized by spatially varying distributions and nonstationary behaviors. With the development of modern radars with high range and velocity resolution, approaches leveraging multi-domain features of radar echoes have gained significant attention, utilizing phase, Doppler, and time-frequency domain characteristics to distinguish targets from sea clutter. Many researchers have actively engaged in developing approaches that utilize hand-crafted features and machine learning classifiers [1, 2, 3, 4, 5, 6]. However, these methods rely heavily on domain expertise and hand-crafted heuristics, often struggling to capture high-dimensional and complex patterns in the signal feature.
With the rapid development of deep learning technology, deep learning (DL) and neural networks (NN) have gained huge popularity in the computer vision (CV) society and natural language processing (NLP) society. Moreover, DL and NN approaches are also introduced for small marine target detection. These methods primarily utilize radar signal features to train deep neural networks, eliminating the need for precise mathematical models of clutter and targets. As a result, they somewhat reduce the impact of complex environmental conditions on target detection performance by learning to extract meaningful and distinctive features in a high-dimensional space. Some studies use convolutional neural networks (CNNs) [7, 8, 9, 10, 11, 12, 13], recurrent neural networks (RNNs) [14], and graph neural networks (GNNs) [15, 16, 17]. Among all the methods mentioned above, detectors based on sequence features and RNNs [14] or CNNs [10] offer an excellent trade-off between inference speed and detection performance.
Despite recent progress, compact NN models with small parameter scales and randomly initialized weights often exhibit limited learning capabilities and unstable training processes. These limitations lead to poor detection performance, particularly in scenarios with scarce training data or with significant environmental variability. Extensive recent studies [18, 19, 20, 21, 22, 23, 24] have shown that pre-trained LLM can be a good starting point for wireless signal and time series modeling, as they are pre-trained on massive sequential datasets. Compared to random initialized weights of small models, this starting point may significantly improve model optimization stability and effectiveness. Moreover, the self-attention mechanism in LLM exhibits structural similarity to Principal Component Analysis (PCA) [18], demonstrating strong generality in extracting key features from high-dimensional sequence data. This insight opens a very promising direction for radar target detection by fine-tuning pre-trained LLMs, offering the potential for performance improvements.
Our preliminary study [25] shows that fine-tuning a pre-trained GPT2 model to analyze sequence features yields state-of-the-art detection performance. Extensive ablation experiments further confirm that removing the GPT2 model or substituting it with a randomly initialized Transformer encoder significantly degrades the detection rate.
Although powerful and promising, some challenges still exist. First, the large scale of LLM requires lightweight and efficient fine-tuning strategies, which are essential for practical deployment, especially in resource-constrained settings. Second, and more critically, for radar applications, adapting LLMs to process sequence features in low signal-to-clutter ratio (SCR) conditions is far from a routine task, often leading to severe overfitting issues. This phenomenon is fundamentally attributed to the intrinsic characteristics of sequence features:
  • Significant Clutter-Induced Noise: In low SCR environments, small targets are frequently obscured by ocean waves. This causes large noise in sequence features, especially in phase and amplitude features, increasing the likelihood of the model overfitting to irrelevant and noisy information.
  • Feature Pattern Convergence Imbalance: Various patterns in different sequence features often exhibit different convergence rates during fine-tuning. For example, simple and regular patterns may rapidly reach convergence during model training, and then exhibit a tendency for overfitting, whereas other valuable patterns may require more iterations to achieve convergence.
To address the aforementioned challenges, this paper introduces RadarLLM, an innovative framework for learning from radar echo signals. First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens. Second, we train a small reference model based on the validation samples and calculate the loss for each feature token of the training samples to estimate its generalization potential. Third, we construct a target detector based on a pre-trained LLM, then fine-tune the model based on low-rank adaptation (LoRA) [26] and a novel preference-aware loss function, which is inspired by the work in NLP [27] and CV [28]. This loss introduces a selective training strategy by assigning training weights to different feature patches according to their learning values, which greatly improving model performance in low SCR environments. Finally, we retrain a classification head to enhance detection performance and control the false alarm rate.
Here we summarize our key contributions as follows:
  1. 1.
    We propose RadarLLM, a novel framework combining a lightweight reference model and a pre-trained LLM for marine radar target detection. To the best of our knowledge, we are the first to show that well-known LLM can be a strong starting point for intelligent radar signal processing.
  2. 2.
    For fine-tuning of the framework, we propose a novel preference-aware loss function which greatly improves RadarLLM’s performance in a low SCR environment. RadarLLM can selectively optimize on useful feature tokens with high learning values while reducing the risk of overfitting.
  3. 3.
    Our extensive experiments confirm the effectiveness of RadarLLM. It attains new state-of-the-art detection performance under sufficient training data, and constrained training data settings on popular marine radar datasets.
The remainder of this paper is structured as follows. Section II briefly summarizes the related work. Section III presents the detailed operating procedure of the proposed RadarLLM. Section IV introduces the experimental results and analysis. Section V concludes our work.
II Related Work
II-A Deep Learning Powered Marine Radar Target Detection
Deep learning models with elaborately crafted architectures have demonstrated great promise in marine target detection. Among them, Chen et al. [7] design a dual-channel CNN (DCCNN)-based structure detector that extracts both amplitude and time-frequency information from signals to achieve target detection. Qu et al. [8] introduce a CNN architecture augmented with asymmetric convolution layers to capture deep features of the time-frequency distribution. Xu et al. [9] propose a target detector in the case of limited training samples, utilizing pre-trained CNN to extract features from time-frequency spectrogram. Wan et al. [14] propose a sequence-based target detection framework that leverages instantaneous phase, Doppler spectrum, and short-time Fourier transform features, in conjunction with a bidirectional long short-term memory (Bi-LSTM) network. Wang et al. [11] propose a target detector based on complex value U-Net (CV-UNet), which realizes clutter suppression based on the amplitude-phase characteristics of radar echoes, and then achieves target detection. Su et al. [17] introduce a graph neural network (GNN) approach for radar target detection, which constructs spatio-temporal adjacency matrices, extracts hierarchical features through convolutional graph operations, and outputs detection results via nonlinear vertex embeddings.
In addition to traditional supervised learning, semi-supervised, unsupervised, and incremental learning approaches are proposed to further advance radar target detection. Wang et al. [29] propose a self-evolving framework for maritime radar target detection using semi-supervised learning, which enhances detection performance through unsupervised sample selection, data augmentation, and model optimization. Xia et al. [10] introduce an unsupervised contrastive learning framework that learns discriminative representations between targets and clutter from unlabeled data. Wang et al. [30] propose an incremental learning-based target detection method, which can continuously adapt the NN model in real time according to environmental changes. These methods provide compelling evidence that deep learning can significantly enhance detection performance without the need for explicit mathematical modeling. However, it is important to recognize certain limitations inherent in these methods. Current methods rely on relatively small NN models, which possess limited feature recognition capabilities and are highly sensitive to environmental variations.
II-B Language Model Powered Signal Processing Task
Recently, LLMs such as ChatGPT and DeepSeek have demonstrated exceptional capabilities in natural language understanding, code generation, math problem solving, and human-like text generation [31]. Some studies also demonstrate their potential in the signal processing (SP) task. For example, Zhou et al. [18] propose a unified time series analysis framework by fine-tuning frozen GPT2 [32] and achieving SOTA performance on various datasets. Liu et al. [24] propose a GPT2-empowered channel prediction framework to improve prediction accuracy. Sheng et al. [33] utilize the GPT2 model to develop an effective and robust beam prediction method. Zheng and Dai [21] propose a multi-task LLM framework to satisfy the requirements of different wireless communication tasks. IOT-LLM and Penetrative AI [34, 35] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks. However, recent studies [36] have raised concerns about the universal applicability of LLM in signal processing tasks. LLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks. In marine radar target detection under low SCR conditions, directly fine-tuning a pre-trained LLM often leads to severe overfitting issues, limiting its parameter transfer capability. We counteract this shortfall with a preference-aware loss through selectively training on useful feature patches.
III Proposed Method
Figure 1: Overview of our RadarLLM framework.
III-A Overview of the RadarLLM Framework
In this study, we present RadarLLM, a LLM–empowered framework for marine radar target detection. The framework introduces a selective patch-level fine-tuning strategy that counteracts the severe overfitting problem caused by the low SCR environment. The overall pipeline of RadarLLM is illustrated in Fig. 1 and shown in Algorithm 1.
First, five complementary sequence features are extracted from radar echoes and partitioned into uniform patches to form a multi-patch representation. Next, a lightweight reference model rapidly scores each patch to quantify its generalization potential. Then, we fine-tune a pre-trained LLM backbone via a preference-aware loss and Low-Rank Adaptation (LoRA). Finally, an autoencoder-based classification head is re-trained to enhance detection performance and control the false alarm rate. The detailed design and rationale of each module are elaborated in the following subsections.
Algorithm 1 RadarLLM
1:
2:Radar echo signals $\{x(n)\}_{n=1}^{N}$ and Target labels $\{y_{b}\}_{b=1}^{B}$
3:Pre-trained LLM parameters $\theta_{\text{LLM}}$
4:Stage 1: Sequence Feature Extraction
  1. 1.
    Extract IP, DSE, SMS, Amp, and DP features from the observation vector, respectively, see Eq.(3), (4), (5),(7), and (9).
  2. 2.
    Patch the five features into feature tokens, see Fig. 3.
5:Stage 2: Reference Model Training
  1. 1.
    Train a lightweight reference model $\theta_{r}$ based on validation samples and CE loss.
  2. 2.
    Compute the loss for each feature token of the training samples based on the predicted output probabilities of reference model, see Eq.(20).
6:Stage 3: Fine-tune LLM for Target Detection
  1. 1.
    Compute the loss for each feature token of the training samples based on the predicted output probabilities of current model, see Eq.(21).
  2. 2.
    Fine-tune LLM for target detection via preference-aware loss function, see Eq.(22) and Eq.(24).
7:Stage 4: Classification Head Retraining
  1. 1.
    Train autoencoder classification head based on the weighted sum of reconstruction loss and CE loss, see Eq.(30), Eq.(31), and Eq.(32).
  2. 2.
    Adjust detection threshold $\eta$ for controllable FAR, see Eq.(33).
8:
9:Fine-tuned RadarLLM model with controllable FAR.
III-B Stage 1: Sequence Feature Extraction
When the radar transmits coherent pulses toward the sea surface, it receives a sequence of echo signals from each range resolution cell. In clutter-dominated cells, these echoes primarily consist of sea surface backscatter and thermal noise, whereas in target-present cells, additional reflections from the target are superimposed. The received echo sequence $x$ from a resolution unit can be decomposed into observation vectors ${x}_{i}$ through the following segmentation:
$x_{i}=[x(M\cdot(i-1)+m)]_{m=1}^{N},\quad i=1,2,$... (1)
where $N$ represents the length of the observation window and $M$ denotes the sliding step between consecutive observation vectors. Target detection is formulated as a binary hypothesis testing problem:
[…]
where $H_{0}$ is the null hypothesis that indicates target absence in received signal $x(n)$, and $H_{1}$ is the alternative hypothesis that indicates target presence. In the cell under test, $s(n)$ and $c(n)$ denote target echos and sea clutter echos, respectively. Meanwhile, in the reference cells, $x_{p}(n)$ and $c_{p}(n)$ denote received echos and clutter echos, respectively. $P$ is the number of reference cells. To effectively distinguish between the target echo and clutter echo, five unique sequence features are extracted from the time, phase, Doppler, and time frequency domains, as described below:
  • Instantaneous Phase (IP): The instantaneous phase of the radar echo signal $x(n)$ is calculated as follow:
    $\phi(n)=\arg[x(n)],$ (3)
    where $\arg(.)$ represents the phase of a complex variable.
  • Doppler Spectrum Entropy (DSE): For the echo signal $x(n)$, its Doppler amplitude spectrum can be represented as:
    $F(f_{d})=\frac{1}{\sqrt{N}}\left|\sum_{n=0}^{N-1}x(n)\exp\left(-j2\pi f_{d}nT_{s}\right)\right|,$ (4)
    where $f_{d}(-\frac{1}{2T_{s}}\leq f_{d}\leq\frac{1}{2T_{s}})$ represents the Doppler frequency, and $T_{s}$ represents the pulse repetition interval for the radar system. Furthermore, the entropy of the Doppler spectrum is calculated as:
    $\text{DSE}(F)=-\tilde{F}\left(f_{d}\right)\log\tilde{F}\left(f_{d}\right),$ (5)
    where $\tilde{F}\left(f_{d}\right)=\frac{F(f_{d})}{{\textstyle\sum_{f_{d}}}F(f_{d})}$.
  • Short-Time Fourier Transform Magnitude Spectrum (SMS): Due to the non-stationary characteristics of the radar echo signal, short-time fourier transform (STFT) provides a more comprehensive understanding compared to Doppler transform, which is calculated as:
    $S(k,m)=\sum_{n=-\infty}^{\infty}x(n)w(n-m)e^{-j2\pi\frac{kn}{\Omega}},$ (6)
    where $w(.)$ is the window function, $\Omega$ is the number of frequency bins, $m$ represents the time index, and $k$ is the frequency index. To make the features suitable for input into a sequence neural network, we compute the time-averaged magnitude of $S(k,m)$ and convert it into a decibel (dB) scale. Specifically, for each frequency index $k$, the following computation is performed:
    $\text{SMS}(k)=10\log_{10}(\sum_{m=0}^{N-1}|S(k,m)|).$ (7)
  • Amplitude (Amp): The instantaneous amplitude is computed by calculating the magnitude of the complex radar echo signal:
    $A(n)=|x(n)|.$ (8)
  • Doppler Phase (DP): The Doppler phase is extracted from the Doppler transform of echo signal:
    $\text{DP}(f_{d})=\arg(\sum_{n=0}^{N-1}x(n)\exp\left(-j2\pi f_{d}nT_{s}\right)).$ (9)
As shown in Fig. 2, the five sequence features reveal clear distinctions between target and clutter samples.
(a) IP feature
(b) DSE feature
(c) SMS feature
(d) Amp feature
(e) DP feature (Target)
(f) DP feature (Clutter)
Figure 2: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX database.
III-C Structure of the RadarLLM Network
Figure 3: An illustration of the patching operation.
Initially, for a batch of $B$ samples, the input feature matrix is constructed by concatenating the IP feature $F_{\mathrm{IP}}$, DSE feature $F_{\mathrm{DSE}}$, SMS feature $F_{\mathrm{SMS}}$, amplitude feature $F_{\mathrm{Amp}}$, and DP features $F_{\mathrm{DP}}$.
$F=\text{Concat}(F_{\mathrm{IP}},F_{\mathrm{DSE}},F_{\mathrm{SMS}},F_{\mathrm{Amp}},F_{\mathrm{DP}}).$ (10)
Each feature in F{F} is subsequently partitioned into non-overlapping patches of length $L$. If the final patch is shorter than $L$, zero padding is applied to ensure uniform length. The resulting $K$ patches are concatenated to construct the final input tensor ${F}^{\mathrm{P}}\in\mathbb{R}^{B\times K\times L}$, where $K=5\left\lceil\frac{N}{L}\right\rceil$. An illustration of this patching process is provided in Fig. 3.
To adapt the feature patches to the input format required by the pre-trained language model, we employ a fully connected layer to project ${F}^{\mathrm{P}}$ into an embedded representation ${F}^{\mathrm{EB}}\in\mathbb{R}^{B\times K\times L^{\prime}}$. This is followed by the addition of a positional encoding ${E}^{\mathrm{pos}}$, which incorporates either relative or absolute positional information into the embedded features:
${E}^{\mathrm{pos}}_{b,k,2l}=\sin(\frac{k}{10000^{2l/L^{\prime}}}),\ {E}^{\mathrm{pos}}_{b,k,2l+1}=\cos(\frac{k}{10000^{2l/L^{\prime}}}),$ (11)
where $b$ represents the $b$-th sample, $k$ represents the position index of the token, and $l$ denotes the feature dimension index. The positional encoding ${E}^{\mathrm{pos}}$ is added to the embedding ${F}^{\mathrm{EB}}$ to produce a enriched embedding:
${F}^{\mathrm{PE}}={F}^{\mathrm{EB}}+{E}^{\mathrm{pos}}.$ (12)
The enriched embedding ${F}^{\mathrm{PE}}$ is passed through initial several layers of the GPT2 encoder for further feature extraction. To efficiently fine-tune the backbone model, we apply Low-Rank Adaptation (LoRA) [26] to the query ($\mathbf{W}_{Q}$) and value matrices ($\mathbf{W}_{V}$) in the multi-head attention layers and directly fine-tune the Layer Normalization (LayerNorm) parameters. The LoRA fine-tuning process is illustrated in Fig. 4. This strategy introduces lightweight, trainable low-rank matrices to the fixed pre-trained weights, substantially reducing the computational and memory requirements compared to full parameter fine-tuning. Formally, the adapted model generates rich contextual feature representations as follows:
${F}^{\mathrm{LLM}}=\operatorname{LLM}_{\mathrm{LoRA}}\left({F}^{\mathrm{PE}}\right)\in\mathbb{R}^{B\times K\times L^{\prime}},$ (13)
where $\operatorname{LLM}_{\mathrm{LoRA}}(\cdot)$ represents the backbone network of the language model encoder enhanced with LoRA modules. To stabilize the training process, a Layer Normalization (LayerNorm) operation is applied to the output representation ${F}^{\mathrm{LLM}}$.
$F^{\mathrm{LN}}=\mathrm{LayerNorm}({F}^{\mathrm{LLM}}),$ (14)
where $\mathrm{LayerNorm}(\cdot)$ is defined element-wise for each position $k$ and feature dimension $l$ as:
$F_{b,k,l}^{\mathrm{LN}}=\gamma_{b,l}\cdot\frac{F_{b,k,l}^{\mathrm{LLM}}-\mu_{b,l}}{\sqrt{\sigma_{b,l}^{2}+\epsilon}}+\beta_{b,l},$ (15)
where $\mu_{b,l}$ and $\sigma_{b,l}^{2}$ is computed as:
$\mu_{b,l}=\frac{1}{K}\sum_{k=1}^{K}F_{b,k,l}^{\mathrm{LLM}},\quad\sigma_{b,l}^{2}=\frac{1}{K}\sum_{k=1}^{K}\left(F_{b,k,l}^{\mathrm{LLM}}-\mu_{b,l}\right)^{2}.$ (16)
Here, $\gamma_{b,l}$ and $\beta_{b,l}$ are learnable parameters, and $\epsilon$ is a small constant for numerical stability. $F^{\mathrm{LN}}$are then passed through two FC layers to obtain the final output:
$F^{\mathrm{fi}}=FC(FC(F^{\mathrm{LLM}}))\in\mathbb{R}^{B\times K\times 2}.$ (17)
Figure 4: Illustration of the LoRA fine-tuning process.
III-D Stage 2: Compute Loss for Each Token Based on the Reference Model
The architecture of the reference model is illustrated in Fig. 5. As described in Section III-C, the input patches are first projected into a latent Transformer space of dimension $D$ and augmented with learnable additive positional embeddings. This process yields the enriched representation $F^{\mathrm{PE}^{\prime}}\in\mathbb{R}^{B\times K\times D}$, which is subsequently processed by a Transformer encoder. Within the encoder, the multi-head attention (MHA) mechanism functions as follows: the input $F^{\mathrm{PE}^{\prime}}$ is linearly projected into $H$ independent attention heads. For each head $h$, query, key, and value matrices $\mathbf{W}_{Q_{h}}\in\mathbb{R}^{B\times D\times d_{k}}$ $\mathbf{W}_{K}\in\mathbb{R}^{B\times D\times d_{k}}$ $\mathbf{W}_{V}\in\mathbb{R}^{B\times D\times d_{v}}$ are used to compute:
$\mathbf{Q}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{Q_{h}},\quad\mathbf{K}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{K_{h}},\quad\mathbf{V}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{V_{h}},$ (18)
where $d_{k}=d_{v}=D/H$. The attention output of each head is then computed as:
$\mathbf{O}_{h}=\operatorname{Softmax}\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}{}^{\bm{\top}}}{\sqrt{d_{k}}}\right)\mathbf{V}_{h}.$ (19)
Outputs from all heads are concatenated along the feature dimension and projected back to dimension $D$. The MHA block further incorporates Batch Normalization layers and a feed-forward network (FFN) equipped with residual connections, resulting in the final encoded representation $F^{\mathrm{TF}}\in\mathbb{R}^{B\times K\times D}$. Subsequently, LayerNorm is applied, followed by two fully connected layers to yield the final output $F^{\mathrm{fi}^{\prime}}\in\mathbb{R}^{B\times K\times 2}$.
Figure 5: Overview of the reference model.
III-E Stage 3: Fine-tune RadarLLM Network Based on the Preference-aware Loss
III-E1 Methodology
To improve fine-tuning of LLMs for radar target detection in low SCR scenarios, we propose a token-level preference-aware loss function. In such conditions, many feature patches are either severely corrupted by noise or trivially easy, offering limited value for effective learning. Uniformly optimizing all patches risks overfitting by forcing the model to learn from uninformative or misleading inputs. To address this, we introduce a dynamic weighting mechanism that prioritizes tokens with higher learning value, those expected to contribute more to generalization. Specifically, our approach utilizes a dual-model framework consisting of a lightweight reference model $\theta_{r}$ and a high-capacity target model $\theta_{t}$, which leverages LLMs.
  • Reference Model: The reference model is built with a lightweight Transformer encoder and trained solely on validation samples using standard cross-entropy loss. By leveraging the reference model, we can estimate the generalization potential of each feature token in the training samples. A lower reference loss indicates that the token can be effectively captured by a lightweight model, suggesting strong generalization potential. In contrast, a higher reference loss typically reflects excessive noise, making the token less useful for improving detection performance.
  • Target Model: A significantly larger model built upon a pre-trained LLM, parameterized by $\theta_{t}$, as described in Section III-C.
For a batch of training samples, we compute the token-level cross-entropy losses for both models. Given a feature token $F_{b,k}^{\mathrm{fi}}$, the corresponding two losses are as follows:
  • Reference Loss:
    $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})=-\sum_{i}y_{b,k}^{i}\log(\hat{y}_{b,k}^{i,\theta_{r}}),$ (20)
    where $\hat{y}_{b,k}^{i,\theta_{r}}$ is the prediction from $\theta_{r}$.
  • Target Loss:
    $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})=-\sum_{i}y_{b,k}^{i}\log(\hat{y}_{b,k}^{i,\theta_{t}}),$ (21)
    where $\hat{y}_{b,k}^{i,\theta_{t}}$ is generated by the fine-tuned much larger model $\theta_{t}$.
The core of our method lies in computing a token importance score $s_{b,k}$ based on the discrepancy between the target loss and reference loss:
$s_{b,k}=\mathrm{ReLU}(\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})-\alpha\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})),$ (22)
where $\alpha$ is a scaling coefficient, and $\mathrm{ReLU}(\cdot)$ ensures that only tokens with positive excess loss are trained. Tokens with low importance score are not prioritized for optimization because they have already been well learned, as indicated by a low target loss $\mathcal{L}_{\theta_{{t}},b,k}$, or with large noise, as reflected by a high reference loss $\mathcal{L}_{\theta_{{r}},b,k}$. The $\mathrm{ReLU}(\cdot)$ function is defined as:
$\operatorname{ReLU}(x)=\left\{\begin{array}[]{ll}x&x>0\\ 0&x\leq 0\end{array}\right$. (23)
More detailedly, this strategy results in the following selective training scheme:
  • If the target loss is much larger than the reference loss, the token is deemed with large learning value, and is therefore prioritized for training.
  • If the reference loss is large, the token shows weak generalization potential and a high risk of overfitting; therefore, its training weight is reduced during optimization.
  • If the target loss is already small, the token has been sufficiently learned, and further training brings little gain, the training weight is also reduced to avoid redundant updates.
The final training loss is a weighted sum over all feature tokens across the mini-batch, where each token’s contribution is scaled by a token-level importance score:
$\mathcal{L}_{\mathrm{final}}=\frac{1}{B*K}\sum_{b=1}^{B}\sum_{k=1}^{K}s_{b,k}\cdot\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k}).$ (24)
where $B$ and $K$ denote the batch size and the number of feature tokens, respectively.
III-E2 Theoretical Insight of the Effectiveness of Feature Token Importance
We begin by formalizing the task of selecting informative feature patches for training. At each step $t$, given a candidate set $B_{t}$, our objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ achieves the maximal reduction in generalization loss on the test set $D_{\mathrm{te}}$. This can be formally formulated as the following optimization problem:
$\arg\min_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}-\log p\left(y^{\text{te }}\mid F^{\text{te }};D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right).$ (25)
Directly optimizing Eq.(25) is computationally prohibitive, as it entails retraining the model for each candidate feature patch. To derive a tractable surrogate, we reformulate Eq.(25) using Bayes’ theorem:
$-\log p\!\bigl{(}y^{\mathrm{te}}\mid F^{\mathrm{te}};D_{t}\!\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\bigr{)}$
$\quad=-\log\frac{p\!\bigl{(}y,y^{\mathrm{te}}\mid F,F^{\mathrm{te}},D_{t}\bigr{)}}{p\!\bigl{(}y\mid F;D_{t}\bigr{)}}.$ (26)
Assuming that each label is statistically independent of the other pairs in the corpus when conditioned on its own feature vector, the joint term in the numerator can be expressed as:
$p\!\bigl{(}y,y^{\mathrm{te}}\mid F,F^{\mathrm{te}},D_{t}\bigr{)}\\ =p\!\bigl{(}y\mid F;F^{\mathrm{te}},y^{\mathrm{te}},D_{t}\bigr{)}\,p\!\bigl{(}y^{\mathrm{te}}\mid F^{\mathrm{te}};F,D_{t}\bigr{)}\\ =p\!\bigl{(}y\mid F;y^{\mathrm{te}},F^{\mathrm{te}},D_{t}\bigr{)}\,p\!\bigl{(}y^{\mathrm{te}}\mid F^{\mathrm{te}};D_{t}\bigr{)}.$ (27)
Substituting Eq.(27) into Eq.(26) gives:
$\log p\!\left(y^{\text{te}}\mid F^{\text{te}};\,D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right)$ (28)
$=\log\frac{p\left(y\mid F;y^{\text{te }},F^{\text{te }},D_{t}\right)p\left(y^{\text{te }}\mid F^{\text{te }},D_{t}\right)}{p\left(y\mid F;D_{t}\right)}$
$\;\propto\;\;\log p\left(y\mid F;y^{\mathrm{te}},F^{\mathrm{te}},D_{t}\right)-\log p\left(y\mid F;D_{t}\right)$
$\;\propto\;\;\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t},D_{\text{te}}\right].$
In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the test samples in Eq.(25) can be approximated by the following objective:
$\arg\max_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right].$ (29)
Eq.(29) indicates that feature patches with larger loss differences contribute more significantly to improve the model performance on test set; thus, the loss difference can be interpreted as the learning value of each feature patch. However, since accessing the test set for retraining on $D_{t}\cup D_{\mathrm{test}}$ is prohibited and computationally impractical, we approximate the second term using a lightweight reference model trained solely on validation set $D_{\mathrm{val}}$. Consequently, the practical token weight is derived as shown in Eq.(22).
III-F Stage 4: Retrain Classification Head Based on Autoencoder Network
As illustrated in Fig. 1, the output of the LLM-based model has a shape of $[B,K,2]$, representing predictions across multiple tokens. This token-level output cannot be directly used for detection due to the absence of a unified decision representation. To address this, we propose retraining a classifier based on the learned representation $F^{\mathrm{LN}}$. Specifically, we employ an autoencoder-based classification head designed to extract the most informative features from the high-dimensional input, thereby enabling effective and robust decision-making. The proposed autoencoder-based classification head comprises three key components: (1) an encoder, consisting of convolutional layers followed by fully connected layers, which extracts compact and discriminative feature representations from $F^{\mathrm{LN}}$; (2) a decoder, symmetrically constructed with convolutional and fully connected layers, which reconstructs the input from the latent space to promote feature consistency and preservation; and (3) a classification head, which operates on the latent vector to perform the final classification. The autoencoder is trained with a dual-objective loss that combines reconstruction loss and classification loss. The reconstruction loss ensures that the latent representation retains critical information by minimizing the discrepancy between the original input $F^{\mathrm{LN}}$ and its reconstruction $\hat{F}^{\mathrm{LN}}$:
$\mathcal{L}_{\text{recon }}=\left\|F^{\mathrm{LN}}-\hat{F}^{\mathrm{LN}}\right\|_{2}^{2}.$ (30)
The classification loss ensures accurate predictions by aligning the predicted probability distribution $\hat{y}_{b}^{i}$ with the ground-truth label $y_{b}^{i}$.
$\mathcal{L}_{\text{ce}}=-\frac{1}{B}\sum_{b=1}^{B}\sum_{i=1}^{2}y_{b}^{i}\log\hat{y}_{b}^{i}.$ (31)
To jointly optimize both the reconstruction and classification objectives, we adopt a task uncertainty weighting strategy [37], which adaptively adjusts the contribution of each loss term based on its inherent uncertainty. The total loss function is expressed as:
[…]
where $\sigma_{\text{recon}}$ and $\sigma_{\text{ce}}$ are learnable parameters representing the task-dependent uncertainties for the reconstruction and classification losses, respectively. This formulation allows the model to balance the learning dynamics of both tasks in a principled and data-driven manner.
To control the false alarm rate during inference, we sort the first element in the softmax output for all test clutter samples in descending order, yielding a one-dimensional array $O_{\mathrm{sorted}}$. The detection threshold $\eta$ is then determined based on the desired false alarm rate $P_{fa}^{\text{d}}$ as follows:
$\begin{array}[]{l}\eta=O_{\mathrm{sorted}}(x),\ x=\lceil P_{fa}^{\text{d}}\times N_{\text{clutter}}\rceil,\end{array}$ (33)
where $O_{\mathrm{sorted}}(x)$ denotes the $x$-th largest output in the sorted array, $N_{\text{clutter}}$ is the total number of test clutter samples, and $\lceil\cdot\rceil$ is the ceiling operation to ensure integer indexing.
Figure 6: Overview of the autoencoder model.
IV Experiments
In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarLLM’s detection performance under three settings: sufficient training data, constrained training data, and zero-shot scenarios.
IV-A Experimental Setup
IV-A1 Dataset
We employ eight benchmark data sequences (Data1–Data8) from the Intelligent PIxel Processing X-band (IPIX) radar archive, summarized in Table I. All datasets were recorded in 1993 by the IPIX radar on Canada’s east coast and include full polarization measurements (HH, HV, VH, and VV). Each set contains 14 range cells sampled at 1 kHz, with 131 072 complex echoes per cell, yielding an observation window of 0.512 s per sample. Returns from the primary range cell constitute target echoes, whereas returns from the remaining clutter-only cells represent sea clutter. To facilitate performance analysis across varying operating conditions, Data1–Data3 correspond to low signal-to-clutter ratio (SCR) scenes, while Data4–Data8 represent high-SCR scenes.
To ensure sufficient training data, we employ overlapped segmentation following the partition rule in Eq. (1), with parameters set to $M=32$ for target cells and $M=128$ for clutter cells. This process generates 4,079 target samples and more than 9,000 clutter samples per dataset. The samples are divided into three groups: (1) a training set using the first 20% of observation time for both the target and clutter cells, (2) a validation set covering 20% to 35% of the observation time, and (3) a test set containing the remaining samples.
TABLE I: Overview of the IPIX Database.
Label Name File name Primary Secondary SCR
1 17 19931107_135603_starea 9 8,10,11 Low
2 25 19931108_213827_starea 7 6,8 Low
3 26 19931108_220902_starea 7 6,8 Low
4 54 19931111_163625_starea 8 7,9,10 High
5 280 19931118_023604_stareC0000 8 7,9,10 High
6 283 19931118_035737_stareC0000 10 8,9,11,12 High
7 311 19931118_162658_stareC0000 7 6,8,9 High
8 320 19931118_174259_stareC0000 7 6,8,9 High
IV-A2 Baselines
To validate the superiority of RadarLLM, eight deep learning-based marine radar target detection methods are implemented as baselines. Among them, RNN, Bi-LSTM, GRU, Transformer, PatchTST, and ResNet18 represent lightweight models that utilize sequence features; OFA uses a large language model backbone to process sequence features; and ADN18 utilizes time-frequency maps and enhanced CNNs for detection.
  • RNN [14]: RNN is a sequential neural architecture with temporal feedback loops. In experiments, the number of RNN layers and the hidden layer nodes are set to 2 and 6, respectively.
  • Bi-LSTM [14]: Bi-LSTM enhances the traditional Long Short-Term Memory (LSTM) by incorporating bidirectional processing. In our implementation, the number of LSTM layers and the number of hidden units per layer are set to 2 and 6, respectively.
  • GRU [38]: GRU employs simplified gating mechanisms compared to LSTM. In experiments, the number of GRU layers and the hidden layer nodes are set to 2 and 8, respectively.
  • Transformer [39]: Transformer evolutionizes sequence processing through self-attention mechanisms. In experiments, the Transformer encoder is configured with a model dimensionality of 32, comprising 2 layers and 8 attention heads. The feed-forward network (FFN) within each layer is designed with a hidden size of 128.
  • PatchTST [40]: PatchTST segments the sequence features into local patches and leverages a Transformer encoder to model them. In experiments, the Transformer encoder is configured with a model dimensionality of 64, comprising 2 layers and 16 attention heads. The feed-forward network (FFN) within each layer is designed with a hidden size of 128.
  • ResNet18 [10]: ResNet18 is a convolutional architecture adapted for analyzing time series by modifying input channels, leveraging residual blocks to capture temporal patterns.
  • OFA [18]: OFA patches input sequence features, feeds them through a frozen GPT2 model with trainable positional embeddings and LayerNorm, and uses a final linear layer for classification.
  • ADN18 [41]: ADN18 enhances CNN model with asymmetric convolutional kernels and attention mechanism to adaptively capture spatial-temporal features in spectrograms derived from Short-Time Fourier Transform.
IV-A3 Evaluation Metrics
In maritime target detection, achieving a high detection rate at an extremely low false alarm rate is imperative, as misinterpreting sea clutter as targets can have significant consequences, particularly in military operations. We evaluate the detection performance of different detectors at a false alarm rate of 0.005, considering our test set only comprises 6,000 clutter samples. Selecting a lower false alarm rate could result in unstable comparisons due to the limited number of clutter samples.
IV-A4 Implementation Details
The experimentally relevant hyperparameters and details are shown in Table II. For patching, the size of feature patches is set to 48. For the pre-trained LLM, the smallest version of GPT2 with $F=768$ feature dimension and the first $L=4$ layers of the GPT2 encoder are deployed. The network hyperparameters of the Autoencoder are shown in Table III. The value of $\alpha$ in Eq.(22) is set to 0.9. All signal feature extraction operations are performed with MATLAB R2016a software. All network training experiments are conducted on a system equipped with an E5-2695v3 CPU, an NVIDIA 3090Ti GPU, and 64 GB of RAM.
TABLE II: Hyperparameters for network training.
Parameter Value
Batch size 64
Eval batch size 400
Epochs for LLM fine-tuning 500
Epochs for Autoencoder training 300
Optimizer Adam (betas=(0.9,0.999))
Learning rate for LLM fine-tuning 0.0001
Learning rate for Autoencoder training 0.00001
TABLE III: Hyperparameters for Autoencoder network.
Block Layer Filter Stride Padding Output
[…]
Conv6+ReLU 16,1*1 1 Same 55*16
Flatten+Fully-Connected1 - - - 64
Fully-Connected2 - - - 20
Decoder Fully-Connected3+Reshape - - - 55*16
Conv1+ReLU 32,1*1 1 1 55*32
Conv2+ReLU 64,1*1 1 1 55*64
Conv3+ReLU 128,1*1 1 1 55*128
Conv4+ReLU 256,1*1 1 1 55*256
Conv5+ReLU 512,1*1 1 1 55*512
Conv6+ReLU 768,1*1 1 1 55*768
Classifier GeLU+LayerNorm - - - 20
Fully-Connected4 - - - 2
IV-B Comparison of Detection Performance, Speed, and Network Parameters Under Sufficient Training Samples
(1). HH
(2). HV
(3). VH
(4). VV
Figure 7: Comparison of model detection performance on various datasets.
TABLE IV: Detection performance of different methods when the false alarm rate is 0.005. The best and second-best detection results are highlighted in red and blue, respectively.
Detector $H$ HV VH $V$ AVG
RNN [14] 46.04 46.66 38.18 17.04 36.98
Bi-LSTM [14] 70.39 73.15 76.48 60.89 72.48
GRU [38] 76.46 77.55 73.96 67.48 73.86
Transformer [39] 52.95 71.73 67.32 55.67 61.92
PatchTST [40] 75.23 75.72 77.49 66.59 73.76
Resnet18 [10] 66.24 80.96 82.31 73.97 75.87
OFA [18] 74.19 80.38 79.71 70.36 76.16
ADN18 [41] 79.74 83.50 82.01 78.02 80.82
RadarLLM 82.51 84.81 84.61 75.90 81.96
TABLE V: Network parameters (training parameters/total parameters) and interference time of different methods.
Detector Average interference time on one dataset(s) Total network parameters(M) Training network parameters(M)
RNN [14] 5.3587 0.0003 0.0003
Bi-LSTM [14] 5.6028 0.0016 0.0016
GRU [38] 5.7746 0.0008 0.0008
Transformer [39] 6.0678 0.0611 0.0611
PatchTST [40] 5.3727 0.0939 0.0939
Resnet18 [10] 5.4434 3.8454 3.8454
ADN18 [41] 160.9018 14.3315 14.3315
RadarLLM 5.6561 69.3932 2.4567
Table IV reports the detection rates achieved by RadarLLM and the eight baseline detectors introduced in Section IV-A2. All results are obtained with a fixed false alarm rate of 0.005 on the eight IPIX datasets. Compared to the best competing sequence feature detector, RadarLLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively. Figure IV presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates. We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone. Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in average detection rate.
Furthermore, we compare RadarLLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment. The comparison results are summarized in Table V. Although RadarLLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning. Notably, RadarLLM achieves an exceptionally low average inference latency. Specifically, it requires only 5.6561 seconds to process over 9,000 test samples, making it 28.44 times faster than the ADN18 method. This efficiency is primarily attributed to our patching strategy, which substantially reduces the number of feature tokens, and consequently, the computational overhead. Moreover, the intrinsic inference acceleration of the GPT architecture further enhances this performance gain. Together, these advantages make RadarLLM a promising solution for real-time marine target detection.
IV-C Comparison of Detection Performance Under Limited Training Samples With Baselines
Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge. Considering this issue, we conduct experiments under constrained training sample conditions to assess the effectiveness of RadarLLM. The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
In Table VI, we report the detection performance of RadarLLM and several baseline methods under scenarios with limited training data. Compared with the experimental setup described in Section IV-A1, the advantages of RadarLLM become more pronounced. Specifically, RadarLLM achieves a substantial increase of 19.66% in the average detection rate over the previously proposed sequence-feature based detector [14]. This marked improvement can be mainly attributed to three key factors:
  • First, leveraging pre-trained knowledge embedded in the large language model (LLM) provides RadarLLM with a beneficial parameter initialization, facilitating efficient optimization even when training data is limited. This prior knowledge significantly reduces the data dependency typical of conventional training strategies.
  • Second, the substantial parameter scale of RadarLLM inherently grants robustness to distributional variations, enabling superior generalization compared to smaller-scale neural networks. Its extensive capacity allows it to capture complex data structures and efficiently handle variability across different radar detection scenarios.
  • Third, the integration of a preference-aware loss function notably improves RadarLLM’s detection performance, especially in challenging scenarios characterized by low signal-to-clutter ratios (SCR).
TABLE VI: Detection performance under limited training samples of different methods when the false alarm rate is 0.005. The best and second-best detection results are highlighted in red and blue, respectively.
Detector $H$ HV VH $V$ AVG
RNN [14] 35.49 30.13 31.82 15.14 28.15
Bi-LSTM [14] 61.96 67.46 65.39 42.19 59.25
GRU [38] 70.43 74.14 74.20 59.47 69.56
Transformer [39] 70.22 64.56 65.31 51.15 62.81
PatchTST [40] 68.31 72.30 75.69 61.77 69.52
Resnet18 [10] 63.71 78.33 78.08 71.05 72.79
OFA [18] 73.64 76.81 77.44 63.33 72.81
ADN18 [41] 75.45 81.50 80.81 70.97 77.18
RadarLLM 79.26 82.31 81.51 72.54 78.91
(1). Data1 HH
(2). Data1 HV
(3). Data1 VH
(4). Data1 VV
(5). Data2 HH
(6). Data2 HV
(7). Data2 VH
(8). Data2 VV
(9). Data3 HH
(10). Data3 HV
(11). Data3 VH
(12). Data3 VV
Figure 8: Comparison of cross-entropy loss and preference-aware loss on low SCR datasets. We compare detection rates across different datasets and epochs, highlighting the improvements achieved by the preference-aware loss function.
IV-D Ablation Study
IV-D1 Ablation Study on the Backbone Choice
To validate the effectiveness of using pre-trained LLM, the ablation study is conducted in this section by removing or replacing the pre-trained LLM:
  1. 1.
    RadarLLM(0): In this variant, the pre-trained LLM is entirely removed, while all other components of the framework are kept unchanged.
  2. 2.
    RadarLLM(T): In this variant, the pre-trained LLM is replaced with a randomly initialized transformer block.
  3. 3.
    RadarLLM(R): In this variant, the network architecture and fine-tuning procedure of the original LLM are maintained, but its pre-trained parameters are replaced with random initialization.
In Table VII, we show the results of the ablation study. It can be seen that the removal or exchange of pre-trained LLM results in notable performance degradation, indicating the necessity of pretrained LLM for high detection performance.
IV-D2 Ablation Study on the Loss Function
In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets. First, to facilitate a direct and quantitative comparison between the two loss functions, we aggregate the outputs of all feature tokens by applying a softmax normalization function followed by averaging operation. The resulting probabilities are then used for target detection. In Fig. 9, we evaluate the detection rate on the test samples of different low SCR datasets at different training epochs. It can be seen that preference-aware loss achieves an improvement of around 1%-15% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets. In Table VIII, we show the detection performance of the reference model. It can be seen that the optimal detection performance of the fine-tuned model outperforms the reference model by approximately 3%-18% across various datasets. Notably, despite this remarkably significant performance gap, leveraging the reference model to score individual feature tokens still yields significant performance gains.
Although the preceding results are promising, a direct comparison between the proposed preference-aware (PA) loss and the conventional end to end cross-entropy (CE) loss remains lacking. To bridge this gap, we evaluate multiple model variants: RadarLLM(PA) vs. RadarLLM(CE), and PatchTST(PA) vs. PatchTST(CE). Models labeled with (PA) adopt the proposed token-level reweighting strategy based on learning value, while those with (CE) rely on standard cross-entropy loss. For further comparison, we include a sample-level reweighting baseline, denoted as WCE [28], where training weights are adjusted at the sample level.
As summarized in Table IX, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions. For instance, RadarLLM(PA) achieves an average gain of 9.9% over RadarLLM(CE), while PatchTST(PA) improves by 7.0% over its CE-based counterpart. Notably, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
TABLE VII: Results of ablation study on the backbone choice.
$H$ HV VH $V$ AVG
RadarLLM 82.71 83.84 83.91 75.23 81.42
RadarLLM(0) 77.51 79.38 80.35 69.97 76.80
RadarLLM(T) 78.25 79.77 81.16 72.13 77.83
RadarLLM(R) 75.19 75.44 75.81 65.15 72.90
TABLE VIII: Detection performance of reference model on three low SCR datasets.
Reference model $H$ HV VH $V$
Data1 46.68 45.21 46.75 18.48
Data2 34.87 50.75 48.41 42.51
Data3 47.39 59.82 61.85 44.72
TABLE IX: Results of ablation study on the loss function.
Data1 Data2 Data3 AVG
RadarLLM(PA) 52.06(+16.8) 54.87(+5.4) 65.81(+7.5) 57.58(+9.9)
RadarLLM(WCE) 43.51(+8.8) 53.95(+4.9) 64.55(+6.8) 54.84(+7.7)
RadarLLM(CE) 34.73 48.96 57.80 47.16
PatchTST(PA) 47.48(+8.1) 50.29(+5.9) 62.42(+7.0) 53.40(+7.0)
PatchTST(WCE) 46.68(+7.8) 48.96(+5.0) 57.80(+2.9) 51.15(+5.2)
PatchTST(CE) 38.91 43.92 54.89 45.91
IV-E Analysis of the Effect of Hyperparameters on Experimental Results
TABLE X: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarGPT with different number of patch size.)
Average DR Average IT(s)
RadarLLM(with a patch size of 32) 79.44 5.9259
RadarLLM(with a patch size of 48) 81.96 5.6561
RadarLLM(with a patch size of 64) 79.09 5.6222
We first investigate the impact of patch size on RadarLLM’s performance. The results of the ablation study are summarized in Table X. Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes. These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarLLM.
We then investigate the impact of the number of GPT2 layers on RadarLLM’s performance. Our ablation study on GPT2 layer configurations is presented in Table XI. Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones. These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the cross-modal transfer capabilities of the language model, while excess layers suffer from overfitting due to the task-irrelevant network parameters in deeper layers. (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal. Specifically, the inference speed of GPT2(8) remains approximately 91% of that of GPT2(0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
TABLE XI: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of GPT2 layers. (RadarLLM(n) represents RadarLLM with n LLM layers.))
Average DR Average IT(s) Total/Training NP(M)
RadarLLM(0) 77.41 5.3484 41.0410/2.4445
RadarLLM(2) 81.29 5.4812 55.2175/2.4506
RadarLLM(4) 81.96 5.6561 69.3932/2.4567
RadarLLM(6) 80.42 5.7331 83.5690/2.4628
RadarLLM(8) 80.01 5.8655 97.7447/2.4690
IV-F Case Study of the Effectiveness of Feature Token Importance
To further demonstrate the effectiveness of the proposed preference-aware loss function, we conduct a case study analyzing the training weight dynamics of feature tokens for two representive sequence features: Short-Time Fourier Transform Magnitude Spectrum (SMS) and Instantaneous Phase (IP) features.
The SMS feature exhibits a distinctive rise and fall pattern indicative of target presence, primarily characterized by feature tokens 26 and 27. As illustrated in Fig. 9, these tokens initially receive high attention weights during training, underscoring the model’s emphasis on learning these salient patterns. However, as training progresses, the model rapidly assimilates these features, resulting in a marked decline in their weights after approximately 10 epochs. This adaptive behavior reflects the model’s ability to dynamically reallocate attention from well-learned patterns to more complex or informative features, thus mitigating the risk of overfitting to easily recognizable signal features.
In contrast, the IP feature demonstrates a more complex structure, encompassing both moderately varying patterns associated with target echos (corresponding to tokens 0, 3, and 4) and rapidly fluctuating patterns caused by clutter-induced noise (corresponding to tokens 8, 9, and 10). These six tokens, arranged from left to right, align with the gray-lighted segments in the IP feature map illustrated in Fig. 9. During training, tokens 0, 3, and 4 consistently receive elevated attention weights, typically exceeding 1, indicating the model’s sustained emphasis on these informative yet challenging-to-learn feature tokens. In contrast, tokens 8, 9, and 10 are assigned substantially lower weights, which highlights the model’s ability to suppress excessive learning from noisy and less informative feature tokens.
(1). SMS feature
(2). SMS feature token weight dynamics
(3). IP feature
(4). IP feature token weight dynamics
Figure 9: Dynamics of feature token weights during training: (a) SMS feature token weight dynamics, (b) IP feature token weight dynamics. The graphs show how the attention weights assigned to individual feature tokens evolve over epochs, emphasizing the model’s ability to adaptively focus on informative features and suppress noisy tokens.
V Conclusion and Future Work
In this study, we propose RadarLLM, a novel framework that integrates a pre-trained LLM with a lightweight reference model for small marine target detection. Employing LoRA adaptation and preference-aware loss, we selectively optimize different feature patches according to their learning values in a lightweight form, effectively reducing model overfitting in low SCR environments. Extensive experimental results demonstrate that RadarLLM significantly outperforms state-of-the-art baselines across sufficient training data, and limited training data settings in various detection scenarios. These results highlight the considerable potential of LLMs in radar signal processing, especially for challenging small marine target detection. Looking ahead, several promising research directions emerge:
  1. 1.
    Future studies should investigate the use of more sophisticated open-source pre-trained LLMs, such as Meta’s LLaMA2 or LLaMA3 [42]. These much larger models may offer enhanced parameter transfer capabilities compared to GPT2.
  2. 2.
    Although the present approach demonstrates the effectiveness of the selected reference model, there remains considerable potential to explore alternative architectures. Future investigations could examine models such as Mamba [43] or leverage iterative training frameworks, which may further enhance the performance and training efficiency of the reference model.
  3. 3.
    To address the storage demands of using large language models (LLMs), we suggest employing model compression strategies, including quantization, pruning, and knowledge distillation. These methods aim to reduce model size while maintaining performance, facilitating the deployment of LLMs in resource-constrained environments.

### v2

@@
 Nevertheless, directly fine-tuning pre-trained LLMs on marine target detection tasks tends to suffer from pronounced overfitting, particularly in challenging low signal-to-clutter ratio (SCR) scenarios.
-This overfitting primarily stems from the model’s tendency to memorize spurious or noisy feature patterns rather than learning discriminative structures that generalize well to unseen data.
+This overfitting is mainly due to the model’s tendency to memorize noisy feature patterns rather than learning discriminative structures that generalize well to unseen data.
 To address this challenge, we introduce RadarLLM, a novel fine-tuning framework that utilizes an effective preference-aware loss.
@@
 We theoretically demonstrate the effectiveness of the evaluated learning values by transforming the problem as selecting useful feature tokens.
-Extensive experiments on real-world marine radar datasets show that 1) the proposed loss function is much better than the original one, with particularly significant gains in challenging low SCR scenarios and 2) RadarLLM consistently outperforms state-of-the-art baselines across diverse detection scenarios, with particularly notable gains under limited training data conditions.
+Extensive experiments on real-world marine radar datasets show that 1) the proposed loss function outperforms the original one, showing particularly significant improvements under challenging low SCR conditions, with an average performance gain of 9.9% and 2) RadarLLM consistently outperforms state-of-the-art baselines in diverse detection scenarios, with particularly notable gains under limited training data conditions.
 Index Terms:
@@
 For example, simple and regular patterns may rapidly reach convergence during model training, and then exhibit a tendency for overfitting, whereas other valuable patterns may require more iterations to achieve convergence.
-To address the aforementioned challenges, this paper introduces RadarLLM, an innovative framework for learning from radar echo signals.
+To address the aforementioned challenges, this paper introduces RadarLLM, an innovative framework designed for learning from radar echo signals.
 First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens.
 Second, we train a small reference model based on the validation samples and calculate the loss for each feature token of the training samples to estimate its generalization potential.
-Third, we construct a target detector based on a pre-trained LLM, then fine-tune the model based on low-rank adaptation (LoRA) [26] and a novel preference-aware loss function, which is inspired by the work in NLP [27] and CV [28].
-This loss introduces a selective training strategy by assigning training weights to different feature patches according to their learning values, which greatly improving model performance in low SCR environments.
-Finally, we retrain a classification head to enhance detection performance and control the false alarm rate.
+Then, we construct a target detector based on a pre-trained LLM and fine-tune the model using a lightweight finetuning module, along with a novel preference-aware loss function inspired by the work in NLP [26] and CV [27].
+This loss function introduces a selective training strategy by assigning different training weights to feature patches according to their online evaluated learning values, significantly improving model performance in low SCR environments.
+Finally, we retrain the classification head to enhance detection performance and control the false alarm rate.
 Here we summarize our key contributions as follows:
@@
     We propose RadarLLM, a novel framework combining a lightweight reference model and a pre-trained LLM for marine radar target detection.
-To the best of our knowledge, we are the first to show that well-known LLM can be a strong starting point for intelligent radar signal processing.
+To the best of our knowledge, we are the first to show that the well-known LLMs can be a strong starting point for intelligent radar signal processing.
   2. 2.
-    For fine-tuning of the framework, we propose a novel preference-aware loss function which greatly improves RadarLLM’s performance in a low SCR environment.
-RadarLLM can selectively optimize on useful feature tokens with high learning values while reducing the risk of overfitting.
+    For the fine-tuning of the LLMs, we propose a novel preference-aware loss function.
+This loss function can selectively optimize useful feature tokens with high learning values, significantly reducing the risk of model overfitting.
   3. 3.
     Our extensive experiments confirm the effectiveness of RadarLLM.
-It attains new state-of-the-art detection performance under sufficient training data, and constrained training data settings on popular marine radar datasets.
+It attains new state-of-the-art detection performance under sufficient training data and constrained training data settings on popular marine radar datasets.
 The remainder of this paper is structured as follows.
@@
 Wan et al.
-[14] propose a sequence-based target detection framework that leverages instantaneous phase, Doppler spectrum, and short-time Fourier transform features, in conjunction with a bidirectional long short-term memory (Bi-LSTM) network.
+[14] propose a sequence-based target detection framework that leverages instantaneous phase, Doppler spectrum, and short-time Fourier transform features, in conjunction with a bidirectional long-short-term memory (Bi-LSTM) network.
 Wang et al.
-[11] propose a target detector based on complex value U-Net (CV-UNet), which realizes clutter suppression based on the amplitude-phase characteristics of radar echoes, and then achieves target detection.
+[11] propose a target detector based on complex value U-Net (CV-UNet), which performs clutter suppression based on the amplitude-phase characteristics of radar echoes, and then achieves target detection.
 Su et al.
@@
 These methods provide compelling evidence that deep learning can significantly enhance detection performance without the need for explicit mathematical modeling.
-However, it is important to recognize certain limitations inherent in these methods.
-Current methods rely on relatively small NN models, which possess limited feature recognition capabilities and are highly sensitive to environmental variations.
+Despite their initial promise, these methods are inherently constrained by their dependence on relatively compact neural networks.
+Such models exhibit limited representational capacity, which hinders their ability to learn complex, high-dimensional feature representations.
+Consequently, they are particularly susceptible to performance degradation in the face of diverse and nonstationary environmental conditions.
+These challenges motivate us to first explore the use of LLMs for radar target detection.
 II-B Language Model Powered Signal Processing Task
@@
 In marine radar target detection under low SCR conditions, directly fine-tuning a pre-trained LLM often leads to severe overfitting issues, limiting its parameter transfer capability.
-We counteract this shortfall with a preference-aware loss through selectively training on useful feature patches.
+To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
 III Proposed Method
@@
 III-A Overview of the RadarLLM Framework
-In this study, we present RadarLLM, a LLM–empowered framework for marine radar target detection.
-The framework introduces a selective patch-level fine-tuning strategy that counteracts the severe overfitting problem caused by the low SCR environment.
-The overall pipeline of RadarLLM is illustrated in Fig.
-1 and shown in Algorithm 1.
-First, five complementary sequence features are extracted from radar echoes and partitioned into uniform patches to form a multi-patch representation.
-Next, a lightweight reference model rapidly scores each patch to quantify its generalization potential.
-Then, we fine-tune a pre-trained LLM backbone via a preference-aware loss and Low-Rank Adaptation (LoRA).
-Finally, an autoencoder-based classification head is re-trained to enhance detection performance and control the false alarm rate.
-The detailed design and rationale of each module are elaborated in the following subsections.
-Algorithm 1 RadarLLM
+In this study, we present RadarLLM, an LLM–empowered framework tailored for the marine radar target detection task. RadarLLM training protocol is illustrated in Fig. 1 and shown in Algorithm 1.
+To enable lightweight yet effective fine-tuning of LLMs for radar target detection, the proposed RadarLLM framework integrates four tightly coupled modules into a unified optimization pipeline.
+First, the sequence feature extraction and patching module (Section III-B) derives five discriminative sequence features from the multi-domain transformation of radar echo signals.
+Then a patching operation is applied to capture local semantic patterns while reducing computational complexity.
+Second, a lightweight finetuning module (Section III-C) is introduced to efficiently adapt the LLM backbone to the radar task, preserving its universal knowledge contained while minimizing parameter updates.
+Third, a preference-aware loss function (Section III-D), which utilizes a lightweight reference model, enables selective training on informative feature patches and effectively alleviates overfitting problem.
+Finally, an autoencoder-based classification head (Section III-E) is retrained to enhance the detection rate and maintain a controllable false alarm rate.
+As illustrated in Fig.
+1, the LLM-based model first employs an input embedding layer to project the feature patches into the same latent space as the pre-trained LLM.
+The embedded representations are then processed by the initial encoder layers of the pre-trained LLM to extract more discriminative features.
+Subsequently, a layer normalization operation refines these features, yielding the normalized representation $F^{\mathrm{LN}}$, which enhances consistency and stability.
+Finally, an output linear layer generates token-level detection results.
+For the structure of the reference model, a standard Transformer encoder [36] is utilized to process the input feature patches and produce token-level detection outputs, as illustrated in Fig.
+2.
+Algorithm 1 RadarLLM Training Protocol
 1:
@@
 9:Fine-tuned RadarLLM model with controllable FAR.
-III-B Stage 1: Sequence Feature Extraction
+Figure 2: Overview of the reference model.
+III-B Sequence Feature Extraction and Patching
 When the radar transmits coherent pulses toward the sea surface, it receives a sequence of echo signals from each range resolution cell.
@@
     $\text{DP}(f_{d})=\arg(\sum_{n=0}^{N-1}x(n)\exp\left(-j2\pi f_{d}nT_{s}\right)).$ (9)
-As shown in Fig. 2, the five sequence features reveal clear distinctions between target and clutter samples.
+As illustrated in Fig.
+3, the five extracted sequence features exhibit pronounced discriminability between target and clutter samples.
+To better structure the input for downstream processing, the extracted features are partitioned into non-overlapping patches.
+This patching approach offers several key advantages: it helps retain local semantic information, reduces the computational and memory burden of attention mechanisms.
+For a mini-batch of $B$ samples, the input feature matrix is constructed by concatenating the IP feature $F_{\mathrm{IP}}$, DSE feature $F_{\mathrm{DSE}}$, SMS feature $F_{\mathrm{SMS}}$, amplitude feature $F_{\mathrm{Amp}}$, and DP features $F_{\mathrm{DP}}$.
+$F=\text{Concat}(F_{\mathrm{IP}},F_{\mathrm{DSE}},F_{\mathrm{SMS}},F_{\mathrm{Amp}},F_{\mathrm{DP}}).$ (10)
+Each feature in F{F} is subsequently partitioned into non-overlapping patches of length $L$.
+If the final patch is shorter than $L$, zero padding is applied to ensure uniform length.
+The resulting $K$ patches are concatenated to construct the final input tensor ${F}^{\mathrm{P}}\in\mathbb{R}^{B\times K\times L}$, where $K=5\left\lceil\frac{N}{L}\right\rceil$.
+An illustration of this patching process is provided in Fig.
+4.
 (a) IP feature
@@
 Figure 3: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX database.
-III-C Structure of the RadarLLM Network
 Figure 4: An illustration of the patching operation.
-Initially, for a batch of $B$ samples, the input feature matrix is constructed by concatenating the IP feature $F_{\mathrm{IP}}$, DSE feature $F_{\mathrm{DSE}}$, SMS feature $F_{\mathrm{SMS}}$, amplitude feature $F_{\mathrm{Amp}}$, and DP features $F_{\mathrm{DP}}$.
-$F=\text{Concat}(F_{\mathrm{IP}},F_{\mathrm{DSE}},F_{\mathrm{SMS}},F_{\mathrm{Amp}},F_{\mathrm{DP}}).$ (10)
-Each feature in F{F} is subsequently partitioned into non-overlapping patches of length $L$.
-If the final patch is shorter than $L$, zero padding is applied to ensure uniform length.
-The resulting $K$ patches are concatenated to construct the final input tensor ${F}^{\mathrm{P}}\in\mathbb{R}^{B\times K\times L}$, where $K=5\left\lceil\frac{N}{L}\right\rceil$.
-An illustration of this patching process is provided in Fig.
-3.
-To adapt the feature patches to the input format required by the pre-trained language model, we employ a fully connected layer to project ${F}^{\mathrm{P}}$ into an embedded representation ${F}^{\mathrm{EB}}\in\mathbb{R}^{B\times K\times L^{\prime}}$.
-This is followed by the addition of a positional encoding ${E}^{\mathrm{pos}}$, which incorporates either relative or absolute positional information into the embedded features:
-${E}^{\mathrm{pos}}_{b,k,2l}=\sin(\frac{k}{10000^{2l/L^{\prime}}}),\ {E}^{\mathrm{pos}}_{b,k,2l+1}=\cos(\frac{k}{10000^{2l/L^{\prime}}}),$ (11)
-where $b$ represents the $b$-th sample, $k$ represents the position index of the token, and $l$ denotes the feature dimension index.
-The positional encoding ${E}^{\mathrm{pos}}$ is added to the embedding ${F}^{\mathrm{EB}}$ to produce a enriched embedding:
-${F}^{\mathrm{PE}}={F}^{\mathrm{EB}}+{E}^{\mathrm{pos}}.$ (12)
-The enriched embedding ${F}^{\mathrm{PE}}$ is passed through initial several layers of the GPT2 encoder for further feature extraction.
-To efficiently fine-tune the backbone model, we apply Low-Rank Adaptation (LoRA) [26] to the query ($\mathbf{W}_{Q}$) and value matrices ($\mathbf{W}_{V}$) in the multi-head attention layers and directly fine-tune the Layer Normalization (LayerNorm) parameters.
-The LoRA fine-tuning process is illustrated in Fig.
-4.
-This strategy introduces lightweight, trainable low-rank matrices to the fixed pre-trained weights, substantially reducing the computational and memory requirements compared to full parameter fine-tuning.
-Formally, the adapted model generates rich contextual feature representations as follows:
-${F}^{\mathrm{LLM}}=\operatorname{LLM}_{\mathrm{LoRA}}\left({F}^{\mathrm{PE}}\right)\in\mathbb{R}^{B\times K\times L^{\prime}},$ (13)
-where $\operatorname{LLM}_{\mathrm{LoRA}}(\cdot)$ represents the backbone network of the language model encoder enhanced with LoRA modules.
-To stabilize the training process, a Layer Normalization (LayerNorm) operation is applied to the output representation ${F}^{\mathrm{LLM}}$.
-$F^{\mathrm{LN}}=\mathrm{LayerNorm}({F}^{\mathrm{LLM}}),$ (14)
-where $\mathrm{LayerNorm}(\cdot)$ is defined element-wise for each position $k$ and feature dimension $l$ as:
-$F_{b,k,l}^{\mathrm{LN}}=\gamma_{b,l}\cdot\frac{F_{b,k,l}^{\mathrm{LLM}}-\mu_{b,l}}{\sqrt{\sigma_{b,l}^{2}+\epsilon}}+\beta_{b,l},$ (15)
-where $\mu_{b,l}$ and $\sigma_{b,l}^{2}$ is computed as:
-$\mu_{b,l}=\frac{1}{K}\sum_{k=1}^{K}F_{b,k,l}^{\mathrm{LLM}},\quad\sigma_{b,l}^{2}=\frac{1}{K}\sum_{k=1}^{K}\left(F_{b,k,l}^{\mathrm{LLM}}-\mu_{b,l}\right)^{2}.$ (16)
-Here, $\gamma_{b,l}$ and $\beta_{b,l}$ are learnable parameters, and $\epsilon$ is a small constant for numerical stability.
-$F^{\mathrm{LN}}$are then passed through two FC layers to obtain the final output:
-$F^{\mathrm{fi}}=FC(FC(F^{\mathrm{LLM}}))\in\mathbb{R}^{B\times K\times 2}.$ (17)
+III-C Lightweight Fine-tuning Module for the LLM Backbone
+Inspired by the strategy proposed in [18], we adopt a lightweight fine-tuning module that jointly updates the layer normalization parameters and integrates a Low-Rank Adaptation (LoRA) module [37].
+LoRA is a parameter-efficient adaptation technique that freezes the pre-trained model weights while injecting trainable low-rank decomposition matrices into selected layers.
+This design offers two key advantages.
+First, given the immense parameter scale of LLMs, full fine-tuning is often computationally prohibitive.
+This module mitigates this limitation by updating only compact low-rank matrices, thereby reducing trainable parameters by several orders of magnitude.
+Second, this module alleviates catastrophic forgetting of pre-trained knowledge during adaptation, as the newly acquired information is encoded within a subspace of lower rank than the original representation.
+Consequently, this module enables the finetuned model to retain its pre-trained general-purpose knowledge while efficiently adapting to downstream radar tasks.
+Concretely, we fine-tune the query and value matrices, $\mathbf{W}_{Q}$ and $\mathbf{W}_{V}$, within each attention block of the LLM backbone, as illustrated in Fig.
+5.
+LoRA introduces two trainable rank decomposition matrices, $A$ and $B$, attached to the frozen pre-trained weight matrix $\mathbf{W}$.
+The modified transformation is expressed as:
+$\mathbf{W}\leftarrow\mathbf{W}+AB^{T},$ (11)
+where $\mathbf{W}\in\mathbb{R}^{d{1}\times d{2}}$, $A\in\mathbb{R}^{d_{1}\times r}$, $B\in\mathbb{R}^{d_{2}\times r}$, and $r\ll\min(d_{1},d_{2})$.
+The matrices are initialized as $A\sim\mathcal{N}(0,\sigma^{2})$ and $B=0$.
+During fine-tuning, $\mathbf{W}$ remains frozen while $A$ and $B$ are optimized.
+Owing to the low-rank constraint ($r\ll\min(d_{1},d_{2})$), this approach dramatically reduces the number of trainable parameters compared to full-parameter fine-tuning.
 Figure 5: Illustration of the LoRA fine-tuning process.
-III-D Stage 2: Compute Loss for Each Token Based on the Reference Model
-The architecture of the reference model is illustrated in Fig.
-5.
-As described in Section III-C, the input patches are first projected into a latent Transformer space of dimension $D$ and augmented with learnable additive positional embeddings.
-This process yields the enriched representation $F^{\mathrm{PE}^{\prime}}\in\mathbb{R}^{B\times K\times D}$, which is subsequently processed by a Transformer encoder.
-Within the encoder, the multi-head attention (MHA) mechanism functions as follows: the input $F^{\mathrm{PE}^{\prime}}$ is linearly projected into $H$ independent attention heads.
-For each head $h$, query, key, and value matrices $\mathbf{W}_{Q_{h}}\in\mathbb{R}^{B\times D\times d_{k}}$ $\mathbf{W}_{K}\in\mathbb{R}^{B\times D\times d_{k}}$ $\mathbf{W}_{V}\in\mathbb{R}^{B\times D\times d_{v}}$ are used to compute:
-$\mathbf{Q}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{Q_{h}},\quad\mathbf{K}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{K_{h}},\quad\mathbf{V}_{h}=F^{\mathrm{PE}^{\prime}}\mathbf{W}_{V_{h}},$ (18)
-where $d_{k}=d_{v}=D/H$. The attention output of each head is then computed as:
-$\mathbf{O}_{h}=\operatorname{Softmax}\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}{}^{\bm{\top}}}{\sqrt{d_{k}}}\right)\mathbf{V}_{h}.$ (19)
-Outputs from all heads are concatenated along the feature dimension and projected back to dimension $D$.
-The MHA block further incorporates Batch Normalization layers and a feed-forward network (FFN) equipped with residual connections, resulting in the final encoded representation $F^{\mathrm{TF}}\in\mathbb{R}^{B\times K\times D}$.
-Subsequently, LayerNorm is applied, followed by two fully connected layers to yield the final output $F^{\mathrm{fi}^{\prime}}\in\mathbb{R}^{B\times K\times 2}$.
-Figure 5: Overview of the reference model.
-III-E Stage 3: Fine-tune RadarLLM Network Based on the Preference-aware Loss
-III-E1 Methodology
-To improve fine-tuning of LLMs for radar target detection in low SCR scenarios, we propose a token-level preference-aware loss function.
-In such conditions, many feature patches are either severely corrupted by noise or trivially easy, offering limited value for effective learning.
-Uniformly optimizing all patches risks overfitting by forcing the model to learn from uninformative or misleading inputs.
-To address this, we introduce a dynamic weighting mechanism that prioritizes tokens with higher learning value, those expected to contribute more to generalization.
-Specifically, our approach utilizes a dual-model framework consisting of a lightweight reference model $\theta_{r}$ and a high-capacity target model $\theta_{t}$, which leverages LLMs.
+III-D Preference-aware Loss Function
+III-D1 Methodology
+To enhance the fine-tuning of LLMs for radar target detection under low SCR conditions, we introduce a token-level preference-aware loss function.
+In such scenarios, a substantial proportion of feature patches contribute little information for effective learning.
+Uniform optimization over all feature patches risks overfitting by coercing the model to memorize uninformative or even misleading patterns.
+To mitigate this, we propose a dynamic re-weighting scheme that adaptively assigns higher learning weights to feature patches exhibiting superior generalization potential.
+Specifically, our approach uses a dual-model framework consisting of a lightweight reference model $\theta_{r}$ and a high-capacity target model $\theta_{t}$, which leverages LLMs.
   • Reference Model: The reference model is built with a lightweight Transformer encoder and trained solely on validation samples using standard cross-entropy loss.
@@
 In contrast, a higher reference loss typically reflects excessive noise, making the token less useful for improving detection performance.
-  • Target Model: A significantly larger model built upon a pre-trained LLM, parameterized by $\theta_{t}$, as described in Section III-C.
+  • Target Model: A significantly larger model built upon a pre-trained LLM, parameterized by $\theta_{t}$.
 For a batch of training samples, we compute the token-level cross-entropy losses for both models. Given a feature token $F_{b,k}^{\mathrm{fi}}$, the corresponding two losses are as follows:
@@
 where $B$ and $K$ denote the batch size and the number of feature tokens, respectively.
-III-E2 Theoretical Insight of the Effectiveness of Feature Token Importance
+III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
 We begin by formalizing the task of selecting informative feature patches for training.
@@
 Consequently, the practical token weight is derived as shown in Eq.(14).
-III-F Stage 4: Retrain Classification Head Based on Autoencoder Network
-As illustrated in Fig.
-1, the output of the LLM-based model has a shape of $[B,K,2]$, representing predictions across multiple tokens.
-This token-level output cannot be directly used for detection due to the absence of a unified decision representation.
-To address this, we propose retraining a classifier based on the learned representation $F^{\mathrm{LN}}$.
-Specifically, we employ an autoencoder-based classification head designed to extract the most informative features from the high-dimensional input, thereby enabling effective and robust decision-making.
-The proposed autoencoder-based classification head comprises three key components: (1) an encoder, consisting of convolutional layers followed by fully connected layers, which extracts compact and discriminative feature representations from $F^{\mathrm{LN}}$; (2) a decoder, symmetrically constructed with convolutional and fully connected layers, which reconstructs the input from the latent space to promote feature consistency and preservation; and (3) a classification head, which operates on the latent vector to perform the final classification.
-The autoencoder is trained with a dual-objective loss that combines reconstruction loss and classification loss.
-The reconstruction loss ensures that the latent representation retains critical information by minimizing the discrepancy between the original input $F^{\mathrm{LN}}$ and its reconstruction $\hat{F}^{\mathrm{LN}}$:
+III-D3 Comparison with Existing Loss Designs
+Most existing loss improvements for marine radar target detection address issues such as sample imbalance, hard sample learning, or false alarm control.
+Various solutions have been proposed, including enhanced focal loss that emphasizes learning from hard or moderately difficult samples [38], and Neyman-Pearson criterion-inspired loss that effectively regulates the false alarm rate [39].
+However, these approaches seldom address a crucial issue: mitigating model overfitting in complex real-world scenarios, particularly under low SCR conditions.
+In contrast, the proposed preference-aware loss incorporates a selective training mechanism that directs the model’s attention to transferable and informative patterns, thus substantially reducing the risk of overfitting.
+Moreover, while sample reweighting strategies are often employed to alleviate overfitting by defining a weighting function that maps the training loss to a corresponding sample weight, our method adopts a patch-level reweighting strategy motivated by the uneven distribution of informative patterns across multi-domain radar features.
+This design enables a more fine-grained and flexible learning process.
+In general, the proposed preference-aware loss marks a major and substantial advance in enhancing model generalization across diverse and low-SCR radar detection scenarios.
+III-E False Alarm Controllable Classification Head Retraining
+As shown in Fig.
+1, the LLM-based model outputs $[B,K,2]$ token-level predictions, which lack a unified detection representation.
+To achieve consistent decision-making, we retrain a classifier using the learned feature $F^{\mathrm{LN}}$.
+An autoencoder-based classification head extracts informative features from this high-dimensional input for robust detection.
+It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative features; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector.
+The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
 $\mathcal{L}_{\text{recon }}=\left\|F^{\mathrm{LN}}-\hat{F}^{\mathrm{LN}}\right\|_{2}^{2}.$ (22)
-The classification loss ensures accurate predictions by aligning the predicted probability distribution $\hat{y}_{b}^{i}$ with the ground-truth label $y_{b}^{i}$.
 $\mathcal{L}_{\text{ce}}=-\frac{1}{B}\sum_{b=1}^{B}\sum_{i=1}^{2}y_{b}^{i}\log\hat{y}_{b}^{i}.$ (23)
-To jointly optimize both the reconstruction and classification objectives, we adopt a task uncertainty weighting strategy [37], which adaptively adjusts the contribution of each loss term based on its inherent uncertainty.
-The total loss function is expressed as:
+A task-uncertainty weighting scheme [40] adaptively balances these objectives:
 […]
-where $\sigma_{\text{recon}}$ and $\sigma_{\text{ce}}$ are learnable parameters representing the task-dependent uncertainties for the reconstruction and classification losses, respectively.
-This formulation allows the model to balance the learning dynamics of both tasks in a principled and data-driven manner.
-To control the false alarm rate during inference, we sort the first element in the softmax output for all test clutter samples in descending order, yielding a one-dimensional array $O_{\mathrm{sorted}}$.
+where $\sigma_{\text{recon}}$ and $\sigma_{\text{ce}}$ are learnable task uncertainties that regulate the joint optimization dynamics.
+During inference, false alarms are controlled via threshold-setting using sorted softmax outputs of clutter samples.
 The detection threshold $\eta$ is then determined based on the desired false alarm rate $P_{fa}^{\text{d}}$ as follows:
 $\begin{array}[]{l}\eta=O_{\mathrm{sorted}}(x),\ x=\lceil P_{fa}^{\text{d}}\times N_{\text{clutter}}\rceil,\end{array}$ (25)
-where $O_{\mathrm{sorted}}(x)$ denotes the $x$-th largest output in the sorted array, $N_{\text{clutter}}$ is the total number of test clutter samples, and $\lceil\cdot\rceil$ is the ceiling operation to ensure integer indexing.
-Figure 6: Overview of the autoencoder model.
+where $O_{\mathrm{sorted}}(x)$ is the $x$-th largest output and $N_{\text{clutter}}$ is the total number of test clutter samples.
 IV Experiments
-In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarLLM’s detection performance under three settings: sufficient training data, constrained training data, and zero-shot scenarios.
+In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarLLM’s detection performance under two settings: sufficient training data and constrained training data scenarios.
 IV-A Experimental Setup
@@
 We employ eight benchmark data sequences (Data1–Data8) from the Intelligent PIxel Processing X-band (IPIX) radar archive, summarized in Table I.
-All datasets were recorded in 1993 by the IPIX radar on Canada’s east coast and include full polarization measurements (HH, HV, VH, and VV).
+All datasets were recorded in 1993 by the IPIX radar on the east coast of Canada and include full polarization measurements (HH, HV, VH, and VV).
 Each set contains 14 range cells sampled at 1 kHz, with 131 072 complex echoes per cell, yielding an observation window of 0.512 s per sample.
 Returns from the primary range cell constitute target echoes, whereas returns from the remaining clutter-only cells represent sea clutter.
-To facilitate performance analysis across varying operating conditions, Data1–Data3 correspond to low signal-to-clutter ratio (SCR) scenes, while Data4–Data8 represent high-SCR scenes.
+To facilitate performance analysis in varying operating conditions, Data1–Data3 correspond to low signal-to-clutter ratio (SCR) scenes, while Data4–Data8 represent high-SCR scenes.
 To ensure sufficient training data, we employ overlapped segmentation following the partition rule in Eq.
@@
 This process generates 4,079 target samples and more than 9,000 clutter samples per dataset.
-The samples are divided into three groups: (1) a training set using the first 20% of observation time for both the target and clutter cells, (2) a validation set covering 20% to 35% of the observation time, and (3) a test set containing the remaining samples.
+The samples are divided into three groups: (1) a training set using the first 20% observation time for both the target and clutter cells, (2) a validation set covering 20% to 35% of the observation time, and (3) a test set containing the remaining samples.
 TABLE I: Overview of the IPIX Database.
@@
   • RNN [14]: RNN is a sequential neural architecture with temporal feedback loops. In experiments, the number of RNN layers and the hidden layer nodes are set to 2 and 6, respectively.
-  • Bi-LSTM [14]: Bi-LSTM enhances the traditional Long Short-Term Memory (LSTM) by incorporating bidirectional processing.
+  • Bi-LSTM [14]: Bi-LSTM enhances traditional long-short-term memory (LSTM) by incorporating bidirectional processing.
 In our implementation, the number of LSTM layers and the number of hidden units per layer are set to 2 and 6, respectively.
   • GRU [41]: GRU employs simplified gating mechanisms compared to LSTM. In experiments, the number of GRU layers and the hidden layer nodes are set to 2 and 8, respectively.
-  • Transformer [39]: Transformer evolutionizes sequence processing through self-attention mechanisms.
+  • Transformer [36]: Transformer reshapes sequence processing through self-attention mechanisms.
 In experiments, the Transformer encoder is configured with a model dimensionality of 32, comprising 2 layers and 8 attention heads.
@@
 In maritime target detection, achieving a high detection rate at an extremely low false alarm rate is imperative, as misinterpreting sea clutter as targets can have significant consequences, particularly in military operations.
-We evaluate the detection performance of different detectors at a false alarm rate of 0.005, considering our test set only comprises 6,000 clutter samples.
+We evaluated the detection performance of different detectors at a false alarm rate of 0.005, considering that our test set only comprises 6,000 clutter samples.
 Selecting a lower false alarm rate could result in unstable comparisons due to the limited number of clutter samples.
@@
 We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone.
-Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in average detection rate.
+Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in the average detection rate.
 Furthermore, we compare RadarLLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment.
@@
 Although RadarLLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning.
-Notably, RadarLLM achieves an exceptionally low average inference latency.
-Specifically, it requires only 5.6561 seconds to process over 9,000 test samples, making it 28.44 times faster than the ADN18 method.
+In particular, RadarLLM achieves an exceptionally low average inference latency.
+Specifically, it requires only 5.6561 seconds to process more than 9,000 test samples, which is 28.44 times faster than the ADN18 method.
 This efficiency is primarily attributed to our patching strategy, which substantially reduces the number of feature tokens, and consequently, the computational overhead.
@@
 Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge.
-Considering this issue, we conduct experiments under constrained training sample conditions to assess the effectiveness of RadarLLM.
+Considering this issue, we conducted experiments under constrained training sample conditions to assess the effectiveness of RadarLLM.
 The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
@@
   • Second, the substantial parameter scale of RadarLLM inherently grants robustness to distributional variations, enabling superior generalization compared to smaller-scale neural networks.
-Its extensive capacity allows it to capture complex data structures and efficiently handle variability across different radar detection scenarios.
-  • Third, the integration of a preference-aware loss function notably improves RadarLLM’s detection performance, especially in challenging scenarios characterized by low signal-to-clutter ratios (SCR).
+Its extensive capacity allows it to capture complex data structures and efficiently handle variability in different radar detection scenarios.
+  • Third, the integration of a preference-aware loss function significantly improves RadarLLM detection performance, especially in challenging scenarios characterized by low signal-to-clutter ratios (SCR).
 TABLE VI: Detection performance under limited training samples of different methods when the false alarm rate is 0.005.
@@
 IV-D Ablation Study
-IV-D1 Ablation Study on the Backbone Choice
-To validate the effectiveness of using pre-trained LLM, the ablation study is conducted in this section by removing or replacing the pre-trained LLM:
+IV-D1 Ablation Study on Backbone Architectures and Fine-tuning Strategies
+To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the pre-trained LLM:
   1. 1.
@@
     RadarLLM (R): In this variant, the network architecture and fine-tuning procedure of the original LLM are maintained, but its pre-trained parameters are replaced with random initialization.
-In Table VII, we show the results of the ablation study.
+In Table VII, we show the results of the ablation study on backbone choice.
 It can be seen that the removal or exchange of pre-trained LLM results in notable performance degradation, indicating the necessity of pretrained LLM for high detection performance.
-IV-D2 Ablation Study on the Loss Function
-In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets.
-First, to facilitate a direct and quantitative comparison between the two loss functions, we aggregate the outputs of all feature tokens by applying a softmax normalization function followed by averaging operation.
-The resulting probabilities are then used for target detection.
-In Fig.
-9, we evaluate the detection rate on the test samples of different low SCR datasets at different training epochs.
-It can be seen that preference-aware loss achieves an improvement of around 1%-15% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets.
-In Table VIII, we show the detection performance of the reference model.
-It can be seen that the optimal detection performance of the fine-tuned model outperforms the reference model by approximately 3%-18% across various datasets.
-Notably, despite this remarkably significant performance gap, leveraging the reference model to score individual feature tokens still yields significant performance gains.
-Although the preceding results are promising, a direct comparison between the proposed preference-aware (PA) loss and the conventional end to end cross-entropy (CE) loss remains lacking.
-To bridge this gap, we evaluate multiple model variants: RadarLLM(PA) vs.
-RadarLLM(CE), and PatchTST(PA) vs.
-PatchTST(CE).
-Models labeled with (PA) adopt the proposed token-level reweighting strategy based on learning value, while those with (CE) rely on standard cross-entropy loss.
-For further comparison, we include a sample-level reweighting baseline, denoted as WCE [28], where training weights are adjusted at the sample level.
-As summarized in Table IX, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions.
-For instance, RadarLLM(PA) achieves an average gain of 9.9% over RadarLLM(CE), while PatchTST(PA) improves by 7.0% over its CE-based counterpart.
-Notably, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
 TABLE VII: Results of ablation study on the backbone choice.
 $H$ HV VH $V$ AVG
-RadarLLM 82.71 83.84 83.91 75.23 81.42
+RadarLLM 82.51 84.81 84.61 75.90 81.96
 RadarLLM (0) 77.51 79.38 80.35 69.97 76.80
@@
 RadarLLM (R) 75.19 75.44 75.81 65.15 72.90
-TABLE VIII: Detection performance of reference model on three low SCR datasets.
+Moreover, to verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study as summarized in Table VIII, and the corresponding results are presented in Table IX.
+The results demonstrate that jointly fine-tuning the LoRA modules and Layer Normalization parameters achieves the highest overall performance.
+In contrast, fine-tuning only the Layer Normalization parameters, the LoRA modules, or performing full fine-tuning results in moderate performance degradation.
+These findings indicate that the proposed lightweight fine-tuning strategy effectively balances model adaptability and computational efficiency while preserving the universal knowledge embedded in the LLM backbone.
+TABLE VIII: Ablation Study on Fine-tuning Strategies settings.
+LoRA Layer Norm Other Params
+RadarLLM $\checkmark$ $\checkmark$
+RadarLLM (F) $\checkmark$ $\checkmark$
+RadarLLM (LoRA) $\checkmark$
+RadarLLM (LN) $\checkmark$
+TABLE IX: Results of ablation study on the finetuning strategy.
+$H$ HV VH $V$ AVG
+RadarLLM 82.51 84.81 84.61 75.90 81.96
+RadarLLM (F) 81.69 83.45 84.41 74.07 80.91
+RadarLLM (LoRA) 82.20 84.45 84.56 75.53 81.69
+RadarLLM (LN) 82.33 83.73 83.80 75.12 81.25
+IV-D2 Ablation Study on the Loss Function
+In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets.
+First, to facilitate a direct and quantitative comparison between the two loss functions, we aggregated the outputs of all feature tokens by applying a softmax normalization function followed by averaging operation.
+The resulting probabilities are then used for target detection.
+In Fig.
+7, we evaluate the detection rate on the test samples of different low SCR datasets at different training epochs.
+It can be seen that preference-aware loss achieves an improvement of around 1%-15% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets.
+In Table X, we show the detection performance of the reference model.
+It can be seen that the optimal detection performance of the fine-tuned model outperforms the reference model by approximately 3%-18% across various datasets.
+In particular, despite this remarkably significant performance gap, leveraging the reference model to score individual feature tokens still yields significant performance gains.
+Although the preceding results are promising, a direct comparison between the proposed preference-aware (PA) loss and the conventional end-to-end cross-entropy (CE) loss remains lacking.
+To bridge this gap, we evaluated multiple model variants: RadarLLM (PA) vs.
+RadarLLM (CE), and PatchTST (PA) vs.
+PatchTST (CE).
+Models labeled with (PA) adopt the proposed token-level reweighting strategy based on learning value, whereas those labeled with (CE) rely on standard cross-entropy loss.
+For further comparison, we include a sample-level reweighting baseline, denoted as WCE [27], where training weights are adjusted at the sample level.
+As summarized in Table XI, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions.
+For example, RadarLLM (PA) achieves an average gain of 9.9% over RadarLLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart.
+In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
+TABLE X: Detection performance of reference model on three low SCR datasets.
 Reference model $H$ HV VH $V$
@@
 Data3 47.39 59.82 61.85 44.72
-TABLE IX: Results of ablation study on the loss function.
+TABLE XI: Results of ablation study on the loss function.
 Data1 Data2 Data3 AVG
@@
 IV-E Analysis of the Effect of Hyperparameters on Experimental Results
-TABLE X: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarGPT with different number of patch size.)
+TABLE XII: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of patch size.)
 Average DR Average IT(s)
@@
 We first investigate the impact of patch size on RadarLLM’s performance.
-The results of the ablation study are summarized in Table X.
+The results of the ablation study are summarized in Table XII.
 Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes.
 These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarLLM.
-We then investigate the impact of the number of GPT2 layers on RadarLLM’s performance.
-Our ablation study on GPT2 layer configurations is presented in Table XI.
+We then investigated the impact of the number of GPT2 layers on RadarLLM’s performance.
+Our ablation study on GPT2 layer configurations is presented in Table XIII.
 Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones.
@@
 (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal.
-Specifically, the inference speed of GPT2(8) remains approximately 91% of that of GPT2(0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
-TABLE XI: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of GPT2 layers.
+Specifically, the inference speed of GPT2 (8) remains approximately 91% that of GPT2 (0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
+TABLE XIII: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of GPT2 layers.
 (RadarLLM (n) represents RadarLLM with n LLM layers.))
@@
 To further demonstrate the effectiveness of the proposed preference-aware loss function, we conduct a case study analyzing the training weight dynamics of feature tokens for two representive sequence features: Short-Time Fourier Transform Magnitude Spectrum (SMS) and Instantaneous Phase (IP) features.
-The SMS feature exhibits a distinctive rise and fall pattern indicative of target presence, primarily characterized by feature tokens 26 and 27.
+The SMS feature exhibits a distinctive rise and fall pattern indicative of target presence, mainly characterized by the feature tokens 26 and 27.
 As illustrated in Fig.
-9, these tokens initially receive high attention weights during training, underscoring the model’s emphasis on learning these salient patterns.
+7, these tokens initially receive high attention weights during training, underscoring the model’s emphasis on learning these salient patterns.
 However, as training progresses, the model rapidly assimilates these features, resulting in a marked decline in their weights after approximately 10 epochs.
@@
 These six tokens, arranged from left to right, align with the gray-lighted segments in the IP feature map illustrated in Fig.
-9.
+8.
 During training, tokens 0, 3, and 4 consistently receive elevated attention weights, typically exceeding 1, indicating the model’s sustained emphasis on these informative yet challenging-to-learn feature tokens.
@@
 In this study, we propose RadarLLM, a novel framework that integrates a pre-trained LLM with a lightweight reference model for small marine target detection.
-Employing LoRA adaptation and preference-aware loss, we selectively optimize different feature patches according to their learning values in a lightweight form, effectively reducing model overfitting in low SCR environments.
-Extensive experimental results demonstrate that RadarLLM significantly outperforms state-of-the-art baselines across sufficient training data, and limited training data settings in various detection scenarios.
+Employing LoRA adaptation and preference-aware loss, we selectively optimize different feature patches according to their online evaluated learning values in a lightweight form, effectively reducing model overfitting issues in low SCR environments.
+Extensive experimental results demonstrate that RadarLLM significantly outperforms state-of-the-art baselines across sufficient training data and limited training data settings in various detection scenarios.
 These results highlight the considerable potential of LLMs in radar signal processing, especially for challenging small marine target detection.
@@
   1. 1.
-    Future studies should investigate the use of more sophisticated open-source pre-trained LLMs, such as Meta’s LLaMA2 or LLaMA3 [42].
+    Future studies should investigate the use of more sophisticated open source pre-trained LLMs, such as LLaMA2 or LLaMA3 [44].
 These much larger models may offer enhanced parameter transfer capabilities compared to GPT2.

### v3

@@
-RadarLLM: Adapting Pretrained Large Language Models for Marine Radar Target Detection with Preference-aware Loss
+RadarPLM: Adapting Pretrained Language Models for Marine Radar Target Detection with Preference-aware Loss
 Abstract
-Recent advances in pre-trained large language models (LLMs) have demonstrated their capacities to capture universal knowledge, making them promising general-purpose optimization solvers for wireless signal processing.
-Motivated by these findings, we take the first step towards fine-tuning pre-trained LLMs for the effective analysis of radar signal features in marine target detection tasks.
-Nevertheless, directly fine-tuning pre-trained LLMs on marine target detection tasks tends to suffer from pronounced overfitting, particularly in challenging low signal-to-clutter ratio (SCR) scenarios.
-This overfitting is mainly due to the model’s tendency to memorize noisy feature patterns rather than learning discriminative structures that generalize well to unseen data.
-To address this challenge, we introduce RadarLLM, a novel fine-tuning framework that utilizes an effective preference-aware loss.
-Unlike conventional training strategies that uniformly optimize all feature tokens, this loss function selectively optimizes different feature patches based on their online evaluated learning values, thus guiding the model to focus on the most generalizable patterns during optimization.
-We theoretically demonstrate the effectiveness of the evaluated learning values by transforming the problem as selecting useful feature tokens.
-Extensive experiments on real-world marine radar datasets show that 1) the proposed loss function outperforms the original one, showing particularly significant improvements under challenging low SCR conditions, with an average performance gain of 9.9% and 2) RadarLLM consistently outperforms state-of-the-art baselines in diverse detection scenarios, with particularly notable gains under limited training data conditions.
+Recent advances in pre-trained language models (PLMs) have demonstrated their capabilities in capturing universal knowledge, making them promising applications for radar signal processing.
+Nevertheless, directly fine-tuning PLMs on radar signals is both computationally expensive and prone to overfitting, particularly in low signal-to-clutter ratio (SCR) environments.
+In this paper, we propose a novel fine-tuning framework for PLM-based marine radar target detection.
+First, we design a lightweight adaptation module, enabling parameter-efficient fine-tuning while preserving the pretrained model’s general knowledge.
+Second, a novel preference-aware loss is developed to selectively optimize different feature patches based on their online evaluated learning values, guiding the model to concentrate on the most generalizable feature patterns during optimization.
+Extensive experiments on real-world marine radar datasets demonstrate that the proposed finetuning framework achieves an average performance improvement of 9.9% over the standard approach under low SCR conditions.
+Furthermore, the fine-tuned model, RadarPLM, consistently outperforms state-of-the-art detectors, particularly when training data are limited.
 Index Terms:
-Marine small target detection, radar signal processing, pre-trained large language models (LLMs), preference-aware loss.
+Marine small target detection, radar signal processing, pre-trained language models (PLMs), preference-aware loss.
 I Introduction
@@
 These limitations lead to poor detection performance, particularly in scenarios with scarce training data or with significant environmental variability.
-Extensive recent studies [18, 19, 20, 21, 22, 23, 24] have shown that pre-trained LLM can be a good starting point for wireless signal and time series modeling, as they are pre-trained on massive sequential datasets.
+Extensive recent studies [18, 19, 20, 21, 22, 23, 24] have shown that pre-trained language models (PLMs) can be good starting point for wireless signal and time series modeling, as they are pre-trained on massive sequential datasets.
 Compared to random initialized weights of small models, this starting point may significantly improve model optimization stability and effectiveness.
-Moreover, the self-attention mechanism in LLM exhibits structural similarity to Principal Component Analysis (PCA) [18], demonstrating strong generality in extracting key features from high-dimensional sequence data.
-This insight opens a very promising direction for radar target detection by fine-tuning pre-trained LLMs, offering the potential for performance improvements.
-Our preliminary study [25] shows that fine-tuning a pre-trained GPT2 model to analyze sequence features yields state-of-the-art detection performance.
-Extensive ablation experiments further confirm that removing the GPT2 model or substituting it with a randomly initialized Transformer encoder significantly degrades the detection rate.
+Moreover, the self-attention mechanism in PLM exhibits structural similarity to Principal Component Analysis (PCA) [18], demonstrating strong generality in extracting key features from high-dimensional sequence data.
+This insight opens a very promising direction for radar target detection by fine-tuning PLMs, offering the potential for performance improvements.
 Although powerful and promising, some challenges still exist.
-First, the large scale of LLM requires lightweight and efficient fine-tuning strategies, which are essential for practical deployment, especially in resource-constrained settings.
-Second, and more critically, for radar applications, adapting LLMs to process sequence features in low signal-to-clutter ratio (SCR) conditions is far from a routine task, often leading to severe overfitting issues.
+First, the large scale of PLM requires lightweight and efficient fine-tuning strategies, which are essential for practical deployment, especially in resource-constrained settings.
+Second, for practical radar applications, adapting PLMs to process sequence features in low signal-to-clutter ratio (SCR) conditions often leads to severe overfitting issues.
 This phenomenon is fundamentally attributed to the intrinsic characteristics of sequence features:
-  • Significant Clutter-Induced Noise: In low SCR environments, small targets are frequently obscured by ocean waves.
-This causes large noise in sequence features, especially in phase and amplitude features, increasing the likelihood of the model overfitting to irrelevant and noisy information.
+  • Significant Clutter-Induced Noise: In low SCR environments, small targets are frequently obscured by ocean waves, which inevitably generate mixed-statue segments.
+The mixed statue segments can cause a low SCR and increase the likelihood of the model overfitting to irrelevant and noisy information.
   • Feature Pattern Convergence Imbalance: Various patterns in different sequence features often exhibit different convergence rates during fine-tuning.
-For example, simple and regular patterns may rapidly reach convergence during model training, and then exhibit a tendency for overfitting, whereas other valuable patterns may require more iterations to achieve convergence.
-To address the aforementioned challenges, this paper introduces RadarLLM, an innovative framework designed for learning from radar echo signals.
+For example, simple and regular patterns may rapidly reach convergence during model fine-tuning and then exhibit a tendency for overfitting, whereas other valuable patterns may require more iterations to achieve convergence.
+Consequently, uniform optimization across all feature patterns may cause the model to overfit not only to noisy patterns but also to overly simplistic ones, which fail to capture the more generalizable patterns inherent in radar sequence features.
+To overcome the above-mentioned issues, including the demand for lightweight fine-tuning and the tendency of overfitting under low SCR conditions, we develop a novel framework named RadarPLM.
+The proposed framework effectively bridges the gap between PLMs and radar target detection tasks through lightweight adaptation module and selective training.
 First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens.
-Second, we train a small reference model based on the validation samples and calculate the loss for each feature token of the training samples to estimate its generalization potential.
-Then, we construct a target detector based on a pre-trained LLM and fine-tune the model using a lightweight finetuning module, along with a novel preference-aware loss function inspired by the work in NLP [26] and CV [27].
-This loss function introduces a selective training strategy by assigning different training weights to feature patches according to their online evaluated learning values, significantly improving model performance in low SCR environments.
-Finally, we retrain the classification head to enhance detection performance and control the false alarm rate.
+Next, a radar target detector is constructed based on a PLM, which is fine-tuned through a lightweight adaptation module to efficiently align with the radar detection task.
+For model optimization, we design a preference-aware loss function by incorporating a lightweight reference model, which is inspired by recent advances [25, 26].
+This loss introduces a selective training strategy that dynamically adjusts the learning weights of different feature patches according to their online-evaluated learning values, significantly mitigating model overfitting.
+Finally, the binary classification head is retrained to further improve detection performance.
+Our preliminary study [27] demonstrated that fine-tuning a pre-trained GPT2 model effectively extracts discriminative information from sequence features, achieving state-of-the-art detection performance.
+However, model overfitting remains a challenge.
+To address this, we propose the RadarLLM framework, which introduces a lightweight adaptation module and selective training strategy for more reliable and generalizable adaptation.
 Here we summarize our key contributions as follows:
   1. 1.
-    We propose RadarLLM, a novel framework combining a lightweight reference model and a pre-trained LLM for marine radar target detection.
+    We propose RadarPLM, a novel framework combining a lightweight reference model and a fine-tuned PLM for marine radar target detection.
 To the best of our knowledge, we are the first to show that the well-known LLMs can be a strong starting point for intelligent radar signal processing.
   2. 2.
-    For the fine-tuning of the LLMs, we propose a novel preference-aware loss function.
-This loss function can selectively optimize useful feature tokens with high learning values, significantly reducing the risk of model overfitting.
+    For effective fine-tuning, we develop a lightweight adaptation module together with a novel preference-aware loss function, which jointly reduce computational overhead and effectively mitigate the risk of model overfitting.
   3. 3.
-    Our extensive experiments confirm the effectiveness of RadarLLM.
-It attains new state-of-the-art detection performance under sufficient training data and constrained training data settings on popular marine radar datasets.
+    Our extensive experiments validate that RadarLLM achieves new state-of-the-art detection performance on popular marine radar datasets under both sufficient and limited training data settings.
 The remainder of this paper is structured as follows.
 Section II briefly summarizes the related work.
-Section III presents the detailed operating procedure of the proposed RadarLLM.
+Section III presents the detailed operating procedure of the proposed RadarPLM.
 Section IV introduces the experimental results and analysis.
@@
 Consequently, they are particularly susceptible to performance degradation in the face of diverse and nonstationary environmental conditions.
-These challenges motivate us to first explore the use of LLMs for radar target detection.
+These challenges motivate us to first explore the use of PLMs for radar target detection.
 II-B Language Model Powered Signal Processing Task
@@
 [32] utilize the GPT2 model to develop an effective and robust beam prediction method.
-Zheng and Dai [21] propose a multi-task LLM framework to satisfy the requirements of different wireless communication tasks.
+Zheng and Dai [21] propose a multi-task PLM framework to satisfy the requirements of different wireless communication tasks.
 IOT-LLM and Penetrative AI [33, 34] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks.
-However, recent studies [35] have raised concerns about the universal applicability of LLM in signal processing tasks.
-LLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks.
-In marine radar target detection under low SCR conditions, directly fine-tuning a pre-trained LLM often leads to severe overfitting issues, limiting its parameter transfer capability.
+However, recent studies [35] have raised concerns about the universal applicability of PLM in signal processing tasks.
+PLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks.
+In marine radar target detection under low SCR conditions, directly fine-tuning a PLM often leads to severe overfitting issues, limiting its parameter transfer capability.
 To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
 III Proposed Method
-Figure 1: Overview of our RadarLLM framework.
-III-A Overview of the RadarLLM Framework
-In this study, we present RadarLLM, an LLM–empowered framework tailored for the marine radar target detection task. RadarLLM training protocol is illustrated in Fig. 1 and shown in Algorithm 1.
-To enable lightweight yet effective fine-tuning of LLMs for radar target detection, the proposed RadarLLM framework integrates four tightly coupled modules into a unified optimization pipeline.
+Figure 1: Overview of our RadarPLM framework.
+Figure 2: Overview of the reference model.
+Figure 3: An illustration of the patching operation.
+III-A Overview of the RadarPLM Framework
+The overall framework of RadarPLM is illustrated in Fig.
+1, which consists of four tightly connected stages: (1) sequence feature extraction and patching, (2) reference model training with token-level loss computation, (3) fine-tune LLMs for marine radar target detection, and (4) autoencoder-based binary classification head retraining.
+To enable lightweight and generalizable fine-tuning of PLMs for radar target detection, the proposed RadarPLM framework integrates four tightly coupled modules into a unified optimization pipeline.
 First, the sequence feature extraction and patching module (Section III-B) derives five discriminative sequence features from the multi-domain transformation of radar echo signals.
-Then a patching operation is applied to capture local semantic patterns while reducing computational complexity.
-Second, a lightweight finetuning module (Section III-C) is introduced to efficiently adapt the LLM backbone to the radar task, preserving its universal knowledge contained while minimizing parameter updates.
-Third, a preference-aware loss function (Section III-D), which utilizes a lightweight reference model, enables selective training on informative feature patches and effectively alleviates overfitting problem.
-Finally, an autoencoder-based classification head (Section III-E) is retrained to enhance the detection rate and maintain a controllable false alarm rate.
-As illustrated in Fig.
-1, the LLM-based model first employs an input embedding layer to project the feature patches into the same latent space as the pre-trained LLM.
-The embedded representations are then processed by the initial encoder layers of the pre-trained LLM to extract more discriminative features.
-Subsequently, a layer normalization operation refines these features, yielding the normalized representation $F^{\mathrm{LN}}$, which enhances consistency and stability.
-Finally, an output linear layer generates token-level detection results.
-For the structure of the reference model, a standard Transformer encoder [36] is utilized to process the input feature patches and produce token-level detection outputs, as illustrated in Fig.
+Then a patching module is devised to capture local semantic patterns while reducing computational complexity.
+Second, a lightweight adaptation module (Section III-C) is introduced to efficiently adapt the PLM backbone to the radar task, retaining its universal knowledge while minimizing parameter updates.
+Third, a preference-aware loss function (Section III-D), leveraging a lightweight reference model, enables selective optimization and mitigates overfitting.
+As shown in Fig.
+1, the PLM-based model utilizes an input embedding layer to project feature patches into the PLM’s encoder feature space.
+The embedded representations are processed by the initial encoder layers to extract more discriminative features.
+A layer normalization operation refines these features, producing the normalized representation $F^{\mathrm{LN}}$, which enhances consistency and stability.
+Finally, two output linear layers generate token-level detection results.
+For the reference model structure, a standard Transformer encoder [36] processes the input feature patches to produce token-level detection outputs, as illustrated in Fig.
 2.
-Algorithm 1 RadarLLM Training Protocol
-1:
-2:Radar echo signals $\{x(n)\}_{n=1}^{N}$ and Target labels $\{y_{b}\}_{b=1}^{B}$
-3:Pre-trained LLM parameters $\theta_{\text{LLM}}$
-4:Stage 1: Sequence Feature Extraction
-  1. 1.
-    Extract IP, DSE, SMS, Amp, and DP features from the observation vector, respectively, see Eq.(3), (4), (5),(7), and (9).
-  2. 2.
-    Patch the five features into feature tokens, see Fig. 4.
-5:Stage 2: Reference Model Training
-  1. 1.
-    Train a lightweight reference model $\theta_{r}$ based on validation samples and CE loss.
-  2. 2.
-    Compute the loss for each feature token of the training samples based on the predicted output probabilities of reference model, see Eq.(12).
-6:Stage 3: Fine-tune LLM for Target Detection
-  1. 1.
-    Compute the loss for each feature token of the training samples based on the predicted output probabilities of current model, see Eq.(13).
-  2. 2.
-    Fine-tune LLM for target detection via preference-aware loss function, see Eq.(14) and Eq.(16).
-7:Stage 4: Classification Head Retraining
-  1. 1.
-    Train autoencoder classification head based on the weighted sum of reconstruction loss and CE loss, see Eq.(22), Eq.(23), and Eq.(24).
-  2. 2.
-    Adjust detection threshold $\eta$ for controllable FAR, see Eq.(25).
-8:
-9:Fine-tuned RadarLLM model with controllable FAR.
-Figure 2: Overview of the reference model.
+An autoencoder-based binary classification head (Section III-E) is retrained to further improve detection rate.
 III-B Sequence Feature Extraction and Patching
@@
 As illustrated in Fig.
-3, the five extracted sequence features exhibit pronounced discriminability between target and clutter samples.
+4, the five extracted sequence features exhibit pronounced discriminability between target and clutter samples.
 To better structure the input for downstream processing, the extracted features are partitioned into non-overlapping patches.
@@
 An illustration of this patching process is provided in Fig.
-4.
+3.
 (a) IP feature
@@
 Figure 4: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX database.
-Figure 4: An illustration of the patching operation.
-III-C Lightweight Fine-tuning Module for the LLM Backbone
-Inspired by the strategy proposed in [18], we adopt a lightweight fine-tuning module that jointly updates the layer normalization parameters and integrates a Low-Rank Adaptation (LoRA) module [37].
-LoRA is a parameter-efficient adaptation technique that freezes the pre-trained model weights while injecting trainable low-rank decomposition matrices into selected layers.
+III-C Lightweight Fine-tuning Module for the PLM Backbone
+Adapting large-scale PLMs to specialized downstream applications like radar target detection presents significant challenges, necessitating a lightweight fine-tuning module.
+This necessity stems from two primary considerations.
+First, the immense parameter scale of PLMs renders full fine-tuning computationally expensive and often impractical for resource-constrained environments.
+Second, and more critically for radar target detection, available datasets are typically limited in size.
+Fine-tuning an entire model on such small datasets introduces the risks of catastrophic forgetting, where the model’s valuable pre-trained knowledge is overwritten, and overfitting, due to data scarcity.
+Drawing inspiration from recent work [18, 37], we propose a lightweight fine-tuning module designed to balance this trade-off between adaptability and efficiency.
+Our approach is guided by two hypotheses, consistent with [18, 37]: (1) achieving strong performance does not require updating all network parameters; instead, updating a small subset, such as bias terms or specific layers, while keeping the majority frozen is sufficient.
+(2) The weight updates during adaptation possess a low ”intrinsic rank.” Consequently, instead of modifying the entire weight matrix, we can capture task-specific knowledge within a much smaller, low-rank subspace.
+Our module implements these insights by jointly updating the layer normalization parameters and incorporating a Low-Rank Adaptation (LoRA) component.
+Concretely, we apply LoRA to the query and value projection matrices, $\mathbf{W}_{Q}$ and $\mathbf{W}_{V}$, within each attention block of the PLM backbone, as illustrated in Fig.
+5.
+For a given frozen weight matrix $\mathbf{W}$, LoRA introduces two low-rank trainable matrices, $A$ and $B$, and modifies the transformation as follows:
+$\mathbf{W}\leftarrow\mathbf{W}+AB^{T},$ (11)
+where $\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$, $A\in\mathbb{R}^{d_{1}\times r}$, $B\in\mathbb{R}^{d_{2}\times r}$, and $r\ll\min(d_{1},d_{2})$.
+The low-rank decomposition effectively projects the task-specific adaptation into a compact subspace, greatly reducing the parameter footprint while retaining original pre-trained knowledge.
+The matrices are initialized as $A\sim\mathcal{N}(0,\sigma^{2})$ and $B=0$, ensuring stable convergence and preventing interference with pre-trained weights during the early training phase.
+During fine-tuning, only $A$ and $B$ are updated, while $\mathbf{W}$ remains fixed.
+Owing to the low-rank constraint: $r\ll\min(d_{1},d_{2})$, the number of parameters for fine-tuning $A$ and $B$, i.e., $rd_{1}+rd_{2}$, is significantly less than that of the full weight matrix, $d_{1}d_{2}$.
+The reduced parameter size thus makes the fine-tuning much efficient.
 This design offers two key advantages.
-First, given the immense parameter scale of LLMs, full fine-tuning is often computationally prohibitive.
-This module mitigates this limitation by updating only compact low-rank matrices, thereby reducing trainable parameters by several orders of magnitude.
-Second, this module alleviates catastrophic forgetting of pre-trained knowledge during adaptation, as the newly acquired information is encoded within a subspace of lower rank than the original representation.
+First, this module mitigates the computationally expensive limitation of full fine-tuning by updating only a very small subset of the network parameters and compact low-rank matrices, thereby reducing trainable parameters by several orders of magnitude.
+Second, this module alleviates catastrophic forgetting of pre-trained knowledge during adaptation, as the newly acquired information is encoded within a subspace rather than the original representation.
 Consequently, this module enables the finetuned model to retain its pre-trained general-purpose knowledge while efficiently adapting to downstream radar tasks.
-Concretely, we fine-tune the query and value matrices, $\mathbf{W}_{Q}$ and $\mathbf{W}_{V}$, within each attention block of the LLM backbone, as illustrated in Fig.
-5.
-LoRA introduces two trainable rank decomposition matrices, $A$ and $B$, attached to the frozen pre-trained weight matrix $\mathbf{W}$.
-The modified transformation is expressed as:
-$\mathbf{W}\leftarrow\mathbf{W}+AB^{T},$ (11)
-where $\mathbf{W}\in\mathbb{R}^{d{1}\times d{2}}$, $A\in\mathbb{R}^{d_{1}\times r}$, $B\in\mathbb{R}^{d_{2}\times r}$, and $r\ll\min(d_{1},d_{2})$.
-The matrices are initialized as $A\sim\mathcal{N}(0,\sigma^{2})$ and $B=0$.
-During fine-tuning, $\mathbf{W}$ remains frozen while $A$ and $B$ are optimized.
-Owing to the low-rank constraint ($r\ll\min(d_{1},d_{2})$), this approach dramatically reduces the number of trainable parameters compared to full-parameter fine-tuning.
 Figure 5: Illustration of the LoRA fine-tuning process.
@@
 III-D1 Methodology
-To enhance the fine-tuning of LLMs for radar target detection under low SCR conditions, we introduce a token-level preference-aware loss function.
-In such scenarios, a substantial proportion of feature patches contribute little information for effective learning.
-Uniform optimization over all feature patches risks overfitting by coercing the model to memorize uninformative or even misleading patterns.
-To mitigate this, we propose a dynamic re-weighting scheme that adaptively assigns higher learning weights to feature patches exhibiting superior generalization potential.
-Specifically, our approach uses a dual-model framework consisting of a lightweight reference model $\theta_{r}$ and a high-capacity target model $\theta_{t}$, which leverages LLMs.
-  • Reference Model: The reference model is built with a lightweight Transformer encoder and trained solely on validation samples using standard cross-entropy loss.
-By leveraging the reference model, we can estimate the generalization potential of each feature token in the training samples.
-A lower reference loss indicates that the token can be effectively captured by a lightweight model, suggesting strong generalization potential.
-In contrast, a higher reference loss typically reflects excessive noise, making the token less useful for improving detection performance.
-  • Target Model: A significantly larger model built upon a pre-trained LLM, parameterized by $\theta_{t}$.
-For a batch of training samples, we compute the token-level cross-entropy losses for both models. Given a feature token $F_{b,k}^{\mathrm{fi}}$, the corresponding two losses are as follows:
-  • Reference Loss:
+To address the prevalent issue of overfitting in PLM-based model fine-tuning, we introduce a preference-aware loss function to enable selective training, a novel strategy designed to guide the model towards learning generalizable feature patterns.
+During fine-tuning, feature patches dominated by clutter-induced noise or anomalies can mislead the model and hinder effective learning.
+Conventional optimization methods, which treat all feature patterns indiscriminately with a standard cross-entropy loss, are prone to memorizing these non-generalizable, instance-specific characteristics, leading to severe overfitting.
+In contrast, our proposed selective training strategy follows a ”preference-aware” philosophy.
+The core idea is to evaluate the learning value of each feature patch during training and use this evaluation to reweight its contribution to the final loss.
+This approach enables the model to focus on feature patterns with high learning values while minimizing the impact of noisy or anomalous patterns.
+By preventing the model from overfitting to these non-generalizable patterns, this strategy enhances model robustness and mitigates overfitting.
+Specifically, for each feature token $F_{b,k}^{\mathrm{fi}}$ within a training batch, we compute two distinct loss values:
+  • Reference Loss $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})$: This loss is computed using a lightweight reference model $\theta_{r}$, which has been trained on the validation set.
+It serves as a stable proxy for evaluating the feature patch’s inherent difficulty or noise level:
     $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})=-\sum_{i}y_{b,k}^{i}\log(\hat{y}_{b,k}^{i,\theta_{r}}),$ (12)
-    where $\hat{y}_{b,k}^{i,\theta_{r}}$ is the prediction from $\theta_{r}$.
-  • Target Loss:
+    where $\hat{y}_{b,k}^{i,\theta_{r}}$ is the prediction from lightweight reference model $\theta_{r}$.
+  • Target Loss $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})$: his loss is calculated using the predictions from the current PLM-based model $\theta_{t}$ being fine-tuned, which reflects how well the model currently fits the feature patch:
     $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})=-\sum_{i}y_{b,k}^{i}\log(\hat{y}_{b,k}^{i,\theta_{t}}),$ (13)
-    where $\hat{y}_{b,k}^{i,\theta_{t}}$ is generated by the fine-tuned much larger model $\theta_{t}$.
-The core of our method lies in computing a token importance score $s_{b,k}$ based on the discrepancy between the target loss and reference loss:
+    where $\hat{y}_{b,k}^{i,\theta_{t}}$ represents the prediction from the current PLM-based model $\theta_{t}$.
+The key to our method is computing an evaluated learning value $s_{b,k}$ based on the discrepancy between the target and reference losses:
 $s_{b,k}=\mathrm{ReLU}(\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})-\alpha\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})),$ (14)
-where $\alpha$ is a scaling coefficient, and $\mathrm{ReLU}(\cdot)$ ensures that only tokens with positive excess loss are trained.
-Tokens with low importance score are not prioritized for optimization because they have already been well learned, as indicated by a low target loss $\mathcal{L}_{\theta_{{t}},b,k}$, or with large noise, as reflected by a high reference loss $\mathcal{L}_{\theta_{{r}},b,k}$.
-The $\mathrm{ReLU}(\cdot)$ function is defined as:
-$\operatorname{ReLU}(x)=\left\{\begin{array}[]{ll}x&x>0\\ 0&x\leq 0\end{array}\right$. (15)
-More detailedly, this strategy results in the following selective training scheme:
-  • If the target loss is much larger than the reference loss, the token is deemed with large learning value, and is therefore prioritized for training.
-  • If the reference loss is large, the token shows weak generalization potential and a high risk of overfitting; therefore, its training weight is reduced during optimization.
-  • If the target loss is already small, the token has been sufficiently learned, and further training brings little gain, the training weight is also reduced to avoid redundant updates.
+where $\alpha$ is a scaling coefficient, and the $\mathrm{ReLU}(\cdot)$ function, defined as $\mathrm{ReLU}(x)=\max(0,x)$, ensures that only tokens with positive excess loss are fine-tuned.
+This loss difference naturally leads to a selective training scheme with desirable properties:
+  • High Importance (Unlearned): If $\mathcal{L}_{\theta_{{t}}}$ is much larger than $\alpha\mathcal{L}_{\theta_{{t}}}$, the feature token is identified as containing valuable, yet unlearned patterns.
+  • Low Importance (Noise/Anomaly): If $\alpha\mathcal{L}_{\theta_{{r}}}$ is large (indicating with high inherent noise), it effectively dampens $s_{b,k}$, even if $\mathcal{L}_{\theta_{{t}}}$ is large, preventing the model from overfitting to noisy or anomalous data.
+  • Low Importance (Well Learned): If $\mathcal{L}_{\theta_{{t}}}$ is already small, the feature token is considered well learned, and $s_{b,k}$ will also be small, preventing redundant updates to easy patterns.
 The final training loss is a weighted sum over all feature tokens across the mini-batch, where each token’s contribution is scaled by a token-level importance score:
@@
 III-D2 Theoretical Insight of the Effectiveness of Feature Token Importance
-We begin by formalizing the task of selecting informative feature patches for training.
-At each step $t$, given a candidate set $B_{t}$, our objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ achieves the maximal reduction in generalization loss on the test set $D_{\mathrm{te}}$.
-This can be formally formulated as the following optimization problem:
+A.
+Problem Definition: Inspired by [26], we begin by formalizing an optimization problem of selecting useful feature patches for model fine-tuning.
+At each training step $t$, given a candidate set $B_{t}$, the objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ maximally reduces the generalization loss on the unseen test set $D_{\mathrm{te}}$.
+This can be expressed as the following optimization problem:
 $\arg\min_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}-\log p\left(y^{\text{te }}\mid F^{\text{te }};D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right).$ (16)
-Directly optimizing Eq.(17) is computationally prohibitive, as it entails retraining the model for each candidate feature patch.
-To derive a tractable surrogate, we reformulate Eq.(17) using Bayes’ theorem:
+However, directly optimizing Eq.(16) is computationally prohibitive, as it requires retraining the model for each candidate patch.
+To enable a practical surrogate, we reformulate this objective using probabilistic decomposition.
+B. Derivation of the Surrogate Objective: Applying Bayes’ theorem, the term in Eq.(16) can be decomposed as:
 $-\log p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};D_{t}\!\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\bigr)$
 $\quad=-\log\frac{p\!\bigl(y,y^{\mathrm{te}}\mid F,F^{\mathrm{te}},D_{t}\bigr)}{p\!\bigl(y\mid F;D_{t}\bigr)}.$ (17)
-Assuming that each label is statistically independent of the other pairs in the corpus when conditioned on its own feature vector, the joint term in the numerator can be expressed as:
+Assuming that each label is statistically independent of all other labels in the corpus when conditioned on its corresponding feature vector, the joint probability term in the numerator can be expressed as follows:
 $p\!\bigl(y,y^{\mathrm{te}}\mid F,F^{\mathrm{te}},D_{t}\bigr)\\ =p\!\bigl(y\mid F;F^{\mathrm{te}},y^{\mathrm{te}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};F,D_{t}\bigr)\\ =p\!\bigl(y\mid F;y^{\mathrm{te}},F^{\mathrm{te}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};D_{t}\bigr).$ (18)
-Substituting Eq.(19) into Eq.(18) gives:
+Substituting Eq.(18) into Eq.(17) and taking the logarithm yields:
 $\log p\!\left(y^{\text{te}}\mid F^{\text{te}};\,D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right)$ (19)
@@
 $\arg\max_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right].$ (20)
-Eq.(21) indicates that feature patches with larger loss differences contribute more significantly to improve the model performance on test set; thus, the loss difference can be interpreted as the learning value of each feature patch.
-However, since accessing the test set for retraining on $D_{t}\cup D_{\mathrm{test}}$ is prohibited and computationally impractical, we approximate the second term using a lightweight reference model trained solely on validation set $D_{\mathrm{val}}$.
-Consequently, the practical token weight is derived as shown in Eq.(14).
+As exact Bayesian inference is intractable in neural networks, we fit the models with SGD instead.
+Here, $\mathcal{L}[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}]$ denotes the training loss on the patch using the current model trained on $D_{t}$, whereas $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right]$ represents the excepted irreducible holdout loss obtained if the model were trained on both the training set and the test set.
+C.
+Conclusion: Eq.(20) reveals that the gap between the training loss and the irreducible holdout loss quantifies the reducible component of the generalization error of each feature patch.
+Consequently, this loss difference can be rigorously interpreted as the learning value of each feature patch, representing how much its inclusion in model training is expected to improve performance on unseen data.
+However, accessing the test set for training on $D_{t}\cup D_{\mathrm{test}}$ is both prohibited and computationally impractical.
+To address this limitation, the validation set is adopted as a statistically sound and feasible proxy for the unseen data distribution.
+Accordingly, we approximate $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right]$ using the loss computed by a lightweight reference model trained on the validation set $D_{\mathrm{val}}$ [26].
+Consequently, the resulting token weight derived in Eq.
+(14) provides a practical and theoretically grounded measure of each feature patch’s learning value.
 III-D3 Comparison with Existing Loss Designs
@@
 In general, the proposed preference-aware loss marks a major and substantial advance in enhancing model generalization across diverse and low-SCR radar detection scenarios.
-III-E False Alarm Controllable Classification Head Retraining
+III-E Autoencoder-based Binary Classification Head Retraining
 As shown in Fig.
-1, the LLM-based model outputs $[B,K,2]$ token-level predictions, which lack a unified detection representation.
-To achieve consistent decision-making, we retrain a classifier using the learned feature $F^{\mathrm{LN}}$.
-An autoencoder-based classification head extracts informative features from this high-dimensional input for robust detection.
-It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative features; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector.
+1 and Section III-A, the PLM-based model outputs $[B,K,2]$ token-level predictions, which lack a unified detection representation.
+To achieve consistent decision-making, we retrain a binary classifier head using the learned representation $F^{\mathrm{LN}}$.
+An autoencoder-based biary classification head extracts informative features from this high-dimensional input for robust detection.
+It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative representation; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector.
 The autoencoder is optimized by a dual-objective loss combining reconstruction and classification terms:
@@
 where $\sigma_{\text{recon}}$ and $\sigma_{\text{ce}}$ are learnable task uncertainties that regulate the joint optimization dynamics.
-During inference, false alarms are controlled via threshold-setting using sorted softmax outputs of clutter samples.
+During inference, false alarm rate is controlled via threshold-setting using sorted softmax outputs of clutter samples.
 The detection threshold $\eta$ is then determined based on the desired false alarm rate $P_{fa}^{\text{d}}$ as follows:
 $\begin{array}[]{l}\eta=O_{\mathrm{sorted}}(x),\ x=\lceil P_{fa}^{\text{d}}\times N_{\text{clutter}}\rceil,\end{array}$ (24)
-where $O_{\mathrm{sorted}}(x)$ is the $x$-th largest output and $N_{\text{clutter}}$ is the total number of test clutter samples.
+where $O_{\mathrm{sorted}}(x)$ is the $x$-th largest output and $N_{\text{clutter}}$ is the total number of test clutter samples. The overall procedure of the RadarPLM is presented in Algorithm 1.
+Algorithm 1 RadarPLM Training Protocol
+1:
+2:Radar echo signals $\{x(n)\}_{n=1}^{N}$ and Target labels $\{y_{b}\}_{b=1}^{B}$
+3:Pre-trained LLM parameters $\theta_{\text{LLM}}$
+4:Stage 1: Sequence Feature Extraction
+  1. 1.
+    Extract IP, DSE, SMS, Amp, and DP features from the observation vector, respectively, see Eq.(3), (4), (5),(7), and (9).
+  2. 2.
+    Patch the five features into feature tokens, see Fig. 3.
+5:Stage 2: Reference Model Training
+  1. 1.
+    Train a lightweight reference model $\theta_{r}$ based on validation samples and CE loss.
+  2. 2.
+    Compute the loss for each feature token of the training samples based on the predicted output probabilities of reference model, see Eq.(12).
+6:Stage 3: Fine-tune PLM for Target Detection
+  1. 1.
+    Compute the loss for each feature token of the training samples based on the predicted output probabilities of current model, see Eq.(13).
+  2. 2.
+    Fine-tune PLM for target detection via preference-aware loss function, see Eq.(14) and Eq.(15).
+7:Stage 4: Binary Classification Head Retraining
+  1. 1.
+    Train autoencoder-based binary classification head based on the weighted sum of reconstruction loss and CE loss, see Eq.(21), Eq.(22), and Eq.(23).
+  2. 2.
+    Adjust detection threshold $\eta$ for controllable FAR, see Eq.(24).
+8:
+9:Fine-tuned RadarLLM model with controllable FAR.
 IV Experiments
-In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarLLM’s detection performance under two settings: sufficient training data and constrained training data scenarios.
+TABLE I: Overview of the IPIX Database.
+Label Name File name Primary Secondary SCR
+1 17 19931107_135603_starea 9 8,10,11 Low
+2 25 19931108_213827_starea 7 6,8 Low
+3 26 19931108_220902_starea 7 6,8 Low
+4 54 19931111_163625_starea 8 7,9,10 High
+5 280 19931118_023604_stareC0000 8 7,9,10 High
+6 283 19931118_035737_stareC0000 10 8,9,11,12 High
+7 311 19931118_162658_stareC0000 7 6,8,9 High
+8 320 19931118_174259_stareC0000 7 6,8,9 High
+TABLE II: Hyperparameters for network training.
+Parameter Value
+Batch size 64
+Eval batch size 400
+Epochs for LLM fine-tuning 500
+Epochs for Autoencoder training 300
+Optimizer Adam (betas=(0.9,0.999))
+Learning rate for PLM fine-tuning 0.0001
+Learning rate for Autoencoder training 0.00001
+TABLE III: Hyperparameters for Autoencoder network.
+Block Layer Filter Stride Padding Output
+[…]
+[…]
+[…]
+[…]
+[…]
+Conv6+ReLU 16,1*1 1 Same 55*16
+Flatten+Fully-Connected1 - - - 64
+Fully-Connected2 - - - 20
+Decoder Fully-Connected3+Reshape - - - 55*16
+Conv1+ReLU 32,1*1 1 1 55*32
+Conv2+ReLU 64,1*1 1 1 55*64
+Conv3+ReLU 128,1*1 1 1 55*128
+Conv4+ReLU 256,1*1 1 1 55*256
+Conv5+ReLU 512,1*1 1 1 55*512
+Conv6+ReLU 768,1*1 1 1 55*768
+Classifier GeLU+LayerNorm - - - 20
+Fully-Connected4 - - - 2
+In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarPLM’s detection performance under two settings: sufficient training data and limited training data scenarios.
 IV-A Experimental Setup
@@
 The samples are divided into three groups: (1) a training set using the first 20% observation time for both the target and clutter cells, (2) a validation set covering 20% to 35% of the observation time, and (3) a test set containing the remaining samples.
-TABLE I: Overview of the IPIX Database.
-Label Name File name Primary Secondary SCR
-1 17 19931107_135603_starea 9 8,10,11 Low
-2 25 19931108_213827_starea 7 6,8 Low
-3 26 19931108_220902_starea 7 6,8 Low
-4 54 19931111_163625_starea 8 7,9,10 High
-5 280 19931118_023604_stareC0000 8 7,9,10 High
-6 283 19931118_035737_stareC0000 10 8,9,11,12 High
-7 311 19931118_162658_stareC0000 7 6,8,9 High
-8 320 19931118_174259_stareC0000 7 6,8,9 High
 IV-A2 Baselines
-To validate the superiority of RadarLLM, eight deep learning-based marine radar target detection methods are implemented as baselines.
-Among them, RNN, Bi-LSTM, GRU, Transformer, PatchTST, and ResNet18 represent lightweight models that utilize sequence features; OFA uses a large language model backbone to process sequence features; and ADN18 utilizes time-frequency maps and enhanced CNNs for detection.
+To validate the superiority of RadarPLM, eight deep learning-based marine radar target detection methods are implemented as baselines.
+Among them, RNN, Bi-LSTM, GRU, Transformer, PatchTST, and ResNet18 represent lightweight models that process sequence features; OFA uses a PLM backbone to process sequence features; and ADN18 utilizes time-frequency maps and enhanced CNNs for detection.
   • RNN [14]: RNN is a sequential neural architecture with temporal feedback loops. In experiments, the number of RNN layers and the hidden layer nodes are set to 2 and 6, respectively.
@@
 All network training experiments are conducted on a system equipped with an E5-2695v3 CPU, an NVIDIA 3090Ti GPU, and 64 GB of RAM.
-TABLE II: Hyperparameters for network training.
-Parameter Value
-Batch size 64
-Eval batch size 400
-Epochs for LLM fine-tuning 500
-Epochs for Autoencoder training 300
-Optimizer Adam (betas=(0.9,0.999))
-Learning rate for LLM fine-tuning 0.0001
-Learning rate for Autoencoder training 0.00001
-TABLE III: Hyperparameters for Autoencoder network.
-Block Layer Filter Stride Padding Output
-[…]
-[…]
-[…]
-[…]
-[…]
-Conv6+ReLU 16,1*1 1 Same 55*16
-Flatten+Fully-Connected1 - - - 64
-Fully-Connected2 - - - 20
-Decoder Fully-Connected3+Reshape - - - 55*16
-Conv1+ReLU 32,1*1 1 1 55*32
-Conv2+ReLU 64,1*1 1 1 55*64
-Conv3+ReLU 128,1*1 1 1 55*128
-Conv4+ReLU 256,1*1 1 1 55*256
-Conv5+ReLU 512,1*1 1 1 55*512
-Conv6+ReLU 768,1*1 1 1 55*768
-Classifier GeLU+LayerNorm - - - 20
-Fully-Connected4 - - - 2
-IV-B Comparison of Detection Performance, Speed, and Network Parameters Under Sufficient Training Samples
+IV-B Performance Comparison on Sufficient Training Samples
 (1). HH
@@
 Figure 6: Comparison of model detection performance on various datasets.
-TABLE IV: Detection performance of different methods when the false alarm rate is 0.005. The best and second-best detection results are highlighted in red and blue, respectively.
-Detector $H$ HV VH $V$ AVG
+TABLE IV: Comparison of detection rates (%) on sufficient training samples.
+Method $H$ HV VH $V$ AVG
 RNN [14] 46.04 46.66 38.18 17.04 36.98
@@
 ADN18 [43] 79.74 83.50 82.01 78.02 80.82
-RadarLLM 82.51 84.81 84.61 75.90 81.96
-TABLE V: Network parameters (training parameters/total parameters) and interference time of different methods.
-Detector Average interference time on one dataset(s) Total network parameters(M) Training network parameters(M)
+RadarPLM 82.51 84.81 84.61 75.90 81.96
+(All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.)
+TABLE V: Network parameters (training parameters / total parameters) and average inference time of different methods.
+Detector Avg. inference time (s) Total Params (M) Trainable Params (M)
 RNN [14] 5.3587 0.0003 0.0003
@@
 ADN18 [43] 160.9018 14.3315 14.3315
-RadarLLM 5.6561 69.3932 2.4567
+RadarLLM (ours) 5.6561 69.3932 2.4567
 Table IV reports the detection rates achieved by RadarLLM and the eight baseline detectors introduced in Section IV-A2.
@@
 Compared to the best competing sequence feature detector, RadarLLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively.
-Figure IV presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates.
+Figure LABEL:DT presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates.
 We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone.
 Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in the average detection rate.
-Furthermore, we compare RadarLLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment.
+Furthermore, we compare RadarPLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment.
 The comparison results are summarized in Table V.
-Although RadarLLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning.
-In particular, RadarLLM achieves an exceptionally low average inference latency.
+Although RadarPLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning.
+In particular, RadarPLM achieves an exceptionally low average inference latency.
 Specifically, it requires only 5.6561 seconds to process more than 9,000 test samples, which is 28.44 times faster than the ADN18 method.
@@
 Moreover, the intrinsic inference acceleration of the GPT architecture further enhances this performance gain.
-Together, these advantages make RadarLLM a promising solution for real-time marine target detection.
-IV-C Comparison of Detection Performance Under Limited Training Samples With Baselines
+Together, these advantages make RadarPLM a promising solution for real-time marine target detection.
+IV-C Performance Comparison on Limited Training Samples
 Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge.
-Considering this issue, we conducted experiments under constrained training sample conditions to assess the effectiveness of RadarLLM.
+Considering this issue, we conducted experiments under constrained training sample conditions to assess the effectiveness of RadarPLM.
 The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
-In Table VI, we report the detection performance of RadarLLM and several baseline methods under scenarios with limited training data.
-Compared with the experimental setup described in Section IV-A1, the advantages of RadarLLM become more pronounced.
-Specifically, RadarLLM achieves a substantial increase of 19.66% in the average detection rate over the previously proposed sequence-feature based detector [14].
+In Table VI, we report the detection performance of RadarPLM and several baseline methods under scenarios with limited training data.
+Compared with the experimental setup described in Section IV-A1, the advantages of RadarPLM become more pronounced.
+Specifically, RadarPLM achieves a substantial increase of 19.66% in the average detection rate over the previously proposed sequence-feature based detector [14].
 This marked improvement can be mainly attributed to three key factors:
-  • First, leveraging pre-trained knowledge embedded in the large language model (LLM) provides RadarLLM with a beneficial parameter initialization, facilitating efficient optimization even when training data is limited.
-This prior knowledge significantly reduces the data dependency typical of conventional training strategies.
-  • Second, the substantial parameter scale of RadarLLM inherently grants robustness to distributional variations, enabling superior generalization compared to smaller-scale neural networks.
+  • First, The embedded knowledge from pre-trained parameters endows RadarPLM with a highly advantageous parameter initialization.
+This acts as a powerful springboard for optimization, enabling effective convergence even in data-limited regimes and drastically mitigating the data-hungry nature of conventional training paradigms.
+  • Second, the substantial parameter scale of RadarPLM inherently grants robustness to distributional variations, enabling superior generalization compared to smaller-scale neural networks.
 Its extensive capacity allows it to capture complex data structures and efficiently handle variability in different radar detection scenarios.
-  • Third, the integration of a preference-aware loss function significantly improves RadarLLM detection performance, especially in challenging scenarios characterized by low signal-to-clutter ratios (SCR).
-TABLE VI: Detection performance under limited training samples of different methods when the false alarm rate is 0.005.
-The best and second-best detection results are highlighted in red and blue, respectively.
-Detector $H$ HV VH $V$ AVG
+  • Third, the integration of a preference-aware loss function significantly improves RadarPLM detection performance, especially in challenging scenarios characterized by low SCR.
+TABLE VI: Comparison of detection rates (%) on limited training samples.
+Method $H$ HV VH $V$ AVG
 RNN [14] 35.49 30.13 31.82 15.14 28.15
@@
 ADN18 [43] 75.45 81.50 80.81 70.97 77.18
-RadarLLM 79.26 82.31 81.51 72.54 78.91
-(1). Data1 HH
-(2). Data1 HV
-(3). Data1 VH
-(4). Data1 VV
-(5). Data2 HH
-(6). Data2 HV
-(7). Data2 VH
-(8). Data2 VV
-(9). Data3 HH
-(10). Data3 HV
-(11). Data3 VH
-(12). Data3 VV
-Figure 7: Comparison of cross-entropy loss and preference-aware loss on low SCR datasets.
+RadarPLM 79.26 82.31 81.51 72.54 78.91
+(All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.)
+(1). Data1
+(2). Data2
+(3). Data3
+Figure 7: Comparison of detection rates between the proposed preference-aware loss and the conventional cross-entropy loss.
 We compare detection rates across different datasets and epochs, highlighting the improvements achieved by the preference-aware loss function.
 IV-D Ablation Study
-IV-D1 Ablation Study on Backbone Architectures and Fine-tuning Strategies
-To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the pre-trained LLM:
+IV-D1 Backbone Architectures
+To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the PLM under sufficient training samples setting:
   1. 1.
-    RadarLLM (0): In this variant, the pre-trained LLM is entirely removed, while all other components of the framework are kept unchanged.
+    RadarPLM (0): In this variant, the PLM backbone is entirely removed, while all other components of the framework are kept unchanged.
   2. 2.
-    RadarLLM (T): In this variant, the pre-trained LLM is replaced with a randomly initialized transformer block.
+    RadarPLM (T): In this variant, the PLM backbone is replaced with a randomly initialized transformer block.
   3. 3.
-    RadarLLM (R): In this variant, the network architecture and fine-tuning procedure of the original LLM are maintained, but its pre-trained parameters are replaced with random initialization.
+    RadarPLM (R): In this variant, the network architecture and fine-tuning procedure are maintained, but pre-trained parameters of PLM are replaced with random initialization.
 In Table VII, we show the results of the ablation study on backbone choice.
-It can be seen that the removal or exchange of pre-trained LLM results in notable performance degradation, indicating the necessity of pretrained LLM for high detection performance.
-TABLE VII: Results of ablation study on the backbone choice.
-$H$ HV VH $V$ AVG
-RadarLLM 82.51 84.81 84.61 75.90 81.96
-RadarLLM (0) 77.51 79.38 80.35 69.97 76.80
-RadarLLM (T) 78.25 79.77 81.16 72.13 77.83
-RadarLLM (R) 75.19 75.44 75.81 65.15 72.90
-Moreover, to verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study as summarized in Table VIII, and the corresponding results are presented in Table IX.
-The results demonstrate that jointly fine-tuning the LoRA modules and Layer Normalization parameters achieves the highest overall performance.
-In contrast, fine-tuning only the Layer Normalization parameters, the LoRA modules, or performing full fine-tuning results in moderate performance degradation.
-These findings indicate that the proposed lightweight fine-tuning strategy effectively balances model adaptability and computational efficiency while preserving the universal knowledge embedded in the LLM backbone.
-TABLE VIII: Ablation Study on Fine-tuning Strategies settings.
-LoRA Layer Norm Other Params
-RadarLLM $\checkmark$ $\checkmark$
-RadarLLM (F) $\checkmark$ $\checkmark$
-RadarLLM (LoRA) $\checkmark$
-RadarLLM (LN) $\checkmark$
-TABLE IX: Results of ablation study on the finetuning strategy.
-$H$ HV VH $V$ AVG
-RadarLLM 82.51 84.81 84.61 75.90 81.96
-RadarLLM (F) 81.69 83.45 84.41 74.07 80.91
-RadarLLM (LoRA) 82.20 84.45 84.56 75.53 81.69
-RadarLLM (LN) 82.33 83.73 83.80 75.12 81.25
-IV-D2 Ablation Study on the Loss Function
+It can be seen that the removal or exchange of PLM backbone results in notable performance degradation, indicating the necessity of PLM for high detection performance.
+TABLE VII: Results of ablation study on the backbone choice and fine-tuning strategies.
+RadarPLM(0)–RadarPLM(R) denote backbone ablation results, while RadarPLM(F)–RadarPLM(LN) denote fine-tuning strategy ablation results.
+Model $H$ HV VH $V$ AVG
+RadarPLM 82.51 84.81 84.61 75.90 81.96
+Backbone choice ablation
+RadarPLM (0) 77.51 79.38 80.35 69.97 76.80
+RadarPLM (T) 78.25 79.77 81.16 72.13 77.83
+RadarPLM (R) 75.19 75.44 75.81 65.15 72.90
+Fine-tuning strategy ablation
+RadarPLM (F) 81.69 83.45 84.41 74.07 80.91
+RadarPLM (LoRA) 82.20 84.45 84.56 75.53 81.69
+RadarPLM (LN) 82.33 83.73 83.80 75.12 81.25
+IV-D2 Fine-tuning Strategies
+To verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study under the setting of sufficient training samples, as summarized in Table VIII, and the corresponding results are presented in Table VII.
+In this experiment, four fine-tuning configurations were evaluated: (1) RadarPLM: jointly fine-tuning the LoRA modules and Layer Normalization (LN) parameters, (2) RadarPLM (LoRA): fine-tuning only the LoRA modules, (3) RadarPLM (LN): fine-tuning only the LN parameters, and (4) RadarPLM (F): full fine-tuning of all model parameters.
+As shown in the results, the joint fine-tuning of LoRA and LN achieves notably higher performance, while maintaining a relatively shorter training cost (0.116 h/100 epochs) than full fine-tuning (0.148 h/100 epochs).
+This demonstrates that the proposed design not only enhances model adaptability, but also significantly reduces computational cost.
+TABLE VIII: Ablation study on fine-tuning strategy configurations and their corresponding training time (measured over 100 training epochs).
+Model LoRA Layer Norm Other Params Training Time (hour/100 epochs)
+RadarPLM $\checkmark$ $\checkmark$ 0.116
+RadarPLM (F) $\checkmark$ $\checkmark$ 0.148
+RadarPLM (LoRA) $\checkmark$ 0.116
+RadarPLM (LN) $\checkmark$ 0.112
+IV-D3 Loss Function
 In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets.
@@
 In Fig.
-7, we evaluate the detection rate on the test samples of different low SCR datasets at different training epochs.
-It can be seen that preference-aware loss achieves an improvement of around 1%-15% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets.
-In Table X, we show the detection performance of the reference model.
-It can be seen that the optimal detection performance of the fine-tuned model outperforms the reference model by approximately 3%-18% across various datasets.
-In particular, despite this remarkably significant performance gap, leveraging the reference model to score individual feature tokens still yields significant performance gains.
+7, we present the detection rates on test samples from different low-SCR datasets under HH polarization across multiple training epochs.
+It can be observed that preference-aware loss achieves an improvement of around 1%-12% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets.
+We also include the performance of the reference model (gray line) for comparison.
+it can be seen that the optimal detection performance of the fine-tuned PLM-based model outperforms the reference model by approximately 3%-18% across various datasets.
+Notably, despite this substantial performance gap, employing the reference model to construct the preference-aware loss still leads to significant performance gains, demonstrating that our method effectively realizes a compelling weak-to-strong generalization capability.
 Although the preceding results are promising, a direct comparison between the proposed preference-aware (PA) loss and the conventional end-to-end cross-entropy (CE) loss remains lacking.
@@
 For further comparison, we include a sample-level reweighting baseline, denoted as WCE [26], where training weights are adjusted at the sample level.
-As summarized in Table XI, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions.
-For example, RadarLLM (PA) achieves an average gain of 9.9% over RadarLLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart.
+As summarized in Table IX, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions.
+For example, RadarPLM (PA) achieves an average gain of 9.9% over RadarPLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart.
 In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
-TABLE X: Detection performance of reference model on three low SCR datasets.
-Reference model $H$ HV VH $V$
-Data1 46.68 45.21 46.75 18.48
-Data2 34.87 50.75 48.41 42.51
-Data3 47.39 59.82 61.85 44.72
-TABLE XI: Results of ablation study on the loss function.
-Data1 Data2 Data3 AVG
-RadarLLM (PA) 52.06(+16.8) 54.87(+5.4) 65.81(+7.5) 57.58(+9.9)
-RadarLLM (WCE) 43.51(+8.8) 53.95(+4.9) 64.55(+6.8) 54.84(+7.7)
-RadarLLM (CE) 34.73 48.96 57.80 47.16
+TABLE IX: Results of ablation study on the loss function.
+Method Data1 Data2 Data3 AVG
+RadarPLM (PA) 52.06 (+16.8) 54.87 (+5.4) 65.81 (+7.5) 57.58 (+9.9)
+RadarPLM (WCE) 43.51 (+8.8) 53.95 (+4.9) 64.55 (+6.8) 54.84 (+7.7)
+RadarPLM (CE) 34.73 48.96 57.80 47.16
 PatchTST (PA) 47.48 (+8.1) 50.29 (+5.9) 62.42 (+7.0) 53.40 (+7.0)
@@
 PatchTST (CE) 38.91 43.92 54.89 45.91
+TABLE X: Performance comparison of RadarPLM under different patch sizes and GPT2 layer numbers.
+The results include average detection rate (DR), average inference time (IT), and total/training network parameters (NP, in millions).
+RadarPLM Performance under Different Configurations
+Patch Size Configuration Average DR (%) Average IT (s) Total / Training NP (M)
+32 79.44 5.9259 69.39 / 2.46
+48 81.96 5.6561 69.39 / 2.46
+64 79.09 5.6222 69.39 / 2.46
+GPT2 Layers 0 Layer 77.41 5.3484 41.04 / 2.44
+2 Layers 81.29 5.4812 55.22 / 2.45
+4 Layers 81.96 5.6561 69.39 / 2.46
+6 Layers 80.42 5.7331 83.57 / 2.46
+8 Layers 80.01 5.8655 97.74 / 2.47
 IV-E Analysis of the Effect of Hyperparameters on Experimental Results
-TABLE XII: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of patch size.)
-Average DR Average IT(s)
-RadarLLM (with a patch size of 32) 79.44 5.9259
-RadarLLM (with a patch size of 48) 81.96 5.6561
-RadarLLM (with a patch size of 64) 79.09 5.6222
-We first investigate the impact of patch size on RadarLLM’s performance.
-The results of the ablation study are summarized in Table XII.
+We first investigate the impact of patch size on RadarPLM’s performance.
+The results of the ablation study are summarized in Table X.
 Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes.
-These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarLLM.
-We then investigated the impact of the number of GPT2 layers on RadarLLM’s performance.
-Our ablation study on GPT2 layer configurations is presented in Table XIII.
+These results highlight the importance of selecting an appropriate patch size to optimize the performance of RadarPLM.
+We then investigated the impact of the number of GPT2 layers on RadarPLM’s performance.
+Our ablation study on GPT2 layer configurations is presented in Table X.
 Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones.
-These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the cross-modal transfer capabilities of the language model, while excess layers suffer from overfitting due to the task-irrelevant network parameters in deeper layers.
+These results highlight two key observations: (1) a model with insufficient layers fails to fully leverage the parameter transfer capabilities of the PLM, while excess layers suffer from overfitting due to the task-irrelevant parameters in deeper layers.
 (2) Although utilizing more GPT2 layers significantly increases the scale of model parameters, the impact on inference latency is minimal.
 Specifically, the inference speed of GPT2 (8) remains approximately 91% that of GPT2 (0), as the inherent inference acceleration of the GPT architecture mitigates the computational overhead introduced by deeper layers.
-TABLE XIII: The average detection rate, interfere time on one dataset, and total/training network parameters of RadarLLM with different number of GPT2 layers.
-(RadarLLM (n) represents RadarLLM with n LLM layers.))
-Average DR Average IT(s) Total/Training NP(M)
-RadarLLM (0) 77.41 5.3484 41.0410/2.4445
-RadarLLM (2) 81.29 5.4812 55.2175/2.4506
-RadarLLM (4) 81.96 5.6561 69.3932/2.4567
-RadarLLM (6) 80.42 5.7331 83.5690/2.4628
-RadarLLM (8) 80.01 5.8655 97.7447/2.4690
-IV-F Case Study of the Effectiveness of Feature Token Importance
-To further demonstrate the effectiveness of the proposed preference-aware loss function, we conduct a case study analyzing the training weight dynamics of feature tokens for two representive sequence features: Short-Time Fourier Transform Magnitude Spectrum (SMS) and Instantaneous Phase (IP) features.
-The SMS feature exhibits a distinctive rise and fall pattern indicative of target presence, mainly characterized by the feature tokens 26 and 27.
-As illustrated in Fig.
-7, these tokens initially receive high attention weights during training, underscoring the model’s emphasis on learning these salient patterns.
-However, as training progresses, the model rapidly assimilates these features, resulting in a marked decline in their weights after approximately 10 epochs.
-This adaptive behavior reflects the model’s ability to dynamically reallocate attention from well-learned patterns to more complex or informative features, thus mitigating the risk of overfitting to easily recognizable signal features.
-In contrast, the IP feature demonstrates a more complex structure, encompassing both moderately varying patterns associated with target echos (corresponding to tokens 0, 3, and 4) and rapidly fluctuating patterns caused by clutter-induced noise (corresponding to tokens 8, 9, and 10).
-These six tokens, arranged from left to right, align with the gray-lighted segments in the IP feature map illustrated in Fig.
-8.
-During training, tokens 0, 3, and 4 consistently receive elevated attention weights, typically exceeding 1, indicating the model’s sustained emphasis on these informative yet challenging-to-learn feature tokens.
-In contrast, tokens 8, 9, and 10 are assigned substantially lower weights, which highlights the model’s ability to suppress excessive learning from noisy and less informative feature tokens.
-(1). SMS feature
-(2). SMS feature token weight dynamics
-(3). IP feature
-(4). IP feature token weight dynamics
-Figure 8: Dynamics of feature token weights during training: (a) SMS feature token weight dynamics, (b) IP feature token weight dynamics.
-The graphs show how the attention weights assigned to individual feature tokens evolve over epochs, emphasizing the model’s ability to adaptively focus on informative features and suppress noisy tokens.
-V Conclusion and Future Work
-In this study, we propose RadarLLM, a novel framework that integrates a pre-trained LLM with a lightweight reference model for small marine target detection.
-Employing LoRA adaptation and preference-aware loss, we selectively optimize different feature patches according to their online evaluated learning values in a lightweight form, effectively reducing model overfitting issues in low SCR environments.
-Extensive experimental results demonstrate that RadarLLM significantly outperforms state-of-the-art baselines across sufficient training data and limited training data settings in various detection scenarios.
-These results highlight the considerable potential of LLMs in radar signal processing, especially for challenging small marine target detection.
-Looking ahead, several promising research directions emerge:
-  1. 1.
-    Future studies should investigate the use of more sophisticated open source pre-trained LLMs, such as LLaMA2 or LLaMA3 [44].
-These much larger models may offer enhanced parameter transfer capabilities compared to GPT2.
-  2. 2.
-    Although the present approach demonstrates the effectiveness of the selected reference model, there remains considerable potential to explore alternative architectures.
-Future investigations could examine models such as Mamba [45] or leverage iterative training frameworks, which may further enhance the performance and training efficiency of the reference model.
-  3. 3.
-    To address the storage demands of using large language models (LLMs), we suggest employing model compression strategies, including quantization, pruning, and knowledge distillation.
-These methods aim to reduce model size while maintaining performance, facilitating the deployment of LLMs in resource-constrained environments.
+V Conclusion
+In this study, we propose RadarPLM, a novel fine-tuning framework to adapt PLMs for marine radar target detection.
+We design a lightweight fine-tuning module for lightweight adaptation and a preference-aware loss for selective optimization.
+Extensive experimental results demonstrate that RadarPLM significantly outperforms state-of-the-art baselines across sufficient training data and limited training data settings in various detection scenarios.
+These results highlight the remarkable potential of PLMs as a general optimization solver for radar signal processing.
+Looking ahead, future research will focus on leveraging larger open-source PLMs, and incorporating model compression techniques such as quantization, pruning, and knowledge distillation to further enhance RadarPLM’s scalability and deployment efficiency.

### v4

@@
-RadarPLM: Adapting Pretrained Language Models for Marine Radar Target Detection with Preference-aware Loss
+RadarPLM: Adapting Pre-trained Language Models for Marine Radar Target Detection by Selective Fine-tuning
 Abstract
-Recent advances in pre-trained language models (PLMs) have demonstrated their capabilities in capturing universal knowledge, making them promising applications for radar signal processing.
+Recent advances in pre-trained language models (PLMs) have demonstrated their capabilities in capturing universal knowledge, making them promising for radar signal processing applications.
 Nevertheless, directly fine-tuning PLMs on radar signals is both computationally expensive and prone to overfitting, particularly in low signal-to-clutter ratio (SCR) environments.
 In this paper, we propose a novel fine-tuning framework for PLM-based marine radar target detection.
-First, we design a lightweight adaptation module, enabling parameter-efficient fine-tuning while preserving the pretrained model’s general knowledge.
-Second, a novel preference-aware loss is developed to selectively optimize different feature patches based on their online evaluated learning values, guiding the model to concentrate on the most generalizable feature patterns during optimization.
-Extensive experiments on real-world marine radar datasets demonstrate that the proposed finetuning framework achieves an average performance improvement of 9.9% over the standard approach under low SCR conditions.
-Furthermore, the fine-tuned model, RadarPLM, consistently outperforms state-of-the-art detectors, particularly when training data are limited.
+First, we design a lightweight adaptation module, enabling computationally efficient fine-tuning while preserving the pre-trained model’s general knowledge.
+Second, a novel preference-aware loss is developed to selectively optimize different feature patches based on their online-evaluated learning values, guiding the model to concentrate on those generalizable feature patterns during optimization.
+Finally, a binary classification head is retrained based on autoencoder network to further enhance detection performance.
+Experiments on real-world radar data show that the proposed RadarPLM framework yields at least a 6.35% improvement in detection performance over the existing networks under low SCR conditions.
+Especially, in small training samples cases, the proposed RadarPLM also achieves significant advantage over existing networks owing to the incorporation of the PLM.
 Index Terms:
@@
 I Introduction
-With the growth of global economic integration, maritime transportation has become crucial for international trade and security.
-Real-time and precise detection of maritime targets is essential for intelligent monitoring systems and various applications such as search and rescue, environmental monitoring, and national security.
-Radar technology, known for its ability to operate under diverse weather conditions and around the clock, plays a key role in maritime target detection.
-However, radar systems are challenged by sea clutter, which significantly obscures small targets and complicates detection.
-The complex nature of sea clutter makes it difficult to distinguish weak targets from sea clutter, posing a persistent challenge in radar target detection.
-Detecting small maritime targets, including small boats, frogmen, and fairway buoys, remains an active research area, as these targets have small radar cross sections and are difficult to identify in cluttered environments.
-The early methods for detecting marine targets are typically developed based on the fundamental concept of constant false alarm rate (CFAR), with specific implementations such as CA-CFAR and GO-CFAR.
-These methods operate by estimating clutter statistics from reference cells to dynamically adjust detection thresholds.
-Although computationally efficient, these methods exhibit significant limitations in complex marine environments.
-Empirical studies reveal that their actual false alarm rates (FAR) often deviate drastically from preset parameters, sometimes exceeding theoretical values by orders of magnitude.
-This inadequacy stems from their reliance on simplistic statistical assumptions, which fail to hold in heterogeneous sea clutter scenarios characterized by spatially varying distributions and nonstationary behaviors.
-With the development of modern radars with high range and velocity resolution, approaches leveraging multi-domain features of radar echoes have gained significant attention, utilizing phase, Doppler, and time-frequency domain characteristics to distinguish targets from sea clutter.
-Many researchers have actively engaged in developing approaches that utilize hand-crafted features and machine learning classifiers [1, 2, 3, 4, 5, 6].
-However, these methods rely heavily on domain expertise and hand-crafted heuristics, often struggling to capture high-dimensional and complex patterns in the signal feature.
-With the rapid development of deep learning technology, deep learning (DL) and neural networks (NN) have gained huge popularity in the computer vision (CV) society and natural language processing (NLP) society.
-Moreover, DL and NN approaches are also introduced for small marine target detection.
-These methods primarily utilize radar signal features to train deep neural networks, eliminating the need for precise mathematical models of clutter and targets.
-As a result, they somewhat reduce the impact of complex environmental conditions on target detection performance by learning to extract meaningful and distinctive features in a high-dimensional space.
-Some studies use convolutional neural networks (CNNs) [7, 8, 9, 10, 11, 12, 13], recurrent neural networks (RNNs) [14], and graph neural networks (GNNs) [15, 16, 17].
+Radar-based marine small target detection faces significant challenges due to the complex sea environment, especially under low signal-to-clutter ratio (SCR) conditions.
+Early research, exemplified by constant false alarm rate (CFAR) detectors, hinges on the precise modeling of statistical distributions to distinguish between targets and clutter.
+However, accurately characterizing the non-stationary and non-Gaussian distribution of sea clutter is inherently difficult, while distribution mismatches often result in excessive false alarm rates.
+Furthermore, as these traditional techniques typically exploit only a single amplitude feature, their detection performance is severely compromised in low SCR scenarios.
+With the development of modern wide-band radars featuring high range and velocity resolution, approaches leveraging multi-domain features of radar echoes have gained significant attention.
+A series of hand-crafted features [1, 2, 3, 4, 5], including the phase, Doppler, and time-frequency domain characteristics, have been devised for target detection.
+However, these methods rely heavily on hand-crafted heuristics and often fail to capture the nonlinear structures inherent in radar signal features.
+Deep learning methods have been increasingly applied to small marine radar target detection, as they can directly and effectively learn discriminative nonlinear structures from radar signal features, offering a more flexible and powerful alternative to traditional hand-crafted approaches.
+Some studies use convolutional neural networks (CNNs) [6, 7, 8, 9], recurrent neural networks (RNNs) [10], and graph neural networks (GNNs) [11].
 Among all the methods mentioned above, detectors based on sequence features and RNNs [10] or CNNs [9] offer an excellent trade-off between inference speed and detection performance.
-Despite recent progress, compact NN models with small parameter scales and randomly initialized weights often exhibit limited learning capabilities and unstable training processes.
-These limitations lead to poor detection performance, particularly in scenarios with scarce training data or with significant environmental variability.
-Extensive recent studies [18, 19, 20, 21, 22, 23, 24] have shown that pre-trained language models (PLMs) can be good starting point for wireless signal and time series modeling, as they are pre-trained on massive sequential datasets.
-Compared to random initialized weights of small models, this starting point may significantly improve model optimization stability and effectiveness.
-Moreover, the self-attention mechanism in PLM exhibits structural similarity to Principal Component Analysis (PCA) [18], demonstrating strong generality in extracting key features from high-dimensional sequence data.
-This insight opens a very promising direction for radar target detection by fine-tuning PLMs, offering the potential for performance improvements.
-Although powerful and promising, some challenges still exist.
-First, the large scale of PLM requires lightweight and efficient fine-tuning strategies, which are essential for practical deployment, especially in resource-constrained settings.
-Second, for practical radar applications, adapting PLMs to process sequence features in low signal-to-clutter ratio (SCR) conditions often leads to severe overfitting issues.
-This phenomenon is fundamentally attributed to the intrinsic characteristics of sequence features:
+Recently, the research community has experienced a paradigm shift from lightweight, task-specific neural networks toward large-scale foundation models.
+In particular, pre-trained large language models (LLMs), such as ChatGPT and DeepSeek, have demonstrated strong generalization capabilities across a wide range of complex tasks, including long-form text generation, mathematical reasoning, and code synthesis [12].
+Motivated by these advances, an increasing body of research has extended LLMs beyond the textual domain to handle non-textual data modalities, giving rise to multimodal large language models (MLLMs).
+In parallel, several recent studies [13, 14, 15] have explored the adaptation of pre-trained large language models for time-series analysis and wireless signal processing.
+These works consistently report notable performance improvements over conventional lightweight task-specific models, particularly in scenarios characterized by limited labeled data.
+Building upon these developments, this work investigates the feasibility and effectiveness of leveraging PLMs as a powerful backbone network for marine radar signal processing, aiming to enhance model robustness and generalization under challenging data conditions.
+However, some challenges exist in adapting PLM for marine radar signal processing applications.
+First, owing to the large parameter scale of PLMs, their adaptation to marine radar signal processing necessitates lightweight and efficient fine-tuning strategies, particularly for deployment in resource-constrained environments.
+Such strategies are expected to significantly reduce the number of trainable parameters by several orders of magnitude, while preserving the pre-trained knowledge of PLMs.
+Second, adapting PLMs to process radar signal features in low SCR conditions often leads to severe overfitting issues.
+This phenomenon is fundamentally attributed to the intrinsic characteristics of radar signal features:
   • Significant Clutter-Induced Noise: In low SCR environments, small targets are frequently obscured by ocean waves, which inevitably generate mixed-statue segments.
@@
 Consequently, uniform optimization across all feature patterns may cause the model to overfit not only to noisy patterns but also to overly simplistic ones, which fail to capture the more generalizable patterns inherent in radar sequence features.
-To overcome the above-mentioned issues, including the demand for lightweight fine-tuning and the tendency of overfitting under low SCR conditions, we develop a novel framework named RadarPLM.
-The proposed framework effectively bridges the gap between PLMs and radar target detection tasks through lightweight adaptation module and selective training.
+To overcome the above-mentioned issues, including the demand for lightweight fine-tuning and the tendency of overfitting under low SCR conditions, we develop a novel framework named RadarPLM (based on GPT2, an advanced open-source PLM.).
+The proposed framework effectively bridges the gap between PLMs and radar target detection tasks with a lightweight adaptation module and a selective training strategy.
 First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens.
-Next, a radar target detector is constructed based on a PLM, which is fine-tuned through a lightweight adaptation module to efficiently align with the radar detection task.
-For model optimization, we design a preference-aware loss function by incorporating a lightweight reference model, which is inspired by recent advances [25, 26].
-This loss introduces a selective training strategy that dynamically adjusts the learning weights of different feature patches according to their online-evaluated learning values, significantly mitigating model overfitting.
+Subsequently, a radar target detector is built upon GPT2, a highly advanced open-source PLM, and is fine-tuned through a lightweight adaptation module to efficiently align with the specific radar detection task.
+For model optimization, we design a preference-aware loss function, which is inspired by recent advances [16, 17].
+This loss introduces a selective training strategy dynamically adjusts the learning weights of different feature patches according to their online-evaluated learning values, significantly mitigating model overfitting.
 Finally, the binary classification head is retrained to further improve detection performance.
+Our key contributions are summarized as follows:
+  1. 1.
+    We propose a novel framework RadarPLM based on GPT2 model for marine radar signal processing applications.
+To the best of our knowledge, this work is the first to demonstrate that the recent developed PLMs can serve as a strong starting point for training and optimizing marine radar signal features.
+  2. 2.
+    For fine-tuning, we develop a lightweight adaptation module and design a novel selective training strategy, to reduce finetuning computational overhead and significantly mitigate the risk of model overfitting.
+  3. 3.
+    Extensive experiments validate that RadarPLM achieves new state-of-the-art detection performance on popular marine radar datasets, especially under practical limitations including limited training data and low SCR conditions.
 Our preliminary study [18] demonstrated that fine-tuning a pre-trained GPT2 model effectively extracts discriminative information from sequence features, achieving state-of-the-art detection performance.
-However, model overfitting remains a challenge.
-To address this, we propose the RadarLLM framework, which introduces a lightweight adaptation module and selective training strategy for more reliable and generalizable adaptation.
-Here we summarize our key contributions as follows:
-  1. 1.
-    We propose RadarPLM, a novel framework combining a lightweight reference model and a fine-tuned PLM for marine radar target detection.
-To the best of our knowledge, we are the first to show that the well-known LLMs can be a strong starting point for intelligent radar signal processing.
-  2. 2.
-    For effective fine-tuning, we develop a lightweight adaptation module together with a novel preference-aware loss function, which jointly reduce computational overhead and effectively mitigate the risk of model overfitting.
-  3. 3.
-    Our extensive experiments validate that RadarLLM achieves new state-of-the-art detection performance on popular marine radar datasets under both sufficient and limited training data settings.
+In this paper, we further introduce a lightweight adaptation module and selective training strategy for more lightweight and generalizable adaptation.
 The remainder of this paper is structured as follows.
@@
 II Related Work
-II-A Deep Learning Powered Marine Radar Target Detection
+II-A Marine Radar Target Detection Based on Deep Learning
 Deep learning models with elaborately crafted architectures have demonstrated great promise in marine target detection.
@@
 These challenges motivate us to first explore the use of PLMs for radar target detection.
-II-B Language Model Powered Signal Processing Task
-Recently, LLMs such as ChatGPT and DeepSeek have demonstrated exceptional capabilities in natural language understanding, code generation, math problem solving, and human-like text generation [30].
-Some studies also demonstrate their potential in the signal processing (SP) task.
+II-B Signal Processing Models Based on Pre-trained Language Model
+Recently, PLMs such as GPT2 and LLaMA2 have demonstrated exceptional capabilities in signal processing and communication tasks.
 For example, Zhou et al.
@@
 IOT-LLM and Penetrative AI [27, 28] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks.
-However, recent studies [35] have raised concerns about the universal applicability of PLM in signal processing tasks.
+However, few of these models are specifically devised for radar signal processing applications.
+Moreover, recent studies [29] have raised significant concerns regarding the universal applicability and direct transferability of PLMs to various signal processing tasks.
 PLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks.
@@
 Figure 2: Overview of the reference model.
+(a) IP feature
+(b) DSE feature
+(c) SMS feature
+(d) Amp feature
+(e) DP feature (Target)
+(f) DP feature (Clutter)
+Figure 3: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX database.
 Figure 4: An illustration of the patching operation.
@@
 As illustrated in Fig.
-4, the five extracted sequence features exhibit pronounced discriminability between target and clutter samples.
+3, the five extracted sequence features exhibit pronounced discriminability between target and clutter samples.
 To better structure the input for downstream processing, the extracted features are partitioned into non-overlapping patches.
@@
 An illustration of this patching process is provided in Fig.
-3.
-(a) IP feature
-(b) DSE feature
-(c) SMS feature
-(d) Amp feature
-(e) DP feature (Target)
-(f) DP feature (Clutter)
-Figure 4: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX database.
+4.
 III-C Lightweight Fine-Tuning Module for the PLM Backbone
@@
 This loss difference naturally leads to a selective training scheme with desirable properties:
-  • High Importance (Unlearned): If $\mathcal{L}_{\theta_{{t}}}$ is much larger than $\alpha\mathcal{L}_{\theta_{{t}}}$, the feature token is identified as containing valuable, yet unlearned patterns.
+  • High Importance (Unlearned): If $\mathcal{L}_{\theta_{{t}}}$ is much larger than $\alpha\mathcal{L}_{\theta_{{r}}}$, the feature token is identified as containing valuable, yet unlearned patterns.
   • Low Importance (Noise/Anomaly): If $\alpha\mathcal{L}_{\theta_{{r}}}$ is large (indicating with high inherent noise), it effectively dampens $s_{b,k}$, even if $\mathcal{L}_{\theta_{{t}}}$ is large, preventing the model from overfitting to noisy or anomalous data.
@@
 Problem Definition: Inspired by [17], we begin by formalizing an optimization problem of selecting useful feature patches for model fine-tuning.
-At each training step $t$, given a candidate set $B_{t}$, the objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ maximally reduces the generalization loss on the unseen test set $D_{\mathrm{te}}$.
+At each training step $t$, given a candidate set $B_{t}$, the objective is to select a feature patch $(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}$ whose inclusion in the training set $D_{t}$ maximally reduces the generalization loss on the unseen holdout set $D_{\mathrm{ho}}$.
 This can be expressed as the following optimization problem:
-$\arg\min_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}-\log p\left(y^{\text{te }}\mid F^{\text{te }};D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right).$ (16)
+$\arg\min_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}-\log p\left(y^{\text{ho }}\mid F^{\text{ho }};D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right).$ (16)
 However, directly optimizing Eq.(16) is computationally prohibitive, as it requires retraining the model for each candidate patch.
@@
 B. Derivation of the Surrogate Objective: Applying Bayes’ theorem, the term in Eq.(16) can be decomposed as:
-$-\log p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};D_{t}\!\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\bigr)$
-$\quad=-\log\frac{p\!\bigl(y,y^{\mathrm{te}}\mid F,F^{\mathrm{te}},D_{t}\bigr)}{p\!\bigl(y\mid F;D_{t}\bigr)}.$ (17)
+$-\log p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};D_{t}\!\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\bigr)$
+$\quad=-\log\frac{p\!\bigl(y,y^{\mathrm{ho}}\mid F,F^{\mathrm{ho}},D_{t}\bigr)}{p\!\bigl(y\mid F;D_{t}\bigr)}.$ (17)
 Assuming that each label is statistically independent of all other labels in the corpus when conditioned on its corresponding feature vector, the joint probability term in the numerator can be expressed as follows:
-$p\!\bigl(y,y^{\mathrm{te}}\mid F,F^{\mathrm{te}},D_{t}\bigr)\\ =p\!\bigl(y\mid F;F^{\mathrm{te}},y^{\mathrm{te}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};F,D_{t}\bigr)\\ =p\!\bigl(y\mid F;y^{\mathrm{te}},F^{\mathrm{te}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{te}}\mid F^{\mathrm{te}};D_{t}\bigr).$ (18)
+$p\!\bigl(y,y^{\mathrm{ho}}\mid F,F^{\mathrm{ho}},D_{t}\bigr)\\ =p\!\bigl(y\mid F;F^{\mathrm{ho}},y^{\mathrm{ho}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};F,D_{t}\bigr)\\ =p\!\bigl(y\mid F;y^{\mathrm{ho}},F^{\mathrm{ho}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};D_{t}\bigr).$ (18)
 Substituting Eq.(18) into Eq.(17) and taking the logarithm yields:
-$\log p\!\left(y^{\text{te}}\mid F^{\text{te}};\,D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right)$ (19)
-$=\log\frac{p\left(y\mid F;y^{\text{te }},F^{\text{te }},D_{t}\right)p\left(y^{\text{te }}\mid F^{\text{te }},D_{t}\right)}{p\left(y\mid F;D_{t}\right)}$
-$\;\propto\;\;\log p\left(y\mid F;y^{\mathrm{te}},F^{\mathrm{te}},D_{t}\right)-\log p\left(y\mid F;D_{t}\right)$
-$\;\propto\;\;\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t},D_{\text{te}}\right].$
-In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the test samples in Eq.(16) can be approximated by the following objective:
-$\arg\max_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right].$ (20)
-As exact Bayesian inference is intractable in neural networks, we fit the models with SGD instead.
-Here, $\mathcal{L}[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}]$ denotes the training loss on the patch using the current model trained on $D_{t}$, whereas $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right]$ represents the excepted irreducible holdout loss obtained if the model were trained on both the training set and the test set.
+$\log p\!\left(y^{\text{ho}}\mid F^{\text{ho}};\,D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right)$ (19)
+$=\log\frac{p\left(y\mid F;y^{\text{ho }},F^{\text{ho }},D_{t}\right)p\left(y^{\text{ho }}\mid F^{\text{ho }},D_{t}\right)}{p\left(y\mid F;D_{t}\right)}$
+$\;\propto\;\;\log p\left(y\mid F;y^{\mathrm{ho}},F^{\mathrm{ho}},D_{t}\right)-\log p\left(y\mid F;D_{t}\right)$
+$\;\propto\;\;\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t},D_{\text{ho}}\right].$
+In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the holdout samples in Eq.(16) can be approximated by the following objective:
+$\arg\max_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{ho}}\right].$ (20)
+Here, $\mathcal{L}[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}]$ denotes the training loss on the patch using the current model trained on $D_{t}$, whereas $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{ho}}\right]$ represents the excepted irreducible holdout loss obtained if the model were trained on both the training set and the unseen holdout set.
+This objective reveals that the difference between the training loss and the irreducible holdout loss effectively quantifies the reducible component of the generalization error contributed by each feature patch.
+Consequently, this loss discrepancy can be rigorously interpreted as the learning value of the patch, representing the marginal improvement in performance on unseen data expected from its inclusion in the training process.
 C.
-Conclusion: Eq.(20) reveals that the gap between the training loss and the irreducible holdout loss quantifies the reducible component of the generalization error of each feature patch.
-Consequently, this loss difference can be rigorously interpreted as the learning value of each feature patch, representing how much its inclusion in model training is expected to improve performance on unseen data.
-However, accessing the test set for training on $D_{t}\cup D_{\mathrm{test}}$ is both prohibited and computationally impractical.
-To address this limitation, the validation set is adopted as a statistically sound and feasible proxy for the unseen data distribution.
-Accordingly, we approximate $\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{test}}\right]$ using the loss computed by a lightweight reference model trained on the validation set $D_{\mathrm{val}}$ [26].
-Consequently, the resulting token weight derived in Eq.
-(14) provides a practical and theoretically grounded measure of each feature patch’s learning value.
+Practical Implementation and Approximation: However, realizing the irreducible loss in practice for deep neural networks presents several significant challenges.
+We introduce the following strategic approximations to enable an efficient and tractable implementation for calculating the irreducible loss inspired by [17]:
+  1. 1.
+    The theoretical derivation relies on Bayesian inference and conditioning.
+As exact Bayesian inference is intractable in neural network, we fit the model using Stochastic Gradient Descent (SGD) instead.
+  2. 2.
+    In addition, we adopt the validation set ($D_{\mathrm{val}}$) as the holdout set to serve as a statistically sound proxy for the unseen data distribution.
+  3. 3.
+    Furthermore, since training the PLM-based model on the combined set $D_{t}\cup D_{\mathrm{val}}$ is computationally expensive, we approximate it using the loss computed by a lightweight transformer-based model (reference model) trained on $D_{\mathrm{val}}$.
+D.
+Conclusion: Consequently, the calculated gap between the target loss $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})$ and the approximated irreducible loss $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})$ (derived from the lightweight reference model) serves as the practical measure of each feature patch’s learning value, which forms the basis for the token weighting strategy defined in Eq.(14).
 III-D3 Comparison with Existing Loss Designs
@@
 As shown in Fig.
-1 and Section III-A, the PLM-based model outputs $[B,K,2]$ token-level predictions, which lack a unified detection representation.
-To achieve consistent decision-making, we retrain a binary classifier head using the learned representation $F^{\mathrm{LN}}$.
-An autoencoder-based biary classification head extracts informative features from this high-dimensional input for robust detection.
+1 and Section III-A, the PLM-based network produces token-level outputs of shape $[B,K,2]$, which is inconvenient for final decision.
+To obtain a binary output for each sample, we therefore retrain an autoencoder-based binary classification head on top of feature representation $F^{\mathrm{LN}}$.
 It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative representation; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector.
@@
 2:Radar echo signals $\{x(n)\}_{n=1}^{N}$ and Target labels $\{y_{b}\}_{b=1}^{B}$
-3:Pre-trained LLM parameters $\theta_{\text{LLM}}$
+3:PLM parameters $\theta_{\text{PLM}}$
 4:Stage 1: Sequence Feature Extraction
@@
   2. 2.
-    Compute the loss for each feature token of the training samples based on the predicted output probabilities of reference model, see Eq.(12).
+    Compute the loss for each feature token of the training samples based on the predicted output probabilities of $\theta_{r}$, see Eq.(12).
 6:Stage 3: Fine-tune PLM for Target Detection
@@
   2. 2.
-    Fine-tune PLM for target detection via preference-aware loss function, see Eq.(14) and Eq.(15).
+    Fine-tune PLM for target detection via lightweight adaptation module and preference-aware loss function, see Eq.(14) and Eq.(15).
 7:Stage 4: Binary Classification Head Retraining
@@
 8:
-9:Fine-tuned RadarLLM model with controllable FAR.
+9:Fine-tuned RadarPLM model with controllable FAR.
 IV Experiments
@@
 Eval batch size 400
-Epochs for LLM fine-tuning 500
+Epochs for RadarPLM fine-tuning 500
 Epochs for Autoencoder training 300
@@
 Learning rate for Autoencoder training 0.00001
-TABLE III: Hyperparameters for Autoencoder network.
-Block Layer Filter Stride Padding Output
-[…]
-[…]
-[…]
-[…]
-[…]
-Conv6+ReLU 16,1*1 1 Same 55*16
-Flatten+Fully-Connected1 - - - 64
-Fully-Connected2 - - - 20
-Decoder Fully-Connected3+Reshape - - - 55*16
-Conv1+ReLU 32,1*1 1 1 55*32
-Conv2+ReLU 64,1*1 1 1 55*64
-Conv3+ReLU 128,1*1 1 1 55*128
-Conv4+ReLU 256,1*1 1 1 55*256
-Conv5+ReLU 512,1*1 1 1 55*512
-Conv6+ReLU 768,1*1 1 1 55*768
-Classifier GeLU+LayerNorm - - - 20
-Fully-Connected4 - - - 2
 In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarPLM’s detection performance under two settings: sufficient training data and limited training data scenarios.
@@
 For patching, the size of feature patches is set to 48.
-For the pre-trained LLM, the smallest version of GPT2 with $F=768$ feature dimension and the first $L=4$ layers of the GPT2 encoder are deployed.
-The network hyperparameters of the Autoencoder are shown in Table III.
+For the PLM, the smallest version of GPT2 with $F=768$ feature dimension and the first $L=4$ layers of the GPT2 encoder are deployed.
 The value of $\alpha$ in Eq.(14) is set to 0.9.
@@
 All network training experiments are conducted on a system equipped with an E5-2695v3 CPU, an NVIDIA 3090Ti GPU, and 64 GB of RAM.
-IV-B Performance Comparison on Sufficient Training Samples
+IV-B Performance Comparison
 (1). HH
@@
 (4). VV
-Figure 6: Comparison of model detection performance on various datasets.
-TABLE IV: Comparison of detection rates (%) on sufficient training samples.
-Method $H$ HV VH $V$ AVG
-RNN [14] 46.04 46.66 38.18 17.04 36.98
-Bi-LSTM [14] 70.39 73.15 76.48 60.89 72.48
-GRU [41] 76.46 77.55 73.96 67.48 73.86
-Transformer [36] 52.95 71.73 67.32 55.67 61.92
-PatchTST [42] 75.23 75.72 77.49 66.59 73.76
-ResNet18 [10] 66.24 80.96 82.31 73.97 75.87
-OFA [18] 74.19 80.38 79.71 70.36 76.16
-ADN18 [43] 79.74 83.50 82.01 78.02 80.82
-RadarPLM 82.51 84.81 84.61 75.90 81.96
-(All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.)
-TABLE V: Network parameters (training parameters / total parameters) and average inference time of different methods.
+Figure 6: Comparison of model detection performance on various datasets. The accompanying ROC curves demonstrate that the RadarPLM framework achieves the best overall detection performance.
+TABLE III: Comparison of detection rates (%) on sufficient training samples under different signal-to-clutter ratio (SCR) conditions.
+Method All Data Low SCR Data
+$H$ HV VH $V$ AVG $H$ HV VH $V$ AVG
+RNN [10] 46.04 46.66 38.18 17.04 36.98 9.45 13.00 11.67 4.47 9.65
+Bi-LSTM [10] 70.39 73.15 76.48 60.89 72.48 39.38 42.47 37.48 25.27 36.15
+GRU [35] 76.46 77.55 73.96 67.48 73.86 48.99 50.46 49.79 27.33 44.14
+Transformer [30] 52.95 71.73 67.32 55.67 61.92 32.31 41.89 46.43 15.35 34.00
+PatchTST [36] 75.23 75.72 77.49 66.59 73.76 49.27 49.71 53.18 31.46 45.91
+ResNet18 [9] 66.24 80.96 82.31 73.97 75.87 50.45 56.18 53.73 44.54 51.23
+OFA [13] 74.19 80.38 79.71 70.36 76.16 49.28 54.18 54.62 26.96 46.26
+ADN18 [37] 79.74 83.50 82.01 78.02 80.82 54.76 60.04 31.32 46.11 47.31
+RadarPLM 82.51 84.81 84.61 75.90 81.96 60.03 63.19 62.27 44.82 57.58
+Note: All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.
+TABLE IV: Network parameters (training parameters / total parameters) and average inference time of different methods.
 Detector Avg. inference time (s) Total Params (M) Trainable Params (M)
@@
 ADN18 [37] 160.9018 14.3315 14.3315
-RadarLLM (ours) 5.6561 69.3932 2.4567
-Table IV reports the detection rates achieved by RadarLLM and the eight baseline detectors introduced in Section IV-A2.
+RadarPLM 5.6561 69.3932 2.4567
+Table III reports the detection rates achieved by RadarPLM and the eight baseline detectors introduced in Section IV-A2.
 All results are obtained with a fixed false alarm rate of 0.005 on the eight IPIX datasets.
-Compared to the best competing sequence feature detector, RadarLLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively.
-Figure LABEL:DT presents the complete ROC curves; it is evident that RadarLLM maintains the highest detection probability across the full range of false alarm rates.
-We also compare RadarLLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone.
-Despite using only compact sequence features, RadarLLM still delivers a 1.14% gain in the average detection rate.
+Compared to the best competing sequence feature detector, RadarPLM improves the detection rate by 7.60%, 3.85%, 2.30%, and 1.93% for the HH, HV, VH, and VV polarizations, respectively.
+Fig.
+6 presents the complete ROC curves; it is evident that RadarPLM maintains the highest detection probability across the full range of false alarm rates.
+We also compare RadarPLM with ADN18, which relies on high-dimensional time–frequency maps and an enhanced CNN backbone.
+Despite using only compact sequence features, RadarPLM still delivers a 1.14% gain in the average detection rate.
+More Crucially, RadarPLM demonstrates exceptional robustness in the challenging low SCR conditions.
+As shown in the right section of Table III, while most baselines suffer from severe performance degradation due to strong clutter-induced noise, RadarPLM maintains a remarkable lead.
+It achieves an average detection rate of 57.58% on low SCR datasets, surpassing the second-best method (ResNet18) by 6.35% and the complex ADN18 method by a significant margin of 10.27%.
+This result highlights the critical advantage of our method in learning discriminative representation from noisy sequence features.
 Furthermore, we compare RadarPLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment.
-The comparison results are summarized in Table V.
+The comparison results are summarized in Table IV.
 Although RadarPLM has the highest total network parameters, the number of trainable parameters is comparable to those of other models, as the majority of its weights are frozen during fine-tuning.
@@
 Together, these advantages make RadarPLM a promising solution for real-time marine target detection.
-IV-C Performance Comparison on Limited Training Samples
+IV-C Scarce Training Sample Condition
 Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge.
-Considering this issue, we conducted experiments under constrained training sample conditions to assess the effectiveness of RadarPLM.
-The samples are divided into three groups: (1) a training set using the first 10% of the observation time for both the target and clutter cells, (2) a validation set covering 10% to 15% of the observation time, and (3) a test set containing the remaining samples.
-In Table VI, we report the detection performance of RadarPLM and several baseline methods under scenarios with limited training data.
-Compared with the experimental setup described in Section IV-A1, the advantages of RadarPLM become more pronounced.
+Considering this issue, we conducted experiments under scarce training sample conditions to assess the effectiveness of RadarPLM.
+Specifically, we utilized a data split ratio of 10% (Training samples): 5% (Validation samples): 85% (Test samples), where the number of training samples are severely reduced.
+In Fig.
+7, we present a comparative analysis of all methods under the scarce training sample condition against the standard setting (Abundant Training Samples) described in Section IV-B.
+The results clearly demonstrate that RadarPLM exhibits the second-lowest performance degradation among all assessed methods, highlighting its robustness to scarce data availability.
 Specifically, RadarPLM achieves a substantial increase of 19.66% in the average detection rate over the previously proposed sequence-feature based detector [10].
@@
   • Third, the integration of a preference-aware loss function significantly improves RadarPLM detection performance, especially in challenging scenarios characterized by low SCR.
-TABLE VI: Comparison of detection rates (%) on limited training samples.
-Method $H$ HV VH $V$ AVG
-RNN [14] 35.49 30.13 31.82 15.14 28.15
-Bi-LSTM [14] 61.96 67.46 65.39 42.19 59.25
-GRU [41] 70.43 74.14 74.20 59.47 69.56
-Transformer [36] 70.22 64.56 65.31 51.15 62.81
-PatchTST [42] 68.31 72.30 75.69 61.77 69.52
-ResNet18 [10] 63.71 78.33 78.08 71.05 72.79
-OFA [18] 73.64 76.81 77.44 63.33 72.81
-ADN18 [43] 75.45 81.50 80.81 70.97 77.18
-RadarPLM 79.26 82.31 81.51 72.54 78.91
-(All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.)
-(1). Data1
-(2). Data2
-(3). Data3
-Figure 7: Comparison of detection rates between the proposed preference-aware loss and the conventional cross-entropy loss.
-We compare detection rates across different datasets and epochs, highlighting the improvements achieved by the preference-aware loss function.
-IV-D Ablation Study
-IV-D1 Backbone Architectures
-To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the PLM under sufficient training samples setting:
-  1. 1.
-    RadarPLM (0): In this variant, the PLM backbone is entirely removed, while all other components of the framework are kept unchanged.
-  2. 2.
-    RadarPLM (T): In this variant, the PLM backbone is replaced with a randomly initialized transformer block.
-  3. 3.
-    RadarPLM (R): In this variant, the network architecture and fine-tuning procedure are maintained, but pre-trained parameters of PLM are replaced with random initialization.
-In Table VII, we show the results of the ablation study on backbone choice.
-It can be seen that the removal or exchange of PLM backbone results in notable performance degradation, indicating the necessity of PLM for high detection performance.
-TABLE VII: Results of ablation study on the backbone choice and fine-tuning strategies.
-RadarPLM(0)–RadarPLM(R) denote backbone ablation results, while RadarPLM(F)–RadarPLM(LN) denote fine-tuning strategy ablation results.
+Figure 7: Performance comparison under scarce and abundant training sample conditions.
+The results show that the RadarPLM approach experiences substantially smaller performance degradation under scarce training sample condition while maintaining the best overall detection performance.
+TABLE V: Ablation study on fine-tuning strategy configurations and their corresponding training time (measured over 100 training epochs).
+Model LoRA Layer Norm Other Params Training Time (hour/100 epochs)
+RadarPLM $\checkmark$ $\checkmark$ 0.116
+RadarPLM (F) $\checkmark$ $\checkmark$ 0.148
+RadarPLM (LoRA) $\checkmark$ 0.116
+RadarPLM (LN) $\checkmark$ 0.112
+TABLE VI: Results of ablation study on the backbone choice and fine-tuning strategies.
 Model $H$ HV VH $V$ AVG
@@
 RadarPLM (LN) 82.33 83.73 83.80 75.12 81.25
+Note: RadarPLM(0)–RadarPLM(R) denote backbone ablation results, while RadarPLM(F)–RadarPLM(LN) denote fine-tuning strategy ablation results.
+The best and second-best results are highlighted in red and blue, respectively.
+(1). Data1
+(2). Data2
+(3). Data3
+Figure 8: We compare the detection rates achieved by the proposed preference-aware loss against the conventional cross-entropy (CE) loss across various datasets and training epochs.
+The results demonstrate that while the CE loss exhibits detection rate degradation shortly after a few training epochs (indicating overfitting), the preference-aware loss effectively mitigates overfitting, enabling consistent performance improvement throughout more training epochs.
+TABLE VII: Results of ablation study on the loss function.
+Method Data1 Data2 Data3 AVG
+RadarPLM (PA) 52.06 (+16.8) 54.87 (+5.4) 65.81 (+7.5) 57.58 (+9.9)
+RadarPLM (WCE) 43.51 (+8.8) 53.95 (+4.9) 64.55 (+6.8) 54.84 (+7.7)
+RadarPLM (CE) 34.73 48.96 57.80 47.16
+PatchTST (PA) 47.48 (+8.1) 50.29 (+5.9) 62.42 (+7.0) 53.40 (+7.0)
+PatchTST (WCE) 46.68 (+7.8) 48.96 (+5.0) 57.80 (+2.9) 51.15 (+5.2)
+PatchTST (CE) 38.91 43.92 54.89 45.91
+TABLE VIII: Performance comparison of RadarPLM under different patch sizes and GPT2 layer numbers.
+Parameter Configuration Average DR (%) Average IT (s) Total / Training NP (M)
+Patch Size 32 79.44 5.9259 69.39 / 2.46
+48 81.96 5.6561 69.39 / 2.46
+64 79.09 5.6222 69.39 / 2.46
+GPT2 Layers 0 Layer 76.80 5.3484 41.04 / 2.44
+2 Layers 81.29 5.4812 55.22 / 2.45
+4 Layers 81.96 5.6561 69.39 / 2.46
+6 Layers 80.42 5.7331 83.57 / 2.46
+8 Layers 80.01 5.8655 97.74 / 2.47
+Note: The results include average detection rate (DR), average inference time (IT), and total/training network parameters (NP, in millions).
+The best and second-best results are highlighted in red and blue, respectively.
+IV-D Ablation Study
+IV-D1 Backbone Architectures
+To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the PLM under sufficient training samples setting:
+  1. 1.
+    RadarPLM (0): In this variant, the PLM backbone is entirely removed, while all other components of the framework are kept unchanged.
+  2. 2.
+    RadarPLM (T): In this variant, the PLM backbone is replaced with a randomly initialized transformer block.
+  3. 3.
+    RadarPLM (R): In this variant, the network architecture and fine-tuning procedure are maintained, but pre-trained parameters of PLM are replaced with random initialization.
+In Table VI, we show the results of the ablation study on backbone choice.
+It can be seen that the removal or exchange of PLM backbone results in notable performance degradation, indicating the necessity of PLM for high detection performance.
 IV-D2 Fine-tuning Strategies
-To verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study under the setting of sufficient training samples, as summarized in Table VIII, and the corresponding results are presented in Table VII.
+To verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study under the setting of sufficient training samples, as summarized in Table V, and the corresponding results are presented in Table VI.
 In this experiment, four fine-tuning configurations were evaluated: (1) RadarPLM: jointly fine-tuning the LoRA modules and Layer Normalization (LN) parameters, (2) RadarPLM (LoRA): fine-tuning only the LoRA modules, (3) RadarPLM (LN): fine-tuning only the LN parameters, and (4) RadarPLM (F): full fine-tuning of all model parameters.
@@
 This demonstrates that the proposed design not only enhances model adaptability, but also significantly reduces computational cost.
-TABLE VIII: Ablation study on fine-tuning strategy configurations and their corresponding training time (measured over 100 training epochs).
-Model LoRA Layer Norm Other Params Training Time (hour/100 epochs)
-RadarPLM $\checkmark$ $\checkmark$ 0.116
-RadarPLM (F) $\checkmark$ $\checkmark$ 0.148
-RadarPLM (LoRA) $\checkmark$ 0.116
-RadarPLM (LN) $\checkmark$ 0.112
 IV-D3 Loss Function
@@
 In Fig.
-7, we present the detection rates on test samples from different low-SCR datasets under HH polarization across multiple training epochs.
+8, we present the detection rates on test samples from different low-SCR datasets under HH polarization across multiple training epochs.
 It can be observed that preference-aware loss achieves an improvement of around 1%-12% compared to cross-entropy loss (uniformly optimized on all feature patches) in different training epochs and datasets.
@@
 For further comparison, we include a sample-level reweighting baseline, denoted as WCE [17], where training weights are adjusted at the sample level.
-As summarized in Table IX, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions.
+As summarized in Table VII, the PA loss consistently yields superior performance across all datasets, especially under low SCR conditions.
 For example, RadarPLM (PA) achieves an average gain of 9.9% over RadarPLM (CE), while PatchTST (PA) improves by 7.0% over its CE-based counterpart.
 In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
-TABLE IX: Results of ablation study on the loss function.
-Method Data1 Data2 Data3 AVG
-RadarPLM (PA) 52.06 (+16.8) 54.87 (+5.4) 65.81 (+7.5) 57.58 (+9.9)
-RadarPLM (WCE) 43.51 (+8.8) 53.95 (+4.9) 64.55 (+6.8) 54.84 (+7.7)
-RadarPLM (CE) 34.73 48.96 57.80 47.16
-PatchTST (PA) 47.48 (+8.1) 50.29 (+5.9) 62.42 (+7.0) 53.40 (+7.0)
-PatchTST (WCE) 46.68 (+7.8) 48.96 (+5.0) 57.80 (+2.9) 51.15 (+5.2)
-PatchTST (CE) 38.91 43.92 54.89 45.91
-TABLE X: Performance comparison of RadarPLM under different patch sizes and GPT2 layer numbers.
-The results include average detection rate (DR), average inference time (IT), and total/training network parameters (NP, in millions).
-RadarPLM Performance under Different Configurations
-Patch Size Configuration Average DR (%) Average IT (s) Total / Training NP (M)
-32 79.44 5.9259 69.39 / 2.46
-48 81.96 5.6561 69.39 / 2.46
-64 79.09 5.6222 69.39 / 2.46
-GPT2 Layers 0 Layer 77.41 5.3484 41.04 / 2.44
-2 Layers 81.29 5.4812 55.22 / 2.45
-4 Layers 81.96 5.6561 69.39 / 2.46
-6 Layers 80.42 5.7331 83.57 / 2.46
-8 Layers 80.01 5.8655 97.74 / 2.47
 IV-E Analysis of the Effect of Hyperparameters on Experimental Results
 We first investigate the impact of patch size on RadarPLM’s performance.
-The results of the ablation study are summarized in Table X.
+The results of the ablation study are summarized in Table VIII.
 Using a patch size of 48 achieves the highest detection rate, outperforming configurations with smaller (32) and larger (64) patch sizes.
@@
 We then investigated the impact of the number of GPT2 layers on RadarPLM’s performance.
-Our ablation study on GPT2 layer configurations is presented in Table X.
+Our ablation study on GPT2 layer configurations is presented in Table VIII.
 Activating four GPT2 layers achieves the highest detection rate, outperforming both shallower configurations and deeper ones.
@@
 We design a lightweight fine-tuning module for lightweight adaptation and a preference-aware loss for selective optimization.
-Extensive experimental results demonstrate that RadarPLM significantly outperforms state-of-the-art baselines across sufficient training data and limited training data settings in various detection scenarios.
+Extensive experimental results demonstrate that RadarPLM significantly outperforms state-of-the-art baselines on low SCR datasets and scarce training data settings.
 These results highlight the remarkable potential of PLMs as a general optimization solver for radar signal processing.

### v5

@@
 Nevertheless, directly fine-tuning PLMs on radar signals is both computationally expensive and prone to overfitting, particularly in low signal-to-clutter ratio (SCR) environments.
-In this paper, we propose a novel fine-tuning framework for PLM-based marine radar target detection.
+In this paper, we propose a fine-tuning framework for PLM-based marine radar target detection.
 First, we design a lightweight adaptation module, enabling computationally efficient fine-tuning while preserving the pre-trained model’s general knowledge.
@@
 Experiments on real-world radar data show that the proposed RadarPLM framework yields at least a 6.35% improvement in detection performance over the existing networks under low SCR conditions.
-Especially, in small training samples cases, the proposed RadarPLM also achieves significant advantage over existing networks owing to the incorporation of the PLM.
+Especially, in the small-sample training cases, the proposed RadarPLM also achieves a significant advantage over existing networks owing to the incorporation of the PLM.
 Index Terms:
@@
 In particular, pre-trained large language models (LLMs), such as ChatGPT and DeepSeek, have demonstrated strong generalization capabilities across a wide range of complex tasks, including long-form text generation, mathematical reasoning, and code synthesis [12].
-Motivated by these advances, an increasing body of research has extended LLMs beyond the textual domain to handle non-textual data modalities, giving rise to multimodal large language models (MLLMs).
-In parallel, several recent studies [13, 14, 15] have explored the adaptation of pre-trained large language models for time-series analysis and wireless signal processing.
+Motivated by these advances, an increasing body of research has extended these large models beyond the textual domain to non-textual data modalities, giving rise to multimodal large language models.
+Meanwhile, several recent studies [13, 14] have explored the adaptation of pre-trained large language models for time-series data analysis and wireless-communication signal processing.
 These works consistently report notable performance improvements over conventional lightweight task-specific models, particularly in scenarios characterized by limited labeled data.
-Building upon these developments, this work investigates the feasibility and effectiveness of leveraging PLMs as a powerful backbone network for marine radar signal processing, aiming to enhance model robustness and generalization under challenging data conditions.
-However, some challenges exist in adapting PLM for marine radar signal processing applications.
-First, owing to the large parameter scale of PLMs, their adaptation to marine radar signal processing necessitates lightweight and efficient fine-tuning strategies, particularly for deployment in resource-constrained environments.
-Such strategies are expected to significantly reduce the number of trainable parameters by several orders of magnitude, while preserving the pre-trained knowledge of PLMs.
-Second, adapting PLMs to process radar signal features in low SCR conditions often leads to severe overfitting issues.
-This phenomenon is fundamentally attributed to the intrinsic characteristics of radar signal features:
-  • Significant Clutter-Induced Noise: In low SCR environments, small targets are frequently obscured by ocean waves, which inevitably generate mixed-statue segments.
-The mixed statue segments can cause a low SCR and increase the likelihood of the model overfitting to irrelevant and noisy information.
-  • Feature Pattern Convergence Imbalance: Various patterns in different sequence features often exhibit different convergence rates during fine-tuning.
-For example, simple and regular patterns may rapidly reach convergence during model fine-tuning and then exhibit a tendency for overfitting, whereas other valuable patterns may require more iterations to achieve convergence.
-Consequently, uniform optimization across all feature patterns may cause the model to overfit not only to noisy patterns but also to overly simplistic ones, which fail to capture the more generalizable patterns inherent in radar sequence features.
-To overcome the above-mentioned issues, including the demand for lightweight fine-tuning and the tendency of overfitting under low SCR conditions, we develop a novel framework named RadarPLM (based on GPT2, an advanced open-source PLM.).
+Inspired by these developments, this work investigates the feasibility and effectiveness of leveraging PLMs as a powerful backbone network for marine radar signal processing, aiming to enhance model robustness and generalization under challenging data conditions.
+However, adapting PLMs to marine radar signal processing remains non-trivial, primarily due to two critical challenges: the need for parameter-efficient fine-tuning and the risk of model overfitting under complex real-world environments.
+First, the large parameter scale of PLMs necessitates lightweight and efficient fine-tuning strategies.
+Without such lightweight adaptation, the extensive computational and memory overhead of PLMs fine-tuning would limit their applicability in real-world radar scenarios.
+Lightweight adaptation strategy requires reducing the number of trainable parameters by several orders of magnitude while preserving prior knowledge embedded in PLMs.
+Second, directly adapting PLMs to process radar signal features often leads to severe overfitting.
+This issue arises from the intrinsic characteristics of radar signal features.
+On the one hand, radar echoes typically suffer from low signal-to-noise ratios due to complex environmental noise and sensor limitations.
+Besides, in small marine target detection scenarios, weak target returns are frequently submerged by ocean clutter, producing mixed-status segments that bring noisy or anomalous patterns into the radar signal features.
+On the other hand, different feature patterns exhibit highly imbalanced convergence behaviors during fine-tuning: simple and regular patterns tend to converge rapidly, while more informative but complex patterns remain under-learned.
+Consequently, previous uniform optimization strategy may lead to severe overfitting, not only to noisy or anomalous patterns but also to overly simplistic ones.
+This highlights the necessity of a selective fine-tuning strategy that dynamically adjusts the training emphasis based on the learning value of each feature pattern.
+To overcome the above issues, we develop a novel framework named RadarPLM (based on GPT2, an advanced open-source PLM).
 The proposed framework effectively bridges the gap between PLMs and radar target detection tasks with a lightweight adaptation module and a selective training strategy.
 First, we extract five sequence features from radar echo signals and patch them into multiple feature tokens.
-Subsequently, a radar target detector is built upon GPT2, a highly advanced open-source PLM, and is fine-tuned through a lightweight adaptation module to efficiently align with the specific radar detection task.
-For model optimization, we design a preference-aware loss function, which is inspired by recent advances [16, 17].
-This loss introduces a selective training strategy dynamically adjusts the learning weights of different feature patches according to their online-evaluated learning values, significantly mitigating model overfitting.
-Finally, the binary classification head is retrained to further improve detection performance.
+Subsequently, a radar target detector is built upon GPT2 and is fine-tuned through a lightweight adaptation module to efficiently align with the specific radar detection task.
+To mitigate overfitting, we design a preference-aware loss function, inspired by [15].
+By introducing a selective training strategy to dynamically adjust the learning weights of different feature patches according to their online-evaluated learning values.
+Finally, the binary classification head is trained to further improve detection performance.
 Our key contributions are summarized as follows:
   1. 1.
-    We propose a novel framework RadarPLM based on GPT2 model for marine radar signal processing applications.
-To the best of our knowledge, this work is the first to demonstrate that the recent developed PLMs can serve as a strong starting point for training and optimizing marine radar signal features.
+    We propose a novel framework RadarPLM based on GPT2 model, an advanced open-source PLM, for marine radar signal processing applications.
+To the best of our knowledge, this is the first work to demonstrate that the recent developed PLMs can serve as a strong starting point for model training and optimization on radar signal features.
   2. 2.
-    For fine-tuning, we develop a lightweight adaptation module and design a novel selective training strategy, to reduce finetuning computational overhead and significantly mitigate the risk of model overfitting.
+    For fine-tuning, we develop a lightweight adaptation module and design a novel selective training strategy, to reduce fine-tuning computational overhead and significantly mitigate the risk of model overfitting.
   3. 3.
-    Extensive experiments validate that RadarPLM achieves new state-of-the-art detection performance on popular marine radar datasets, especially under practical limitations including limited training data and low SCR conditions.
-Our preliminary study [18] demonstrated that fine-tuning a pre-trained GPT2 model effectively extracts discriminative information from sequence features, achieving state-of-the-art detection performance.
-In this paper, we further introduce a lightweight adaptation module and selective training strategy for more lightweight and generalizable adaptation.
+    Extensive experiments validate that RadarPLM achieves state-of-the-art detection performance on popular marine radar datasets, especially in low SCR and small-sample training cases.
+Our preliminary study [16] initially demonstrated the promising prospects of PLM for radar signal prcocessing applications.
+In this paper, we further expand that work by introducing a lightweight adaptation module and selective training strategy for more robust and generalizable optimization, achieving better performance.
 The remainder of this paper is structured as follows.
@@
 Section V concludes our work.
+Figure 1: Overview of our RadarPLM framework.
 II Related Work
@@
 These challenges motivate us to first explore the use of PLMs for radar target detection.
-II-B Signal Processing Models Based on Pre-trained Language Model
-Recently, PLMs such as GPT2 and LLaMA2 have demonstrated exceptional capabilities in signal processing and communication tasks.
+II-B Signal Processing Based on Pre-trained Language Models
+Recently, PLMs such as GPT2 and LLaMA2 have demonstrated exceptional capabilities in sequential signal processing and wireless communication tasks.
 For example, Zhou et al.
@@
 Zheng and Dai [24] propose a multi-task PLM framework to satisfy the requirements of different wireless communication tasks.
-IOT-LLM and Penetrative AI [27, 28] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks.
-However, few of these models are specifically devised for radar signal processing applications.
-Moreover, recent studies [29] have raised significant concerns regarding the universal applicability and direct transferability of PLMs to various signal processing tasks.
+IOT-LLM [25] directly utilize raw signal data as input and chain of thought prompts for reasoning on the Internet of Things (IOT) tasks.
+However, none of these models are specifically devised for radar signal processing applications.
+In fact, recent studies [26] have raised significant concerns regarding the universal applicability and direct transferability of PLMs to various signal processing tasks.
 PLM-based approaches do not always deliver significant performance improvements, implying that their pre-trained parameters sometimes fail to transfer to downstream tasks.
-In marine radar target detection under low SCR conditions, directly fine-tuning a PLM often leads to severe overfitting issues, limiting its parameter transfer capability.
+Especially, in marine radar target detection under low SCR conditions, directly fine-tuning a PLM often leads to severe overfitting issues, limiting its parameter transfer capability.
 To mitigate this limitation, we introduce a preference-aware loss that selectively trains on informative feature patches.
 III Proposed Method
-Figure 1: Overview of our RadarPLM framework.
 Figure 2: Overview of the reference model.
-(a) IP feature
-(b) DSE feature
-(c) SMS feature
-(d) Amp feature
-(e) DP feature (Target)
-(f) DP feature (Clutter)
-Figure 3: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX database.
+(a) Instantaneous Phase (IP)
+(b) Doppler Spectrum Entropy (DSE)
+(c) Short-Time Fourier Transform Magnitude Spectrum (SMS)
+(d) Amplitude (Amp)
+(e) Doppler Phase (DP): Target
+(f) Doppler Phase (DP): Clutter
+Figure 3: IP, DSE, SMS, Amp, and DP features for target and sea clutter echo signal on IPIX dataset.
 Figure 4: An illustration of the patching operation.
@@
   • High Importance (Unlearned): If $\mathcal{L}_{\theta_{{t}}}$ is much larger than $\alpha\mathcal{L}_{\theta_{{r}}}$, the feature token is identified as containing valuable, yet unlearned patterns.
-  • Low Importance (Noise/Anomaly): If $\alpha\mathcal{L}_{\theta_{{r}}}$ is large (indicating with high inherent noise), it effectively dampens $s_{b,k}$, even if $\mathcal{L}_{\theta_{{t}}}$ is large, preventing the model from overfitting to noisy or anomalous data.
+  • Low Importance (Noisy/Anomalous patterns): If $\alpha\mathcal{L}_{\theta_{{r}}}$ is large (indicating feature token with inherent noisy/anomalous patterns), it effectively dampens $s_{b,k}$, even if $\mathcal{L}_{\theta_{{t}}}$ is large, preventing model overfitting.
   • Low Importance (Well Learned): If $\mathcal{L}_{\theta_{{t}}}$ is already small, the feature token is considered well learned, and $s_{b,k}$ will also be small, preventing redundant updates to easy patterns.
@@
 $\arg\min_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}-\log p\left(y^{\text{ho }}\mid F^{\text{ho }};D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right).$ (16)
-However, directly optimizing Eq.(16) is computationally prohibitive, as it requires retraining the model for each candidate patch.
+However, directly optimizing (16) is computationally prohibitive, as it requires retraining the model for each candidate patch.
 To enable a practical surrogate, we reformulate this objective using probabilistic decomposition.
-B. Derivation of the Surrogate Objective: Applying Bayes’ theorem, the term in Eq.(16) can be decomposed as:
+B. Derivation of the Surrogate Objective: Applying Bayes’ theorem, the term in (16) can be decomposed as:
 $-\log p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};D_{t}\!\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\bigr)$
@@
 $p\!\bigl(y,y^{\mathrm{ho}}\mid F,F^{\mathrm{ho}},D_{t}\bigr)\\ =p\!\bigl(y\mid F;F^{\mathrm{ho}},y^{\mathrm{ho}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};F,D_{t}\bigr)\\ =p\!\bigl(y\mid F;y^{\mathrm{ho}},F^{\mathrm{ho}},D_{t}\bigr)\,p\!\bigl(y^{\mathrm{ho}}\mid F^{\mathrm{ho}};D_{t}\bigr).$ (18)
-Substituting Eq.(18) into Eq.(17) and taking the logarithm yields:
+Substituting (18) into (17) and taking the logarithm yields:
 $\log p\!\left(y^{\text{ho}}\mid F^{\text{ho}};\,D_{t}\cup(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\right)$ (19)
@@
 $\;\propto\;\;\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\!\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t},D_{\text{ho}}\right].$
-In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the holdout samples in Eq.(16) can be approximated by the following objective:
+In summary, for a model trained on $D_{t}$, the process of identifying the feature patch that minimizes the loss on the holdout samples in (16) can be approximated by the following objective:
 $\arg\max_{(F_{b,k}^{\mathrm{fi}},y_{b}^{k})\in B_{t}}\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{t}\right]-\mathcal{L}\left[y_{b}^{k}\mid F_{b,k}^{\mathrm{fi}};D_{\mathrm{t}},D_{\mathrm{ho}}\right].$ (20)
@@
 D.
-Conclusion: Consequently, the calculated gap between the target loss $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})$ and the approximated irreducible loss $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})$ (derived from the lightweight reference model) serves as the practical measure of each feature patch’s learning value, which forms the basis for the token weighting strategy defined in Eq.(14).
+Conclusion: The calculated gap between the target loss $\mathcal{L}_{\theta_{{t}},b,k}(F^{\mathrm{fi}}_{b,k})$ and the approximated irreducible loss $\mathcal{L}_{\theta_{{r}},b,k}(F^{\mathrm{fi}}_{b,k})$ (derived from the lightweight reference model) serves as a practical measure of each feature patch’s learning value, which forms the token weights in (14).
+Accordingly, the loss gap is employed to adaptively reweight the training objective, guiding the model to focus on learning informative and unlearned feature tokens, while significantly reducing the impact of noisy and trivial patterns.
 III-D3 Comparison with Existing Loss Designs
@@
 1 and Section III-A, the PLM-based network produces token-level outputs of shape $[B,K,2]$, which is inconvenient for final decision.
-To obtain a binary output for each sample, we therefore retrain an autoencoder-based binary classification head on top of feature representation $F^{\mathrm{LN}}$.
+To obtain a binary output for each sample, we therefore retrain an autoencoder-based binary classification head on top of high-dimensional feature representation $F^{\mathrm{LN}}$.
 It comprises: (1) an encoder with convolutional and fully connected layers producing compact discriminative representation; (2) a symmetric decoder ensuring feature consistency through reconstruction; and (3) a classification head operating on the latent vector.
@@
   1. 1.
-    Extract IP, DSE, SMS, Amp, and DP features from the observation vector, respectively, see Eq.(3), (4), (5),(7), and (9).
+    Extract IP, DSE, SMS, Amp, and DP features from the observation vector, respectively, see (3), (4), (5),(7), and (9).
   2. 2.
@@
   2. 2.
-    Compute the loss for each feature token of the training samples based on the predicted output probabilities of $\theta_{r}$, see Eq.(12).
+    Compute the loss for each feature token of the training samples based on the predicted output probabilities of $\theta_{r}$, see (12).
 6:Stage 3: Fine-tune PLM for Target Detection
   1. 1.
-    Compute the loss for each feature token of the training samples based on the predicted output probabilities of current model, see Eq.(13).
+    Compute the loss for each feature token of the training samples based on the predicted output probabilities of current model, see (13).
   2. 2.
-    Fine-tune PLM for target detection via lightweight adaptation module and preference-aware loss function, see Eq.(14) and Eq.(15).
+    Fine-tune PLM for target detection via lightweight adaptation module and preference-aware loss function, see (14) and (15).
 7:Stage 4: Binary Classification Head Retraining
   1. 1.
-    Train autoencoder-based binary classification head based on the weighted sum of reconstruction loss and CE loss, see Eq.(21), Eq.(22), and Eq.(23).
+    Train autoencoder-based binary classification head based on the weighted sum of reconstruction loss and CE loss, see (21), (22), and (23).
   2. 2.
-    Adjust detection threshold $\eta$ for controllable FAR, see Eq.(24).
+    Adjust detection threshold $\eta$ for controllable FAR, see (24).
 8:
@@
 TABLE I: Overview of the IPIX Database.
-Label Name File name Primary Secondary SCR
-1 17 19931107_135603_starea 9 8,10,11 Low
-2 25 19931108_213827_starea 7 6,8 Low
-3 26 19931108_220902_starea 7 6,8 Low
-4 54 19931111_163625_starea 8 7,9,10 High
-5 280 19931118_023604_stareC0000 8 7,9,10 High
-6 283 19931118_035737_stareC0000 10 8,9,11,12 High
-7 311 19931118_162658_stareC0000 7 6,8,9 High
-8 320 19931118_174259_stareC0000 7 6,8,9 High
+Label Name File Name Target Distance Bin SCR
+1 17 19931107_135603_starea 9 Low
+2 25 19931108_213827_starea 7 Low
+3 26 19931108_220902_starea 7 Low
+4 54 19931111_163625_starea 8 High
+5 280 19931118_023604_stareC0000 8 High
+6 283 19931118_035737_stareC0000 10 High
+7 311 19931118_162658_stareC0000 7 High
+8 320 19931118_174259_stareC0000 7 High
 TABLE II: Hyperparameters for network training.
@@
 In this section, we first describe the experimental setup, followed by a comprehensive evaluation of RadarPLM’s detection performance under two settings: sufficient training data and limited training data scenarios.
+TABLE III: Comparison of detection rates (%) on sufficient training samples under different signal-to-clutter ratio (SCR) conditions.
+Method All Data Low SCR Data
+$H$ HV VH $V$ AVG $H$ HV VH $V$ AVG
+RNN [10] 46.04 46.66 38.18 17.04 36.98 9.45 13.00 11.67 4.47 9.65
+Bi-LSTM [10] 70.39 73.15 76.48 60.89 72.48 39.38 42.47 37.48 25.27 36.15
+GRU [32] 76.46 77.55 73.96 67.48 73.86 48.99 50.46 49.79 27.33 44.14
+Transformer [27] 52.95 71.73 67.32 55.67 61.92 32.31 41.89 46.43 15.35 34.00
+PatchTST [33] 75.23 75.72 77.49 66.59 73.76 49.27 49.71 53.18 31.46 45.91
+ResNet18 [9] 66.24 80.96 82.31 73.97 75.87 50.45 56.18 53.73 44.54 51.23
+OFA [13] 74.19 80.38 79.71 70.36 76.16 49.28 54.18 54.62 26.96 46.26
+ADN18 [34] 79.74 83.50 82.01 78.02 80.82 54.76 60.04 31.32 46.11 47.31
+RadarPLM 82.51 84.81 84.61 75.90 81.96 60.03 63.19 62.27 44.82 57.58
+Note: All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.
+(a). HH
+(b). HV
+(c). VH
+(d). VV
+Figure 6: Comparison of model detection performance on various datasets. The accompanying ROC curves demonstrate that the RadarPLM framework achieves the best overall detection performance.
 IV-A Experimental Setup
@@
 To facilitate performance analysis in varying operating conditions, Data1–Data3 correspond to low signal-to-clutter ratio (SCR) scenes, while Data4–Data8 represent high-SCR scenes.
-To ensure sufficient training data, we employ overlapped segmentation following the partition rule in Eq.
-(1), with parameters set to $M=32$ for target cells and $M=128$ for clutter cells.
+To ensure sufficient training data, we employ overlapped segmentation following the partition rule in (1), with parameters set to $M=32$ for target cells and $M=128$ for clutter cells.
 This process generates 4,079 target samples and more than 9,000 clutter samples per dataset.
@@
 For the PLM, the smallest version of GPT2 with $F=768$ feature dimension and the first $L=4$ layers of the GPT2 encoder are deployed.
-The value of $\alpha$ in Eq.(14) is set to 0.9.
+The value of $\alpha$ in (14) is set to 0.9.
 All signal feature extraction operations are performed with MATLAB R2016a software.
 All network training experiments are conducted on a system equipped with an E5-2695v3 CPU, an NVIDIA 3090Ti GPU, and 64 GB of RAM.
-IV-B Performance Comparison
-(1). HH
-(2). HV
-(3). VH
-(4). VV
-Figure 6: Comparison of model detection performance on various datasets. The accompanying ROC curves demonstrate that the RadarPLM framework achieves the best overall detection performance.
-TABLE III: Comparison of detection rates (%) on sufficient training samples under different signal-to-clutter ratio (SCR) conditions.
-Method All Data Low SCR Data
-$H$ HV VH $V$ AVG $H$ HV VH $V$ AVG
-RNN [10] 46.04 46.66 38.18 17.04 36.98 9.45 13.00 11.67 4.47 9.65
-Bi-LSTM [10] 70.39 73.15 76.48 60.89 72.48 39.38 42.47 37.48 25.27 36.15
-GRU [35] 76.46 77.55 73.96 67.48 73.86 48.99 50.46 49.79 27.33 44.14
-Transformer [30] 52.95 71.73 67.32 55.67 61.92 32.31 41.89 46.43 15.35 34.00
-PatchTST [36] 75.23 75.72 77.49 66.59 73.76 49.27 49.71 53.18 31.46 45.91
-ResNet18 [9] 66.24 80.96 82.31 73.97 75.87 50.45 56.18 53.73 44.54 51.23
-OFA [13] 74.19 80.38 79.71 70.36 76.16 49.28 54.18 54.62 26.96 46.26
-ADN18 [37] 79.74 83.50 82.01 78.02 80.82 54.76 60.04 31.32 46.11 47.31
-RadarPLM 82.51 84.81 84.61 75.90 81.96 60.03 63.19 62.27 44.82 57.58
-Note: All results are evaluated at a false alarm rate ($P_{fa}$) of 0.005. The best and second-best results are highlighted in red and blue, respectively.
+IV-B Average Performance Comparison
 TABLE IV: Network parameters (training parameters / total parameters) and average inference time of different methods.
@@
 Despite using only compact sequence features, RadarPLM still delivers a 1.14% gain in the average detection rate.
-More Crucially, RadarPLM demonstrates exceptional robustness in the challenging low SCR conditions.
-As shown in the right section of Table III, while most baselines suffer from severe performance degradation due to strong clutter-induced noise, RadarPLM maintains a remarkable lead.
-It achieves an average detection rate of 57.58% on low SCR datasets, surpassing the second-best method (ResNet18) by 6.35% and the complex ADN18 method by a significant margin of 10.27%.
-This result highlights the critical advantage of our method in learning discriminative representation from noisy sequence features.
 Furthermore, we compare RadarPLM with other baseline models in terms of training complexity, total network parameters, and inference cost to evaluate its practicality for real-world deployment.
@@
 Together, these advantages make RadarPLM a promising solution for real-time marine target detection.
-IV-C Scarce Training Sample Condition
+IV-C Performance Comparison on Low SCR Data
+Owing to the inherent complexity of the marine environment, low-SCR radar echoes are frequently encountered in practical applications.
+In such scenarios, clutter-induced noise gives rise to noisy and anomalous feature patterns, leading to pronounced performance degradation in existing detection algorithms.
+To assess model robustness, we evaluate each approach on more challenging low-SCR datasets (IPIX #17, IPIX #25, and IPIX #26), where the SCR typically falls below 5 dB.
+As shown in the right section of Table III, while most baselines suffer from severe performance degradation due to strong clutter-induced noise, RadarPLM maintains a remarkable lead.
+It achieves an average detection rate of 57.58% on low SCR datasets, surpassing the second-best method (ResNet18) by 6.35% and the complex ADN18 method by a significant margin of 10.27%.
+This result highlights the critical advantage of our method in learning discriminative representation from noisy and anomalous radar signal features.
+IV-D Small-sample Training Condition
 Due to the inherent sparsity of radar targets in real-world scenarios, obtaining high-quality labeled samples for radar target detection presents a substantial challenge.
@@
   • Third, the integration of a preference-aware loss function significantly improves RadarPLM detection performance, especially in challenging scenarios characterized by low SCR.
-Figure 7: Performance comparison under scarce and abundant training sample conditions.
-The results show that the RadarPLM approach experiences substantially smaller performance degradation under scarce training sample condition while maintaining the best overall detection performance.
+Figure 7: Performance comparison under small and abundant sample training conditions.
+The result shows that the RadarPLM approach experiences substantially smaller performance degradation under scarce training sample condition while maintaining the best overall detection performance.
 TABLE V: Ablation study on fine-tuning strategy configurations and their corresponding training time (measured over 100 training epochs).
@@
 The best and second-best results are highlighted in red and blue, respectively.
-(1). Data1
-(2). Data2
-(3). Data3
-Figure 8: We compare the detection rates achieved by the proposed preference-aware loss against the conventional cross-entropy (CE) loss across various datasets and training epochs.
+Figure 8: Comparison of detection rates achieved by the proposed preference-aware loss and the conventional cross-entropy (CE) loss under HH polarization across different datasets ((a) IPIX #17, (b) IPIX #25, and (c) IPIX #26) over training epochs.
 The results demonstrate that while the CE loss exhibits detection rate degradation shortly after a few training epochs (indicating overfitting), the preference-aware loss effectively mitigates overfitting, enabling consistent performance improvement throughout more training epochs.
@@
 The best and second-best results are highlighted in red and blue, respectively.
-IV-D Ablation Study
-IV-D1 Backbone Architectures
+IV-E Ablation Study
+IV-E1 Backbone Architectures
 To validate the effectiveness of the backbone choice, the ablation study is conducted in this section by removing or replacing the PLM under sufficient training samples setting:
@@
 It can be seen that the removal or exchange of PLM backbone results in notable performance degradation, indicating the necessity of PLM for high detection performance.
-IV-D2 Fine-tuning Strategies
+IV-E2 Fine-tuning Strategies
 To verify the effectiveness of the proposed fine-tuning strategy, we conducted an ablation study under the setting of sufficient training samples, as summarized in Table V, and the corresponding results are presented in Table VI.
@@
 This demonstrates that the proposed design not only enhances model adaptability, but also significantly reduces computational cost.
-IV-D3 Loss Function
+IV-E3 Loss Function
 In addition, we compare the fine-tuning effects of preference-aware loss and CE loss on low SCR datasets.
@@
 In particular, token-level reweighting in both PA variants outperforms WCE, highlighting the advantage of fine-grained reweighting in enhancing detection under challenging scenarios.
-IV-E Analysis of the Effect of Hyperparameters on Experimental Results
+IV-F Analysis of the Effect of Hyperparameters on Experimental Results
 We first investigate the impact of patch size on RadarPLM’s performance.
@@
 V Conclusion
-In this study, we propose RadarPLM, a novel fine-tuning framework to adapt PLMs for marine radar target detection.
-We design a lightweight fine-tuning module for lightweight adaptation and a preference-aware loss for selective optimization.
-Extensive experimental results demonstrate that RadarPLM significantly outperforms state-of-the-art baselines on low SCR datasets and scarce training data settings.
-These results highlight the remarkable potential of PLMs as a general optimization solver for radar signal processing.
-Looking ahead, future research will focus on leveraging larger open-source PLMs, and incorporating model compression techniques such as quantization, pruning, and knowledge distillation to further enhance RadarPLM’s scalability and deployment efficiency.
+We propose a fine-tuning framework RadarPLM that adapts the GPT2, an advanced open-source pre-trained language model, for marine radar target detection.
+To ensure lightweight and robust adaptation, we devise a novel lightweight adaptation module, preserving the general knowledge of the GPT2 model while significantly reducing training time and fine-tuning overhead.
+To mitigate overfitting on low-SCR radar signal, we develop a selective training strategy to selectively optimize different feature patches based on their online-evaluated learning values, effectively improving model robustness.
+Extensive experiments on real-world marine radar datasets demonstrate that RadarPLM consistently outperforms SOTA baselines, especially under low SCR conditions, achieving an average detection rate improvement exceeding 6% over the strongest competitors.
+Notably, under small-sample training conditions, RadarPLM surpasses the existing methods by nearly 20% at most, owing to the effective transfer of the universal knowledge in GPT-2 to the marine radar signal processing task.